### v1

# Cubic fourfolds with symplectic automorphisms

Abstract.
We determine projective equations of smooth complex cubic fourfolds with symplectic automorphisms by classifying 6-dimensional projective representations of Laza and Zheng’s 34 groups. In particular, we determine the number of irreducible components for moduli spaces of cubic fourfolds with symplectic actions by these groups. We also discuss the fields of definition of cubic fourfolds in six maximal cases.
1. Introduction
In [LZ22], Laza and Zheng completely classified the groups $\mathrm{Aut}^{s}(X)$ of symplectic automorphisms of complex cubic fourfolds $X$. According to them, we have 34 possible groups as $\mathrm{Aut}^{s}(X)$ with six maximal cases
$3^{4}:\mathrm{A}_{6},\quad\mathrm{A}_{7},\quad 3^{1+4}_{+}:2.2^{2},\quad\mathrm{M}_{10},\quad\mathrm{L}_{2}(11),\quad\mathrm{A}_{3,5}.$
For each group $G$ among the 34 groups, we have the covariant lattice $S_{G}$ defined as the orthogonal complement of the invariant lattice $H^{4}(X,\mathbb{Z})^{G}$. It is uniquely determined (up to isomorphism) by $G$, and the moduli space of cubic fourfolds $X$ such that $G\subset\mathrm{Aut}^{s}(X)$ has dimension $20-\mathrm{rank}\,S_{G}$. However, this moduli space is not necessarily irreducible, corresponding to the fact that there may exist different primitive embeddings $S_{G}\hookrightarrow H^{4}(X,\mathbb{Z})$. For example, there exists the unique cubic fourfold with $\mathrm{Aut}^{s}(X)=G$ for the above maximal groups $G$ except $\mathrm{A}_{7}$ and $\mathrm{M}_{10}$, but two different cubic fourfolds if $\mathrm{Aut}^{s}(X)=\mathrm{A}_{7}$ or $\mathrm{M}_{10}$. Laza and Zheng classified possible finite groups $G=\mathrm{Aut}^{s}(X)$ via embeddings $(G,S_{G})\hookrightarrow(\mathrm{Co}_{0},\mathbb{L})$, where $\mathbb{L}$ is the Leech lattice and $\mathrm{Co}_{0}=\mathrm{O}(\mathrm{L})$ is the Conway group, and the result is based on the classification of the fixed-point lattices for the Leech lattice by Höhn and Mason ([HM16]).
More recently, Yang, Yu and Zhu studied full automorphism groups for cubic fivefolds and fourfolds in [YYZ23]. According to them, there are 15 maximal groups, and a finite group $G$ can acts faithfully on a smooth cubic fourfold if and only if $G$ is a subgroup of these 15 groups. However, explicit equatoins of cubic fourfolds with automorphisms are still unknown in many cases.
In this paper, we study symplectic actions of 34 groups of Laza and Zheng, and detertmine invariant smooth cubic fourfolds. Since any automorphism of a smooth cubic fourfold is given by a projective transformation, our problem is reduced to classification of faithfull projective representations of 34 groups which satisfy symplectic conditions. It is known that complex projective representations of a finite group $G$ lift to linear representations of a Schur cover $G^{*}$. We determine possible projective actions of 34 groups $G$ by considering character tables of $G^{*}$. As the result, we have
Theorem 1.1.
The moduli space of cubic fourfolds with $G\subset\mathrm{Aut}^{s}(X)$ has two irreducible components if $G$ is one of the following groups:
[…]
For other cases, the moduli spaces are irreducible (see Table 1 in sectoin 2 ).
For maximal cases, defining eqautions of corresponding cubic fourfolds are known, and they are defined over $\mathbb{Q}$ in most cases. In the last section, we show
Theorem 1.2.
(1)  If the symplectic automorphisms $\mathrm{Aut}^{s}(X)$ of a cubic fourfold $X$ is maximal, then $X$ is defined over $\mathbb{Q}$ except in the case of $\mathrm{Aut}^{s}(X)=\mathrm{M}_{10}$.
(2)  Let $X_{+}$ and $X_{-}$ be two different cubic fourfolds with $\mathrm{Aut}^{s}(X_{\pm})=\mathrm{M}_{10}$. They are defined over $K=\mathbb{Q}(\sqrt{6})$, and conjugate to each other. Namely, $\mathrm{Gal}(K/\mathbb{Q})$ swiches $X_{+}$ and $X_{-}$.
Our results are heavily dependent on computer system GAP. Also we used databases [AtlasWeb] and [GroupNames] for information of finite groups such as generators and automorphisms.
Notations. We say that $G$ is an extension of $B$ by $A$ and denote $G\cong A.B$ if there is an exact sequence
[…]
(the structure of $A.B$ is not uniquely determined). In particular, it is called a central extension if $\iota(A)$ is included in the center of $G$. We denote $G\cong A:B$ if the exact sequence is split, that is, if $\pi$ has a section.
We use the following notations:
  • $\mathrm{C}_{m},\ \mathrm{D}_{m},\ \mathrm{Q}_{m}$ denote cyclic, dihedral, quaternion groups of order $m$.
  • $\mathrm{C}_{m}$ is often denoted by $m$ (e.g. $4$ denotes $\mathrm{C}_{4}$, and $2^{2}$ denotes $\mathrm{C}_{2}\times\mathrm{C}_{2}$).
  • $p^{1+2n}_{\pm}$ denote extraspecial groups of order $p^{1+2n}$.
  • $\mathrm{S}_{m,n}=\mathrm{S}_{m}\times\mathrm{S}_{n},\ \mathrm{A}_{m,n}=\mathrm{S}_{m,n}\cap\mathrm{A}_{m+n}$, where $\mathrm{A}_{n},\ \mathrm{S}_{n}$ denote alternating, symmetric groups of degree $n$.
  • $\mathrm{diag}(a_{1},\dots,a_{n})$ denotes a diagonal matrix with diagonal entries $a_{1},\dots,a_{n}$.
  • $A\oplus B$ denotes a block matrix $\begin{bmatrix}A&\bf{0}\\ \bf{0}&B\end{bmatrix}$ for square matrices $A$ and $B$
  • $\zeta_{n}=\exp(2\pi i/n),\quad\omega=\zeta_{3}$,  $\widehat{G}=\mathrm{Hom}(G,\mathbb{C}^{\times})$,  $s_{k}(x)=x_{1}^{k}+\cdots+x_{n}^{k}$.
2. Projective actions of finite groups and invariant hypersurfaces
2.1. Schur covering groups
Let $X\subset\mathbb{P}^{n+1}(\mathbb{C})$ be a hypersurface of degree $d$. If $X$ is smooth and $n\geq 2,\ d\geq 3$, then the group $\mathrm{Aut}(X)$ of automorphisms is finite and realized by projective transformations of $\mathbb{P}^{n+1}$ except the case $(n,d)=(2,4)$ ([MM64]). First of all, we recall basic facts on projective representations of finite groups. For details, see [Ka85].
A projective representation of a group $G$ is a homomorphism $\rho:G\rightarrow\mathrm{PGL}_{n}(\mathbb{C})$. If there is a commutative diagram
[…]
of group homomorphisms, we say that $\rho$ is lifted to a linear representation $\tilde{\rho}$. Let us see actions of the alternating group $\mathrm{A}_{7}$ as a typical example. Note that any non-trivial action of a non-commutative simple group on a smooth cubic fourfold is symplectic (see section 2.2).
Example 2.1.
According to [LZ22], there exist two non-isomorphic smooth cubic fourfolds with symplectic action of $\mathrm{A}_{7}$. One is the Clebsch-Segre cubic
(2.1) $\begin{cases}x_{1}+x_{2}+\cdots+x_{7}=0\\ x_{1}^{3}+x_{2}^{3}+\cdots+x_{7}^{3}=0\end{cases}\text{in}\ \mathbb{P}^{6},$
and the symmetric group $\mathrm{S}_{7}$ acts on it as permutations of coordinates (However, the actoin of $(12)\in\mathrm{S}_{7}$ is not symplectic). Another $\mathrm{A}_{7}$-invariant cubic fourfold was identified by Yang, Yu and Zhu in [YYZ23] considering the “triple cover” of $\mathrm{A}_{7}$. Their model is defined over $\mathbb{Q}(\sqrt{15})$, and we give an equation over $\mathbb{Q}$ here. Let $G$ be a subgroup of $\mathrm{GL}_{6}(\mathbb{C})$ generated by two matrices
(2.2) $a=\begin{bmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&\omega&0&0&0\\ 0&0&0&\omega&0&0\\ 0&0&0&0&\omega^{2}&0\\ 0&0&0&0&0&\omega^{2}\\ \end{bmatrix},\quad b=\frac{1}{3}\begin{bmatrix}-1&0&-\omega&0&3\omega^{2}&2\omega^{2}\\ 2&0&2\omega&0&0&-\omega^{2}\\ -\omega^{2}&0&2&3&0&-\omega\\ 2\omega^{2}&0&-1&0&0&2\omega\\ 2\omega&3\omega&-\omega^{2}&0&0&-1\\ -\omega&0&2\omega^{2}&0&0&2\\ \end{bmatrix}.$
We can check that $G\cong 3.\mathrm{A}_{7}$ by GAP. More precisely, the center $\mathrm{C}_{3}$ is generated by $(abab^{-1})^{2}=\mathrm{diag}(\omega,\dots,\omega)$, and we have a commutative diagram:
[…]
The group $G\cong 3.\mathrm{A}_{7}$ is a non-split extension, and this projective action of $\mathrm{A}_{7}$ cannot be lifted to a linear representation of $\mathrm{A}_{7}$ itself. We have the unique invariant cubic 4-fold $X$:
(2.3) $f(x)=2(x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{4}^{3}+x_{5}^{3}+x_{6}^{3})\\ +3(x_{1}^{2}x_{2}+x_{1}x_{2}^{2}+x_{3}^{2}x_{4}+x_{3}x_{4}^{2}+x_{5}^{2}x_{6}+x_{5}x_{6}^{2})\\ +2(x_{2}x_{3}x_{5}+x_{1}x_{4}x_{5}+x_{1}x_{3}x_{6}+x_{2}x_{4}x_{6})\\ +4(x_{1}x_{3}x_{5}+x_{2}x_{4}x_{5}+x_{2}x_{3}x_{6}+x_{1}x_{4}x_{6})=0.$
Considering $X$ by reduction modulo 2, it is easily checked that $X$ is smooth (In fact, $X$ is isomorphic to the Fermat cubic over $\mathbb{F}_{4}$). $\square$
In general, projective representations (over $\mathbb{C}$) of a finite group $G$ can be lifted to linear representations of a central extension of $G$, called a Schur covering group. More precisely, a group $G^{*}$ is called a Schur covering group (or representation group) of $G$ if $G^{*}$ is a central extension
[…]
such that $M\cong\mathrm{H}^{2}(G,\mathbb{C}^{\times})$ and $\iota(M)\subset[G^{*},G^{*}]$. In general, $G^{*}$ is not uniquely determined. The Klein’s four group $\mathrm{C}_{2}\times\mathrm{C}_{2}$ has two Schur covering groups $\mathrm{D}_{8}$ and $\mathrm{Q}_{8}$. However, it is known that all projective representations of $G$ can be lifted to linear representations of one Schur covering group $G^{*}$. An abelian group $M=M(G)$ is called the Schur multiplier of $G$. For example, we have
$M(\mathrm{A}_{n})=\begin{cases}1\quad(n\leq 3)\\ \mathrm{C}_{6}\quad(n=6,7)\\ \mathrm{C}_{2}\quad(\text{others})\end{cases},\quad M(\mathrm{S}_{n})=\begin{cases}1\quad(n\leq 3)\\ \mathrm{C}_{2}\quad(n\geq 4)\end{cases}$
(see [Ka85]). If $M(G)=1$, we have $G^{*}=G$ and any projective representation of $G$ is lifted to a linear representation of $G$ itself.
Remark 2.2.
It is known that $G$ has only one Schur covering group if the orders of $G/[G,G]$ and $M(G)$ are coprime ([Ka85], Theoerm 4.4).
For non-maximal cases of Laza and Zheng’s 34 groups, we have the following table:
Table 1. Schur covers and multipliers
$r(G)$ $G$ GAP-ID $M$ ID* Eq irr
8 2 2,1 1 2,1 (3.1) 1
12 $2^{2}$ 4,2 2 8,3 (3.2) 1
12 3 3,1 1 3,1 (3.3), (3.4) 2
14 4 4,1 1 4,1 (3.5) 1
14 $\mathrm{S}_{3}$ 6,1 1 6,1 (3.6) 1
15 $\mathrm{D}_{8}$ 8,3 2 16,7 (3.7) 1
16 $\mathrm{A}_{3,3}$ 18,4 3 54,8 (4.6), (4.11) 2
16 $\mathrm{D}_{12}$ 12,4 2 24,8 (4.2), (4.3) 2
16 $\mathrm{A}_{4}$ 12,3 2 24,3 (4.20), (4.22) 2
16 $\mathrm{D}_{10}$ 10, 1 1 10, 1 (4.1) 1
17 $\mathrm{S}_{4}$ 24,12 2 48,28 (5.2), (5.3) 1
17 $\mathrm{Q}_{8}$ 8,4 1 8,4 (5.1) 1
18 $23^{1+4}_{+}:$ 2 486, 249 $3^{5}$ (6.1) 1
18 $\mathrm{A}_{4,3}$ 72,43 6 432,258 (6.5), (6.7) 2
18 $\mathrm{A}_{5}$ 60, 5 2 120,5 (6.9), (6.10) 2
18 $43^{2}.$4 36,9 3 108,15 (6.11), (6.12) 2
18 $\mathrm{S}_{3,3}$ 36,10 2 72,22 (6.13) 1
18 $\mathrm{F}_{21}$ 21,1 1 21,1 (6.2) 1
18 $\mathrm{Hol}(5)$ 20, 3 1 20,3 (6.15) 1
18 $\mathrm{QD}_{16}$ 16,8 1 16,8 (6.3) 1
19 $2.23^{1+4}_{+}:$ 2.2 972, 776 $3^{4}$ (7.1) 1
19 $\mathrm{A}_{6}$ 360,118 6 - (2.6), (2.9) 2
19 $\mathrm{L}_{2}(7)$ 168,42 2 336,114 (7.2) 1
19 $\mathrm{S}_{5}$ 120,34 2 240, 89 (7.3), (7.6) 2
19 $\mathrm{M}_{9}$ 72,41 3 216,88 (7.9) 1
19 $\mathrm{N}_{72}$ 72,40 2 144,117 (7.10) 1
19 $\mathrm{T}_{48}$ 48,29 1 48,29 (7.12) 1
where $r(G)=\mathrm{rank}\ S_{G}$ (see [LZ22]), GAP-ID is “IdGroup($G$)” in GAP, $M$ is the Schur multiplier of $G$, and ID* is GAP-ID of a Schur covering group $G^{*}$ given by “SchurCover($G$)” in GAP. Equations of corresponding cubic 4-folds are given in Eq, and irr is the number of irreducible components of the moduli space $\mathcal{M}(G)$ of cubic fourfolds $X$ such that $G\subset\mathrm{Aut}^{s}(X)$.
2.2. Symplectic automorphisms of cubic 4-folds
Let $X$ be a smooth cubic fourfold. Then $\mathrm{Aut}(X)$ acts faithfully on $\mathrm{H}^{4}(X,\mathbb{Z})$. If $g\in\mathrm{Aut}(X)$ acts trivially on $\mathrm{H}^{3,1}(X)\cong\mathbb{C}$, the action of $g$ is called symplectic. Therefore we have an exact sequence
$1\longrightarrow\mathrm{Aut}^{s}(X)\longrightarrow\mathrm{Aut}(X)\longrightarrow\mathrm{Aut}(\mathrm{H}^{3,1}(X))=\mathbb{C}^{\times}.$
for the group $\mathrm{Aut}^{s}(X)$ of symplectic automorphisms. In particular, a non-trivial action of a non-commutative simple group is necessarily symplectic.
In [Fu16], smooth cubic fourfolds with symplectic automorphisms of primary order were classified by Fu. The result was generalized to symplectic automorphisms of any order in [LZ22]. According to them, the order of a symplectic automorphism is one of
$1,2,3,4,5,6,7,8,9,11,12,15$
If a cubic fourfold $X$ has a symplectic automorphism of oder $n=9,11,12,15$, then we have
$\mathrm{Aut}^{s}(X)=3^{4}:\mathrm{A}_{6},\quad\mathrm{L}_{2}(11),\quad 3^{1+4}:2.2^{2},\quad\mathrm{A}_{3,5},$
respectively, and the cubic fourfold $X$ is unique for these cases. Symplectic automorphisms $g$ of order $n\leq 8$ are classified into the following diagonalized normal forms:
Table 2. Normal forms of symplectic automorphisms
$n$ $(k_{1},\cdots,k_{6})$ $|\mathrm{Tr}\,g|$
2 (0,0,0,0,1,1) 2
3 (0,0,0,0,1,2) 3
3 (0,0,0,1,1,1) 3
3 (0,0,1,1,2,2) 0
4 (0,0,2,2,1,3) 0
$n$ $(k_{1},\cdots,k_{6})$ $|\mathrm{Tr}\,g|$
5 (0,0,1,2,3,4) 1
6 (0,3,2,5,4,4) 2
6 (3,3,0,0,2,4) 1
7 (1,5,4,6,2,3) 1
8 (0,4,2,6,1,3) $\sqrt{2}$
where $g=\mathrm{diag}(\zeta_{n}^{k_{1}},\cdots,\zeta_{n}^{k_{6}})$ and $|\mathrm{Tr}\,g|=|\zeta_{n}^{k_{1}}+\cdots+\zeta_{n}^{k_{6}}|$. Note that symplectic automorphisms of order $n\leq 8$ are conjugate to these normal forms as elements of $\mathrm{PGL}_{6}(\mathbb{C})$. For example, let us consider $g=abab^{-1}$ with $a,b$ in Example 2.1. The order of $g$ is 2 in $\mathrm{PGL}_{6}(\mathbb{C})$ since we have $g^{2}=\mathrm{diag}(\omega^{2},\cdots,\omega^{2})$. Eigenvalues of $g$ are
$\omega,\ \omega,\ \omega,\ \omega,\ -\omega,\ -\omega,$
and $g$ is conjugate to $\mathrm{diag}(1,1,1,1,-1,-1)$ in $\mathrm{PGL}_{6}(\mathbb{C})$. Also note that, although we can not define the trace of $g\in\mathrm{PGL}_{6}(\mathbb{C})$, the absolute value $|\mathrm{Tr}\,\tilde{g}|$ for a lift $\tilde{g}\in\mathrm{GL}_{6}(\mathbb{C})$ does not depend on a lifting if $\tilde{g}$ is of finite order. These values are usefull information to classify symplectic actions from character tables.
2.3. Invariants
Let $G$ be one of Laza and Zheng’s 34 groups. For our purpose, we need to consider 6-dimensional linear representations of $G^{*}\cong M.G$ such that $M$ acts as scalar matrices and the corresponding projective action of $G$ is faithful. In fact, we need only triple cover of $G$ as considered in [YYZ23]. Let us see this in general situation. In the following, $G$ denotes a finite group.
Definition 2.3.
For a homomorphism $\rho:G\rightarrow\mathrm{GL}_{n}(\mathbb{C})$ and $\chi\in\widehat{G}$, we say that a homogeneous polynomial $F(x)\in\mathbb{C}[x_{1},\dots,x_{n}]$ is a $\rho(G)$-semi-invariant with character $\chi$ if
$F(\rho(g)x)=\chi(g)F(x)\quad\text{for}\quad{}^{\forall}g\in G.$
If $\chi$ is trivial, we say that $F(x)$ is a $\rho(G)$-invariant.  We denote the vectore space of $\rho(G)$-semi-invariants of degree $d$ with character $\chi$, by $V_{d}(\rho,\chi)$. If $\chi$ is trivial, it is denoted by $V_{d}(\rho)$.
Remark 2.4.
(1)  In the above definition, it is obvious that $\chi$ factors through $\rho$, that is, there exists $\chi_{\rho}\in\widehat{\rho(G)}$ such that $\chi=\chi_{\rho}\circ\rho$. Moreover, $\chi_{\rho}$ is uniquely determined by $\chi$ and $\rho$.
(2)  For a Schur cover $G^{*}=M.G$ of $G$, we have $M\subset[G^{*},G^{*}]$ and any homomorphism from $G^{*}$ to an abelian group factors through $G$. Therefore we have $\widehat{G^{*}}\cong\widehat{G}$.
Now let $\rho:G\rightarrow\mathrm{PGL}_{n}(\mathbb{C})$ be a projective representation, and $\tilde{\rho}:G^{*}\rightarrow\mathrm{GL}_{n}(\mathbb{C})$ be a lifting of $\rho$, where $G^{*}$ is a Schur cover with the multiplier $M\subset G^{*}$. Assume that there is $\chi\in\widehat{G^{*}}$ such that $V_{d}(\tilde{\rho},\chi)\neq 0$.
Lemma 2.5.
In the above situation, we have
(1)  $\tilde{\rho}(m)^{d}=\mathbf{1}_{n}$ for any $m\in M$.
(2)  $\tilde{\rho}(M)\cong\mathrm{C}_{d}$ or $\tilde{\rho}(M)\cong 1$ if $d$ is prime.
Proof.
(1)  By the assumption, $\tilde{\rho}(m)$ is a scalar matrix $\mathrm{diag}(c,\cdots,c)$. For a non-zero $F(x)\in V_{d}(\tilde{\rho},\chi)$, we have
$F(\tilde{\rho}(m)x)=c^{d}F(x)$
and hence $\chi(m)=c^{d}$. Since $\mathrm{Im}\,\chi$ is abelian, we have
$M\subset[G^{*},G^{*}]\subset\mathrm{Ker}\,\chi.$
Therefore we have $c^{d}=\chi(m)=1$, that is, $c=(\zeta_{d})^{k}$ for some $k\in\mathbb{Z}$.
(2) follows from (1). ∎
For example, the Schur multiplier of $\mathrm{A}_{7}$ is $\mathrm{C}_{6}$, and any projective representation $\rho:\mathrm{A}_{7}\rightarrow\mathrm{PGL}_{n}(\mathbb{C})$ lifts to a linear representation $\tilde{\rho}:6.\mathrm{A}_{7}\rightarrow\mathrm{GL}_{n}(\mathbb{C})$. If $A_{7}$ acts on a cubic hypersurface in $\mathbb{P}^{n-1}(\mathbb{C})$ via $\rho$, then $\tilde{\rho}$ factors through $3.\mathrm{A}_{7}=6.\mathrm{A}_{7}/\mathrm{C}_{2}$.
We have the following simple facts that will be used to reduce the amount of calculation of invariants.
Proposition 2.6.
Let $\rho:G\rightarrow\mathrm{GL}_{n}(\mathbb{C})$ be a representation.
(1)  For $\chi\in\widehat{G}$ and $\sigma\in\mathrm{Aut}(G)$, we have
$V_{d}(\rho,\chi)=V_{d}(\rho\circ\sigma,\,\chi\circ\sigma).$
(2)  For $\chi,\,\theta\in\widehat{G}$, we have
$V_{d}(\rho,\,\chi)=V_{d}(\rho\otimes\theta,\,\chi\otimes\theta^{d}).$
In particular, $\rho$ and $\rho\otimes\theta$ have same semi-invariants if $\theta^{d}=1$.
(3)  Let us assume that $d$ and $|\widehat{G}|$ are co-prime to each other. Then, for any $\chi\in\widehat{G}$, there exist $\theta\in\widehat{G}$ such that
$V_{d}(\rho,\,\chi)=V_{d}(\rho\otimes\theta).$
Therefore, any semi-invariant is regarded as an invariant for a representation which is projectively equivalent to $\rho$.
Proof.
(1) and (2) follow from definitions.
(3)  By the assumption, there is $e\in\mathbb{N}$ such that
$ed\equiv 1\mod|\widehat{G}|.$
For $\theta=\chi^{-e}$, we have
$V_{d}(\rho,\,\chi)=V_{d}(\rho\otimes\theta,\,\chi\otimes\theta^{d})=V_{d}(\rho\otimes\theta,\,\chi^{1-ed})=V_{d}(\rho\otimes\theta).$
2.4. Moduli spaces
Let us consider moduli spaces of hypersurfaces with action of finite group $G$ after [YZ20]. For a representation $\rho:G\rightarrow\mathrm{GL}_{n}(\mathbb{C})$, we have the centralizer and the normalizer of $\rho(G)$ in $\mathrm{GL}_{n}(\mathbb{C})$:
$\mathrm{C}_{\rho}(G)$ =$=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ hgh^{-1}=g\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\},$
$\mathrm{N}_{\rho}(G)$ =$=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ h\rho(G)h^{-1}=\rho(G)\}.$
Moreover we define the stabilizer of $V_{d}(\rho,\,\chi)$:
$\mathrm{N}_{\rho,\chi}(G)$ =$=\{h\in\mathrm{N}_{\rho}(G)\ |\ \chi_{\rho}(hgh^{-1})=\chi_{\rho}(g)\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\}$
where $\chi_{\rho}\circ\rho=\chi$ (see Remark 2.4). The quotient group $\mathrm{N}_{\rho}(G)/\mathrm{C}_{\rho}(G)$ is isomorphic to a subgroup of $\mathrm{Aut}(\rho(G))$. Hence the index $|\mathrm{N}_{\rho}(G):\mathrm{C}_{\rho}(G)|$ is finite, and so is $|\mathrm{N}_{\rho,\chi}(G):\mathrm{C}_{\rho}(G)|$. Let $V_{d}(\rho,\,\chi)^{sm}$ be the subset of semi-invariants which define smooth hypersurfaces of degree $d$. Then the quotient $V_{d}(\rho,\,\chi)^{sm}/\mathrm{N}_{\rho,\chi}(G)$ is the moduli space $\mathcal{M}_{d}(\rho,\chi)$ of smooth hypersurfaces with the action of $G$ with respect to $(\rho,\,\chi)$.
Now let $G$ be one of Laza and Zheng’s 34 groups, $G^{*}$ be a Schur cover and $\rho:G^{*}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ be a lifting of a faithful projective actoin of $G$ on $\mathbb{P}^{5}$. If $\rho(G^{*})$ acts simplectically on a smooth cubic fourfold $F(x)=0$ with $F\in V_{3}(\rho,\,\chi)$, we have
(2.5) $20-r(G)=\dim\mathcal{M}_{3}(\rho,\chi)=\dim V_{3}(\rho,\,\chi)-\dim\mathrm{N}_{\rho,\chi}(G^{*}).$
Note that we have $\dim\mathrm{N}_{\rho,\chi}(G^{*})\geq 1$ since ×$\mathbb{C}^{\times}\subset\mathrm{N}_{\rho,\chi}(G^{*})$. Therefore we have the following criterion to exclude unnecessary cases.
Lemma 2.7.
If we have
$\dim V_{3}(\rho,\,\chi)\leq 20-r(G),$
then $F\in V_{3}(\rho,\,\chi)$ does not define a smooth cubic fourfold with a symplectic action of $\rho(G^{*})$.
Remark 2.8.
(1) We have $\dim\mathrm{N}_{\rho,\chi}(G)=\dim\mathrm{C}_{\rho}(G)$, and $\dim\mathrm{C}_{\rho}(G)$ is easily computable. If we have a decomposition
$\rho=m_{1}\rho_{1}+\cdots+m_{k}\rho_{k}$
into irreducible representations $\rho_{i}$, then we have
$\dim\mathrm{C}_{\rho}(G)=\dim(\mathrm{GL}_{m_{1}}(\mathbb{C})\times\cdots\times\mathrm{GL}_{m_{k}}(\mathbb{C}))=m_{1}^{2}+\cdots+m_{k}^{2}$
by Schur’s Lemma.
(2) Obviously, we have
$\mathrm{N}_{\rho\otimes\chi}(G)=\mathrm{N}_{\rho}(G),\qquad\mathrm{C}_{\rho\otimes\chi}(G)=\mathrm{C}_{\rho}(G)$
for $\chi\in\widehat{G}$.
2.5. $\mathrm{A}_{6}$-invariant cubic fourfolds
Let us consider $G=\mathrm{A}_{6}$ as an example. According to [LZ22], we have 1-dimensional family $\mathcal{M}(\mathrm{A}_{6})$ of cubic fourfolds $X$ such that $\mathrm{A}_{6}\subset\mathrm{Aut}^{s}(X)$, but it may have two or more irreducible components.
By Lemma (2.5), an action of the Schur cover $6.\mathrm{A}_{6}$ on a cubic fourfold factors through $3.\mathrm{A}_{6}=6.\mathrm{A}_{6}/\mathrm{C}_{2}$ whose GAP-ID is (1080, 260). All we have to do is classify 6-dimensional linear representations of $3.\mathrm{A}_{6}$ such that the center $\mathrm{C}_{3}$ acts as scalars. Since $\widehat{3.\mathrm{A}_{6}}=\widehat{\mathrm{A}_{6}}$ is trivial, we do not have a semi-invariant with a non-trivial character. Let us look at the character table of $3.\mathrm{A}_{6}$.
$\begin{array}[]{c|rrrrrrrrrrrrrrrrr}\hline\cr\rm class&\rm 1&\rm 3A&\rm 3B&\rm 2A&\rm 6A&\rm 6B&\rm 3C&\rm 3D&\rm 12A&\rm 12B&\rm 4A&\cdots\\ \rm size&1&1&1&45&45&45&120&120&90&90&90&\cdots\\ \hline\cr\chi_{1}&1&1&1&1&1&1&1&1&1&1&1&\cdots\\ \chi_{2}&3&3\overline{\omega}&3\omega&-1&-\overline{\omega}&-\omega&0&0&\overline{\omega}&\omega&1&\\ \chi_{3}&3&3\overline{\omega}&3\omega&-1&-\overline{\omega}&-\omega&0&0&\overline{\omega}&\omega&1&\\ \chi_{4}&3&3\omega&3\overline{\omega}&-1&-\omega&-\overline{\omega}&0&0&\omega&\overline{\omega}&1&\\ \chi_{5}&3&3\omega&3\overline{\omega}&-1&-\omega&-\overline{\omega}&0&0&\omega&\overline{\omega}&1&\\ \chi_{6}&5&5&5&1&1&1&-1&2&-1&-1&-1&\\ \chi_{7}&5&5&5&1&1&1&2&-1&-1&-1&-1&\\ \chi_{8}&6&6\overline{\omega}&6\omega&2&2\overline{\omega}&2\omega&0&0&0&0&0&\\ \chi_{9}&6&6\omega&6\overline{\omega}&2&2\omega&2\overline{\omega}&0&0&0&0&0&\\ \vdots&\vdots&\\ \hline\cr\end{array}$
The following infomations are omitted:
   $\cdot$ irreducible representations whose dimensions are greater than 6,
   $\cdot$ conjugacy classes for order $5$ and $15$.
Classes 1, 3A and 3B form the center $\mathrm{C}_{3}$, and $\chi_{i}\ (i=1,6,7)$ are irreducible characters with $\mathrm{C}_{3}\subset\mathrm{Ker}\,\chi_{i}$, that is, characters of $\mathrm{A}_{6}$ itself rather than the triple cover $3.\mathrm{A}_{6}$. In this case, we have two (non-trivial) characters $\chi_{1}+\chi_{6}$ and $\chi_{1}+\chi_{7}$ of degree 6. These characters are afforded by the following two representations. One is permutatons of projective coordinates $[x_{1}:x_{2}:x_{3}:x_{4}:x_{5}:x_{6}]$ denoted by $\rho$, and another is the composition of $\rho$ and an outer automorphism $\sigma$ of $\mathrm{A}_{6}$ (see Appendix). We have $V_{3}(\rho)=V_{3}(\rho\circ\sigma)$ (see Proposition 2.6) and this vector space is given by
(2.6) $\mathrm{Span}\{x_{1}^{3}+\cdots+x_{6}^{3},\quad(x_{1}^{2}+\cdots+x_{6}^{2})(x_{1}+\cdots+x_{6}),\quad(x_{1}+\cdots+x_{6})^{3}\}.$
This linear system contains the Fermat cubic, and a general menber is smooth. Since $\mathrm{A}_{6}$ is a non-commutative simple group, any action is symplectic (Although $\mathrm{S}_{6}$ acts on these cubics, the action of $(12)\in\mathrm{S}_{6}$ is not symplectic since it is diagonalized into $\mathrm{diag}(1,1,1,1,1,-1)$, see Table 2 ). We have $\dim\mathrm{C}_{\rho}(3.\mathrm{A}_{6})=1^{2}+1^{2}=2$ (see Remark 2.8) and
$\dim V_{3}(\rho)-\dim\mathrm{C}_{\rho}(3.\mathrm{A}_{6})=1=20-r(\mathrm{A}_{6})$
as desired.
Next we consider 6-dimensional representations such that a class 3A acts as a scalar $\omega$ or $\omega^{2}$. We have eight cases:
$2\chi_{2},\quad\chi_{2}+\chi_{3},\quad 2\chi_{3},\quad 2\chi_{4},\quad\chi_{4}+\chi_{5},\quad 2\chi_{5},\quad\chi_{8},\quad\chi_{9}.$
The image of class 4A in $\mathrm{PGL}_{6}(\mathbb{C})$ remains a class of order $4$, and (the absolute value of) the trace must be 0 if the action is symplectic (see Table 2). Since we have
$\chi_{i}(4\mathrm{A})=\begin{cases}1\quad(i=2,3,4,5)\\ 0\quad(i=8,9)\end{cases},$
only $\chi_{8}$ and $\chi_{9}$ are admissible. From the character table, they are complex conjugate to each other. We give corresponding matrix representations $\rho$ and $\overline{\rho}$ as a subgroup of $G\cong 3.\mathrm{A}_{7}$ considered in Example 2.1.
For $a$ and $b$ in (2.2), using GAP, we can check that elements
(2.7) $s=a^{2}(ba^{2}b^{2}a)^{2}a,\quad t=b^{2}a^{2}ba$
generate a subgroup $H\cong 3.\mathrm{A}_{6}$. The $H$-invariant cubic forms are spanned by $f(x)$ in (2.3) and
(2.8) $g(x)=3(x_{1}x_{2}x_{3}+x_{1}x_{2}x_{4}+x_{1}x_{2}x_{5}+x_{1}x_{2}x_{6}+x_{1}x_{3}x_{4}+x_{1}x_{3}x_{5}+x_{1}x_{5}x_{6}\\ +x_{2}x_{3}x_{4}+x_{2}x_{5}x_{6}+x_{3}x_{5}x_{6}+x_{3}x_{4}x_{6}+x_{3}x_{4}x_{5}+x_{4}x_{5}x_{6})\\ +4(x_{1}x_{4}x_{5}+x_{1}x_{3}x_{6}+x_{2}x_{3}x_{5}+x_{2}x_{4}x_{6})\\ +5(x_{1}x_{4}x_{6}+x_{2}x_{3}x_{6}+x_{2}x_{4}x_{5})+8x_{1}x_{3}x_{5}.$
This linear system contains a smooth cubic fourfolds $f(x)=0$, and a general menber is smooth. Moreover we have $H=\overline{H}$ since
$\overline{s}=t^{2}st^{-1}(st)^{2}t,\qquad\overline{t}=st(st^{-1})^{2}stst^{-1}s.$
Namely we have
(2.9) $V(\rho)=V(\overline{\rho})=\mathbb{C}f(x)\oplus\mathbb{C}g(x).$
From the above, we see that $\mathcal{M}(A_{6})$ has two irreducible components.
3. Groups of small $r(G)$
3.1. Groups of $r(G)\leq 14$
Let $G$ be one of small groups
$\mathrm{C}_{2},\quad\mathrm{C}_{2}\times\mathrm{C}_{2},\quad\mathrm{C}_{3},\quad\mathrm{C}_{4},\quad\mathrm{S}_{3}.$
Cubic fourfolds $X$ such that $G\subset\mathrm{Aut}^{s}(X)$ are determined in [Fu16] and [LZ22] as follows.
(i)  $G=\mathrm{C}_{2}$. We may assume that $G$ is generated by $\mathrm{diag}(1,1,1,1,-1,-1)$, and invariant cubics are
(3.1) $F_{1}(x_{1},x_{2},x_{3},x_{4})+x_{5}^{2}L_{1}(x_{1},x_{2},x_{3},x_{4})\\ +x_{5}x_{6}L_{2}(x_{1},x_{2},x_{3},x_{4})+x_{6}^{2}L_{3}(x_{1},x_{2},x_{3},x_{4})=0.$
(ii)  $G=\mathrm{C}_{2}\times\mathrm{C}_{2}$. We may assume that $G$ is generated by $\mathrm{diag}(1,1,1,1,-1,-1)$ and $\mathrm{diag}(1,1,1,-1,-1,1)$, and invariant cubics are
(3.2) $F_{1}(x_{1},x_{2},x_{3})+x_{4}^{2}L_{1}(x_{1},x_{2},x_{3})+x_{5}^{2}L_{2}(x_{1},x_{2},x_{3})\\ +x_{6}^{2}L_{3}(x_{1},x_{2},x_{3})+x_{4}x_{5}x_{6}=0$
(iii)  $G=\mathrm{C}_{3}$. There are two families. One is
(3.3) $F_{1}(x_{1},x_{2},x_{3},x_{4})+x_{5}^{3}+x_{6}^{3}+x_{5}x_{6}L_{1}(x_{1},x_{2},x_{3},x_{4})=0$
and $G$ is generated by $\mathrm{diag}(1,1,1,1,\omega,\omega^{2})$. Another is
(3.4) $F_{1}(x_{1},x_{2})+F_{2}(x_{3},x_{4})+F_{3}(x_{5},x_{6})+\sum_{i=1,2;j=3,4;k=5,6}a_{ijk}x_{i}x_{j}x_{k}=0$
and $G$ is generated by $\mathrm{diag}(1,1,\omega,\omega,\omega^{2},\omega^{2})$.
(iv)  $G=\mathrm{C}_{4}$. We may assume that $G$ is generated by $\mathrm{diag}(1,1,-1,-1,i,-i)$ and invariant cubics are
(3.5) $\mathrm{Span}\{x_{1}N_{1}(x_{3},x_{4}),\ x_{2}N_{2}(x_{3},x_{4}),\ F_{1}(x_{1},x_{2}),\\ x_{5}x_{6}L_{1}(x_{1},x_{2}),\ x_{5}^{2}L_{2}(x_{3},x_{4}),\ x_{6}^{2}L_{3}(x_{3},x_{4})\}.$
(v)  $G=\mathrm{S}_{3}$. We have the following normal forms. The symmetric group $\mathrm{S}_{3}$ acts on projective coordinates
$[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$
by permuting $x_{i}$ and $y_{i}$ simultaneously. Invariant cubics are
(3.6) $\mathrm{Span}\{s_{1}(x)^{3},\ s_{1}(x)s_{2}(x),\ s_{3}(x),\ s_{1}(y)^{3},\ s_{1}(y)s_{2}(y),\ s_{3}(y),\\ s_{1}(x)s_{1}(y)^{2},\ s_{1}(x)s_{2}(y),\ s_{1}(x)^{2}s_{1}(y),\ s_{2}(x)s_{1}(y),\\ \sum_{i=1}^{3}x_{i}y_{i}^{2},\ \sum_{i=1}^{3}x_{i}^{2}y_{i},\ \sum_{\sigma\in\mathrm{A}_{3}}x_{\sigma(1)}y_{\sigma(2)}y_{\sigma(3)},\ \sum_{\sigma\in\mathrm{A}_{3}}x_{\sigma(1)}x_{\sigma(2)}y_{\sigma(3)}\}$
where $s_{i}(x)=x_{1}^{i}+x_{2}^{i}+x_{3}^{i}$ and $\mathrm{A}_{3}$ is the alternating group of degree 3.
3.2. The dihedral group $\mathrm{D}_{8}$
The Schur multiplier of $\mathrm{D}_{8}$ is $\mathrm{C}_{2}$, and we do not need a Schur cover (Lemma 2.5 ). As an abstract group, $\mathrm{D}_{8}$ is
$\mathrm{D}_{8}=\left<a,b\ |\ a^{4}=b^{2}=1,\ bab=a^{-1}\right>,$
and the charactet table is as follows.
$\begin{array}[]{c|rrrrr}\hline\cr\rm class&1&\rm 2A&\rm 2B&\rm 2C&\rm 4A\cr&\rm 1&a^{2}&\{b,ba^{2}\}&\{ba,ba^{3}\}&\{a,a^{3}\}\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&-1&1\cr\chi_{3}&1&1&1&-1&-1\cr\chi_{4}&1&1&-1&1&-1\cr\chi_{5}&2&-2&0&0&0\cr\hline\cr\end{array}.$
We have $\widehat{\mathrm{D}_{8}}=\{\chi_{1},\chi_{2},\chi_{3},\chi_{4}\}\cong\mathrm{C}_{2}\times\mathrm{C}_{2}$, and we do not have to consider semi-invariants by Proposition 2.6 (3). If $\rho:\mathrm{D}_{8}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic conditions, we have
$|\mathrm{Tr}\rho(g)|=\begin{cases}2\quad(\text{the class of}\ g\ \text{is}\ 2\mathrm{A},2\mathrm{B},2\mathrm{C})\\ 0\quad(\text{the class of}\ g\ \text{is}\ 4\mathrm{A})\end{cases},$
and these conditions are satisfied only by
$\theta_{1}=2\chi_{1}+\chi_{3}+\chi_{4}+\chi_{5},\qquad\theta_{2}=2\chi_{2}+\chi_{3}+\chi_{4}+\chi_{5},$
$\theta_{3}=\chi_{1}+\chi_{2}+2\chi_{3}+\chi_{5},\qquad\theta_{4}=\chi_{1}+\chi_{2}+2\chi_{4}+\chi_{5}.$
The outer automorphism group $\mathrm{Out}(\mathrm{D}_{8})=\mathrm{C}_{2}$ is generated by
$\sigma:\mathrm{D}_{8}\longrightarrow\mathrm{D}_{8},\qquad\sigma(a)=a^{3},\quad\sigma(b)=ab.$
It interchanges classes 2B and 2C, and hence $\chi_{3}$ and $\chi_{4}$. Therefore invariants for $\theta_{3}$ and $\theta_{4}$ are same. Let us construct representatoins $\rho_{i}$ affording $\theta_{i}\ (i=1,2,3)$.
We may assume that $a=\mathrm{diag}(1,1,-1,-1,i,-i)$, and then $\rho_{i}$ is given by $a$ and the following $b_{i}$:
$b_{1}$ $=\mathrm{diag}(1,1,1,-1)\oplus\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\quad b_{2}=\mathrm{diag}(-1,-1,1,-1)\oplus\begin{bmatrix}0&1\\ 1&0\end{bmatrix},$
$b_{3}$ $=\mathrm{diag}(1,-1,1,1)\oplus\begin{bmatrix}0&1\\ 1&0\end{bmatrix},$
which satisfy symplectic conditions. Now $V_{3}(\rho_{i})$ is given as a subspace of $b_{i}$-invariants in (3.5). We have
(3.7) $V_{3}(\rho_{1})=\mathrm{Span}\{x_{1}x_{3}^{2},\ x_{1}x_{4}^{2},\ x_{2}x_{3}^{2},\ x_{2}x_{4}^{2},\ F_{1}(x_{1},x_{2}),\ x_{5}x_{6}L_{1}(x_{1},x_{2}),\\ (x_{5}^{2}+x_{6}^{2})x_{3},\ (x_{5}^{2}-x_{6}^{2})x_{4}\},$
(3.8) $V_{3}(\rho_{2})=\mathrm{Span}\{x_{1}x_{3}x_{4},\ x_{2}x_{3}x_{4},\ (x_{5}^{2}-x_{6}^{2})x_{4},\ (x_{5}^{2}+x_{6}^{2})x_{3}\},$
(3.9) $V_{3}(\rho_{3})=\mathrm{Span}\{x_{1}N_{1}(x_{3},x_{4}),\ x_{1}^{3},\ x_{1}x_{2}^{2},\ x_{5}x_{6}x_{1},\ (x_{5}^{2}+x_{6}^{2})L_{2}(x_{3},x_{4})\}.$
A general member is a smooth cubic fourfold only for $V_{3}(\rho_{1})$. A cubic fourfold defined by $F\in V_{3}(\phi_{2})$ contains $\mathbb{P}^{3}=\{x_{3}=x_{4}=0\}$ and singular. Similarly, $F\in V_{3}(\phi_{3})$ vanishes on $\mathbb{P}^{3}=\{x_{1}=x_{5}+ix_{6}=0\}$. We have
$\dim\mathrm{C}_{\rho_{1}}(\mathrm{D}_{8})=2^{2}+1^{2}+1^{2}+1^{2}=7$
and
$\dim V_{3}(\rho_{1})-\dim\mathrm{C}_{\rho_{1}}(\mathrm{D}_{8})=5=20-r(\mathrm{D}_{8}).$
4. Groups of $r(G)=16$
We have the following four groups
$\mathrm{A}_{3,3},\quad\mathrm{D}_{12},\quad\mathrm{A}_{4},\quad\mathrm{D}_{10}.$
4.1. The dihedral groups $\mathrm{D}_{10}$ and $\mathrm{D}_{12}$
For $G=\mathrm{D}_{10},\ \mathrm{D}_{12}$, cubic fourfolds such that $G\subset\mathrm{Aut}^{s}(X)$ are determined in [LZ22] as follows.
(i)  $G=\mathrm{D}_{10}$. For an appropriate choice of coordinates,
(4.1) $F_{1}(x_{1},x_{2})+x_{3}x_{6}L_{1}(x_{1},x_{2})+x_{4}x_{5}L_{2}(x_{1},x_{2})\\ +x_{3}^{2}x_{5}+x_{3}x_{4}^{2}+x_{4}x_{6}^{2}+x_{5}^{2}x_{6}=0.$
An order 5 element in $G$ is $g=\mathrm{diag}(1,1,\zeta_{5},\zeta_{5}^{2},\zeta_{5}^{3},\zeta_{5}^{4})$. Moreover, any smooth cubic fourfolds with a symplectic automorphism of order 5 have this form, and a generic such cubic fourfold has symplectic automorphism group $\mathrm{D}_{10}$ which is generated by $g$ and a permutation coordinates by $(36)(45)$.
(ii)  $G=\mathrm{D}_{12}$. The defining equations of the corresponding cubic fourfolds either belong to
(4.2) $\mathrm{Span}\{x_{1}^{2}x_{3},x_{1}^{2}x_{4},x_{1}x_{2}x_{3},x_{1}x_{2}x_{4},x_{2}^{2}x_{3},x_{2}^{2}x_{4},x_{3}^{3},\\ x_{3}^{2}x_{4},x_{3}x_{4}^{2},x_{3}x_{5}x_{6},x_{4}^{3},x_{4}x_{5}x_{6},x_{5}^{3},x_{6}^{3}\},$
while an order 6 element of $G$ is $\mathrm{diag}(-1,-1,1,1,\omega,\omega^{2})$ or belong to
(4.3) $\mathrm{Span}\{x_{1}^{3},x_{1}x_{2}^{2},x_{1}x_{3}x_{5},x_{1}x_{3}x_{6},x_{2}x_{4}x_{5},x_{2}x_{4}x_{6},x_{3}^{3},x_{3}x_{4}^{2},x_{5}^{3},x_{5}^{2}x_{6},x_{5}x_{6}^{2},x_{6}^{3}\},$
while an order 6 element of $G$ is $\mathrm{diag}(1,-1,\omega,-\omega,\omega^{2},\omega^{2})$. Moreover, a generic cubic fourfold admitting such an order 6 automorphism has symplectic automorphism group $\mathrm{D}_{12}$.
The action of $\mathrm{D}_{12}$ is given by the follwoing way. A general member of (4.2) is normalized into
(4.4) $a_{1}(x_{1}^{2}x_{3}+x_{2}^{2}x_{3})+a_{2}(x_{1}^{2}x_{4}+x_{2}^{2}x_{4})+a_{3}(x_{5}^{6}+x_{6}^{6})+\text{other terms}$
by diagonal actions of (×$(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(-1,-1,1,1,\omega,\omega^{2})$ and a permutation of coordinates by $(12)(56)$.
A general member of (4.3) is normalized into
(4.5) $a_{1}(x_{1}^{3}+x_{3}^{3})+a_{2}(x_{1}x_{2}^{2}+x_{3}x_{4}^{2})+\text{other terms}$
by diagonal actions of (×$(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(1,-1,\omega,-\omega,\omega^{2},\omega^{2})$ and a permutation of coordinates by $(13)(24)$.
4.2. The groups $\mathrm{A}_{3,3}=(\mathrm{S}_{3}\times\mathrm{S}_{3})\cap\mathrm{A}_{6}$
The Schur multiplier of $G=\mathrm{A}_{3,3}$ is $\mathrm{C}_{3}$, and we consider a Schur cover $G^{*}=3.\mathrm{A}_{3,3}$ whose GAP-ID is (54,8). Firstly, we consider linear actions of $A_{3,3}$ itself. The characters teble of $A_{3,3}$ is as follows.
$\begin{array}[]{c|rrrrrr}\rm class&1&\rm 2A&\rm 3A&\rm 3B&\rm 3C&\rm 3D\cr\rm size&1&9&2&2&2&2\cr\hline\cr\chi_{1}&1&1&1&1&1&1\cr\chi_{2}&1&-1&1&1&1&1\cr\chi_{3}&2&0&2&-1&-1&-1\cr\chi_{4}&2&0&-1&2&-1&-1\cr\chi_{5}&2&0&-1&-1&2&-1\cr\chi_{6}&2&0&-1&-1&-1&2\end{array}\quad\begin{array}[]{l}\rm 2A:(12)(45)\\ \rm 3A:(123)\\ \rm 3B:(456)\\ \rm 3C:(123)(456)\\ \rm 3D:(123)(465)\\ \end{array}$
Since we have $\widehat{A_{3,3}}\cong C_{2}$, we do not have to consider semi-invariants. If $\rho:\mathrm{A}_{3,3}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ is faithful and satisfies the symplectic conditions, we have
$|\mathrm{Tr}_{\rho}(g)|=\begin{cases}2\quad(\text{the order of $g$ is 2})\\ 0,\,3\quad(\text{the order of $g$ is 3})\end{cases},$
and these conditions are satisfied only by
$\theta_{ij}=2\chi_{1}+\chi_{i}+\chi_{j},\quad\theta_{ij}\otimes\chi_{2}\qquad(3\leq i<j\leq 6).$
The outer automorphism group $\mathrm{Out}(A_{3,3})$ is isomorphic to $\mathrm{S}_{4}$, and it permutes classes 3A, 3B, $\rm 3A,\ 3B,\ 3C$ and $\rm 3D$. We consider only $\theta_{34}$ and $\theta_{34}\otimes\chi_{2}$ since $\mathrm{Out}(\mathrm{A}_{3,3})$ permutes $\theta_{ij}$ (see Proposition 2.6 (1)).
Remark 4.1.
We see that $\mathrm{Out}(A_{3,3})$ is generated by conjugations by $(12),(13)(24)(36)$, and an automorphism $\sigma$ in Appendix.
A representation $\rho$ affording $\theta_{34}$ is given by permutations of corrdinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is
(4.6) $\mathrm{Span}\{s_{1}(x)^{3},\ s_{2}(x)s_{1}(x),\ s_{3}(x),\ s_{1}(y)^{3},\ s_{2}(y)s_{1}(y),\ s_{3}(y),\\ s_{1}(x)s_{2}(y),\ s_{1}(y)s_{2}(x),\ s_{1}(x)s_{1}(y)^{2},\ s_{1}(x)^{2}s_{1}(y)\},$
which is a subspace of (3.6). Although these are stable under permutatoins by $\mathrm{S}_{3,3}$, the actoin of $(12)\in\mathrm{S}_{3,3}$ is non-symplectic. We have
$\dim\mathrm{C}_{\rho}(\mathrm{A}_{3,3})=2^{2}+1^{2}+1^{2}=6$
and
$\dim V_{3}(\rho)-\dim\mathrm{C}_{\rho}(\mathrm{A}_{3,3})=4=20-r(\mathrm{A}_{3,3}).$
By Proposition 2.6 (2), we have $V_{3}(\rho\otimes\chi_{2})=V_{3}(\rho,\chi_{2})$ and it is
(4.7) $\mathrm{Span}\{(x_{1}-x_{2})(x_{1}-x_{3})(x_{2}-x_{3}),\ (y_{1}-y_{2})(y_{1}-y_{3})(y_{2}-y_{3})\}.$
This is not the case.
Next, we consider $G^{*}=3.A_{3,3}$. Since we have $\widehat{G^{*}}\cong\widehat{G}\cong\mathrm{C}_{2}$, we do not have to consider semi-invariants. Irreducible representaions of $G^{*}$ with a non-trivial action of the Schur multiplier $\mathrm{C}_{3}$ are given as follows.
$\begin{array}[]{c|rrrrrrrrrr}\hline\cr\rm class&\rm 1&\rm 2A&\rm 3A&\rm 3B&\rm 3C&\rm 3D&\rm 3E&\rm 3F&\rm 6A&\rm 6B\cr\rm size&1&9&1&1&6&6&6&6&9&9\cr\hline\cr\chi_{7}&3&1&3\overline{\omega}&3\omega&0&0&0&0&\omega&\overline{\omega}\cr\chi_{8}&3&1&3\omega&3\overline{\omega}&0&0&0&0&\overline{\omega}&\omega\cr\chi_{9}&3&-1&3\omega&3\overline{\omega}&0&0&0&0&-\overline{\omega}&-\omega\cr\chi_{10}&3&-1&3\overline{\omega}&3\omega&0&0&0&0&-\omega&-\overline{\omega}\cr\hline\cr\end{array}$
Classes 1, 3A and 3B form the Schur multiplier, and they act as scalar matrices for
$2\chi_{i}\ (i=7,8,9,10),\quad\chi_{7}+\chi_{10},\quad\chi_{8}+\chi_{9}.$
Since we have
$(\chi_{7}+\chi_{10})(2\mathrm{A})=(\chi_{8}+\chi_{9})(2\mathrm{A})=0,$
$\chi_{7}+\chi_{10}$ and $\chi_{8}+\chi_{9}$ are not suitable. Let $\rho_{i}$ be a representation affording $\chi_{i}$. We have Molien serieses
$\mu(2\rho_{i},\,t)=\sum_{d=0}^{\infty}(\dim V_{d}(2\rho_{i}))t^{d}=1+8t^{3}+\cdots\qquad(i=7,8),$
$\mu(2\rho_{i},\,t)=\sum_{d=0}^{\infty}(\dim V_{d}(2\rho_{i}))t^{d}=1+42t^{6}+\cdots\qquad(i=9,10),$
and there is no invariant cubic for $2\rho_{9}$ and $2\rho_{10}$. Moreover, we have $\chi_{8}=\overline{\chi_{7}}$, and hence $\rho_{8}$ is the complex conjugate of $\rho_{7}$.
Let us consider matrices
(4.8) $a=\begin{bmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{bmatrix},\quad b=\begin{bmatrix}0&\omega&0\\ \omega^{2}&0&0\\ 0&0&1\end{bmatrix},\quad c=\begin{bmatrix}1&0&0\\ 0&\omega&0\\ 0&0&\omega^{2}\end{bmatrix}$
with eigenvalues $(1,1,-1),\ (1,1,-1),\ (1,\omega,\omega^{2})$. These matrices generate a group $H$ with GAP-ID (54,8), and this is one of $\rho_{7}$ and $\rho_{8}$. Note that we have
(4.9) $\overline{a}=a,\quad\overline{b}=c^{2}bc,\quad\overline{c}=aca$
and hence $\overline{H}=H$, that is, $V_{3}(2\rho_{7})=V_{3}(2\rho_{8})$. From eigenvalues, we see that matrices
(4.10) $\tilde{a}=\begin{bmatrix}a&{\bf 0}\\ {\bf 0}&a\end{bmatrix},\quad\tilde{b}=\begin{bmatrix}b&{\bf 0}\\ {\bf 0}&b\end{bmatrix},\quad\tilde{c}=\begin{bmatrix}c&{\bf 0}\\ {\bf 0}&c\end{bmatrix},\quad$
satisfy symplectic conditions, and $V_{3}(2\rho_{7})$ with projective coordinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ is given by
(4.11) ${\rm Span}\{s_{3}(x),\ x_{1}x_{2}x_{3},\ s_{3}(y),\ y_{1}y_{2}y_{3},\ I_{1},\ I_{2},\ I_{3},\ I_{4}\}$
where
$I_{1}$ $=x_{1}y_{1}^{2}+x_{2}y_{2}^{2}+x_{3}y_{3}^{2},\qquad I_{2}=x_{1}y_{2}y_{3}+y_{1}x_{2}y_{3}+y_{1}y_{2}x_{3}$
$I_{3}$ $=x_{1}^{2}y_{1}+x_{2}^{2}y_{2}+x_{3}^{2}y_{3},\qquad I_{4}=y_{1}x_{2}x_{3}+x_{1}y_{2}x_{3}+x_{1}x_{2}y_{3}.$
We have $\dim\mathrm{C}_{\rho_{7}}(3.\mathrm{A}_{3,3})=2^{2}=4$ and
$\dim V_{3}(\rho_{7})-\dim\mathrm{C}_{\rho_{7}}(3.\mathrm{A}_{3,3})=4=20-r(\mathrm{A}_{3,3}).$
Remark 4.2.
We can also consider $3.A_{3,3}$ as a subgroup of $H\cong 3.\mathrm{A}_{6}$ in section 2.5. Here we chose the above matrices (which is not in $H$) to obtain invariants of simple forms.
4.3. The alternating group $\mathrm{A}_{4}$
The Schur multiplier of $\mathrm{A}_{4}$ is $\mathrm{C}_{2}$, and we do not need a Schur cover. The character table of $\mathrm{A}_{4}$ is as follows:
(4.18) $\begin{array}[]{c|cccc}\rm class&1&(12)(34)&(123)&(124)\cr\rm size&1&3&4&4\cr\hline\cr\chi_{1}&1&1&1&1\cr\chi_{2}&1&1&\omega&\overline{\omega}\cr\chi_{3}&1&1&\overline{\omega}&\omega\cr\chi_{4}&3&-1&0&0\cr\end{array}$
We see that only the following characters satisfy the condition on absolute values:
(4.19) $\theta_{i}=3\chi_{i}+\chi_{4}\ (i=1,2,3),\quad\theta_{4}=2\chi_{4},\quad\theta_{5}=\chi_{1}+\chi_{2}+\chi_{3}+\chi_{4}$
(Note that we have $\theta_{i}=\theta_{1}\otimes\chi_{i}$ for $i=1,2,3$). Third symmetric powers of these characters are
$\mathrm{Sym}^{3}\theta_{i}$ $=14\chi_{1}+3\chi_{2}+3\chi_{3}+12\chi_{4}\ (i=1,2,3),$
$\mathrm{Sym}^{3}\theta_{4}$ $=4\chi_{1}+2\chi_{2}+2\chi_{3}+16\chi_{4},$
$\mathrm{Sym}^{3}\theta_{5}$ $=8\chi_{1}+6\chi_{2}+6\chi_{3}+12\chi_{4}.$
Let $\rho_{i}$ be a representation affording $\theta_{i}$. Then we have
$\dim V_{3}(\rho_{i})=14\ (i=1,2,3),\quad\dim V_{3}(\rho_{5})=8,\qquad\dim V_{3}(\rho_{5},\chi_{i})=6\ (i=2,3)$
and we can expect smooth cubics with symplectic action of $\mathrm{A}_{4}$ only for these spaces by Lemma 2.7. By Proposition 2.6 (2), we have
$V_{3}(\rho_{1})=V_{3}(\rho_{1}\otimes\chi_{i},\,\chi_{i}^{3})=V_{3}(\rho_{1}\otimes\chi_{i})=V_{3}(\rho_{i})$
since $\chi_{i}^{3}=1$ for $i=1,2,3$.
Taking projective coordinates $[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by permutations of $x_{i}$, and it satisfies the symplectic condition. The linear space $V_{3}(\rho_{1})$ is
(4.20) $\mathrm{Span}\{s_{1}(x)s_{2}(x),\ s_{2}(x)y_{1},\ s_{2}(x)y_{2},\ s_{3}(x),\ \text{cubic forms in}\ s_{1}(x),y_{1},y_{2}\}$
where $s_{i}(x)=x_{1}^{i}+x_{2}^{i}+x_{3}^{i}+x_{4}^{i}$. A general member of $V_{3}(\rho_{1})$ is smooth since it contains the Fermat cubic. We have $\dim\rho_{1}(\mathrm{A}_{4})=3^{2}+1^{2}=10$ and
$\dim V_{3}(\phi_{1})-\dim\mathrm{C}_{\rho_{1}}(\mathrm{A}_{4})=4=20-r(\mathrm{A}_{4}).$
The representation $\rho_{5}$ is given by
(4.21) $\begin{split}(123)[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]=[x_{2}:x_{3}:x_{1}:x_{4}:\omega y_{1}:\omega^{2}y_{2}],\\ (12)(34)[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]=[x_{2}:x_{1}:x_{4}:x_{3}:y_{1}:y_{2}],\end{split}$
and it satisfies symplectic conditions. The linear space $V_{3}(\rho_{5})$ is spanned by
(4.22) $s_{1}(x)^{3},\quad s_{1}(x)s_{2}(x),\quad s_{3}(x),\quad s_{1}(x)y_{1}y_{2},\quad y_{1}^{3},\quad y_{2}^{3},\quad f(x)y_{2},\quad g(x)y_{1},$
where
$f(x)$ $=(x_{1}-x_{2})(x_{3}-x_{4})-\omega(x_{1}-x_{3})(x_{2}-x_{4})+\omega^{2}(x_{1}-x_{4})(x_{2}-x_{3}),$
$g(x)$ $=(x_{1}-x_{2})(x_{3}-x_{4})-\omega^{2}(x_{1}-x_{3})(x_{2}-x_{4})+\omega(x_{1}-x_{4})(x_{2}-x_{3}).$
A general menber of $V_{3}(\rho_{5})$ is smooth since it contains the Fermat cubic. We have
$\dim\mathrm{C}_{\rho_{5}}(\mathrm{A}_{4})=1^{2}+1^{2}+1^{2}+1^{2}=4$
and
$\dim V_{3}(\rho_{5})-\dim\mathrm{C}_{\rho_{5}}(\mathrm{A}_{4})=4=20-r(\mathrm{A}_{4}).$
Since we have
$\dim V_{3}(\rho_{5},\ \chi_{i})-\dim\mathrm{C}_{\rho_{5}}(\mathrm{A}_{4})=6-4=2\quad(i=2,3),$
a general menber of $V_{3}(\rho_{5},\chi_{i})$ is not smooth for $i=2,3$.
5. Groups of $r(G)=17$
We have two groups $\mathrm{S}_{4}$ and $\mathrm{Q}_{8}$.
5.1. The quaternion group $\mathrm{Q}_{8}$
The Schur multiplier of $\mathrm{Q}_{8}$ is 1, and hence we have $\mathrm{Q}_{8}^{*}=\mathrm{Q}_{8}$. We express $\mathrm{Q}_{8}$ as
$\mathrm{Q}_{8}=\{\pm 1,\ \pm I,\ \pm J,\ \pm K\}$
with usual rules $I^{2}=J^{2}=K^{2}=-1$,  $IJ=-JI=K$. The character table of $\mathrm{Q}_{8}$ is the following:
$\begin{array}[]{c|rrrrr}\rm class&\rm 1&-1&\pm I&\pm J&\pm K\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&1&-1\cr\chi_{3}&1&1&1&-1&-1\cr\chi_{4}&1&1&-1&-1&1\cr\chi_{5}&2&-2&0&0&0\cr\end{array}.$
Since we have $\widehat{\mathrm{Q}_{8}}=\mathrm{C}_{2}\times\mathrm{C}_{2}$, we do not have to consider smi-invariants. If $\rho:\mathrm{Q}_{8}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic conditions, we have
$|\mathrm{Tr}_{\rho}(g)|=\begin{cases}2\quad(g=-1)\\ 0\quad(g=I,J,K)\end{cases},$
and these conditions are satisfied only by
$\theta=\chi_{1}+\chi_{2}+\chi_{3}+\chi_{4}+\chi_{5}.$
A representation $\rho$ affording $\theta$ is given by
$\rho(I)=\mathrm{diag}(1,-1,1,-1)\oplus\begin{bmatrix}0&1\\ -1&0\end{bmatrix},\quad\rho(J)=\mathrm{diag}(1,1,-1,-1)\oplus\begin{bmatrix}i&0\\ 0&-i\end{bmatrix}.$
Note that $\rho(J)$ is the same with the action of $\mathrm{C}_{4}$ for (3.5). The linear space $V_{3}(\rho)$ is given as $\rho(I)$-invariants in (3.5), and spanned by
(5.1) $x_{1}^{3},\ x_{1}x_{2}^{2},\ x_{1}x_{3}^{2},\ x_{1}x_{4}^{2},\ x_{2}x_{3}x_{4},\ x_{2}x_{5}x_{6},\ x_{3}(x_{5}^{2}+x_{6}^{2}),\ x_{4}(x_{5}^{2}-x_{6}^{2}).$
We have
$\dim\mathrm{C}_{\rho}(\mathrm{Q}_{8})=1^{2}+1^{2}+1^{2}+1^{2}+1^{2}=5$
and
$\dim V_{3}(\rho)-\dim\mathrm{C}_{\rho}(\mathrm{Q}_{8})=3=20-r(\mathrm{Q}_{8}).$
5.2. The symmetric group $\mathrm{S}_{4}$
The Schur multiplier of $\mathrm{S}_{4}$ is $\mathrm{C}_{2}$, and we do not have to consider a Schur cover. The character tables of $\mathrm{S}_{4}$ is as follows:
$\begin{array}[]{c|rrrrr}\rm class&1&(12)(34)&(12)&(123)&(1234)\cr\rm size&1&3&6&8&6\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&1&-1\cr\chi_{3}&2&2&0&-1&0\cr\chi_{4}&3&-1&-1&0&1\cr\chi_{5}&3&-1&1&0&-1\cr\end{array}$
Since we have $\widehat{\mathrm{S}_{4}}=\mathrm{C}_{2}$, we do not have to consider smi-invariants. If $\rho:\mathrm{S}_{4}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic condition, we have
$|\mathrm{Tr}_{\rho}(12)(34)|=|\mathrm{Tr}_{\rho}(12)|=2,\quad|\mathrm{Tr}_{\rho}(123)|=0,3,\quad|\mathrm{Tr}_{\rho}(1234)|=0.$
These conditions are satisfied only for
$\theta_{1}=2\chi_{1}+\chi_{2}+\chi_{5},\quad\theta_{2}=\theta_{1}\otimes\chi_{2},\quad\theta_{3}=\chi_{1}+\chi_{3}+\chi_{5},\quad\theta_{4}=\theta_{3}\otimes\chi_{2},$
and their third symmetric powers are
$\mathrm{Sym}^{3}\theta_{1}=9\chi_{1}+5\chi_{2}+3\chi_{3}+4\chi_{4}+8\chi_{5},\quad\mathrm{Sym}^{3}\theta_{2}=\chi_{2}\otimes\mathrm{Sym}^{3}\theta_{1},$
$\mathrm{Sym}^{3}\theta_{3}=6\chi_{1}+2\chi_{2}+6\chi_{3}+4\chi_{4}+8\chi_{5},\quad\mathrm{Sym}^{3}\theta_{4}=\chi_{2}\otimes\mathrm{Sym}^{3}\theta_{3}.$
Let $\rho_{i}:\mathrm{S}_{4}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ be a representation affording $\theta_{i}$. We have
$\dim V_{3}(\rho_{1})=9,\qquad\dim V_{3}(\rho_{2})=\dim V_{3}(\rho_{1},\chi_{2})=5,$
$\dim V_{3}(\rho_{3})=6,\qquad\dim V_{3}(\rho_{4})=\dim V_{3}(\rho_{3},\chi_{2})=2,$
and $\rho_{4}$ is not suitable. Taking projective coordinates $[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by
$(1234)[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]=[x_{2}:x_{3}:x_{4}:x_{1}:y_{1}:-y_{2}],$
$(12)[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]=[x_{2}:x_{1}:x_{3}:x_{4}:y_{1}:-y_{2}],$
and it satisfies symplectic conditions. The linear space $V_{3}(\rho_{1})$ is a $\rho_{1}((12))$-invariant subspace of (4.20):
(5.2) $\mathrm{Span}\{s_{1}(x)^{3},\ s_{1}(x)^{2}y_{1},\ s_{1}(x)y_{1}^{2},\ s_{1}(x)y_{2}^{2},\ y_{1}^{3},\ y_{1}y_{2}^{2},\\ s_{1}(x)s_{2}(x),\ s_{2}(x)y_{1},\ s_{3}(x)\}.$
A general member of $V_{3}(\rho_{1})$ is smooth since it contains a smooth cubic
$s_{3}(x)+y_{1}^{2}y_{2}+y_{2}^{3}=0.$
We have $\dim\mathrm{C}_{\rho_{1}}(\mathrm{S}_{4})=2^{2}+1^{2}+1^{2}=6$ and
$\dim V_{3}(\rho_{1})-\dim\mathrm{C}_{\rho_{1}}(\mathrm{S}_{4})=3=20-r(\mathrm{S}_{4}).$
Since we have
$\mathrm{C}_{\rho_{2}}(\mathrm{S}_{4})=\mathrm{C}_{\rho_{1}\otimes\chi_{1}}(\mathrm{S}_{4})=\mathrm{C}_{\rho_{1}}(\mathrm{S}_{4})$
and $\dim V_{3}(\rho_{2})=5$, the linear space $V_{3}(\rho_{2})$ does not contain a smooth cubic.
The representation $\rho_{3}$ is given by
$(1234)[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]=[x_{2}:x_{3}:x_{4}:x_{1}:-\omega y_{2}:-\omega^{2}y_{1}],$
$(12)[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]=[x_{2}:x_{1}:x_{3}:x_{4}:-\omega^{2}y_{2}:-\omega y_{1}].$
and it satisfies the symplectic conditions. Moreover the restriction of this action to $\mathrm{A}_{4}$ coincides with the action (4.21). The linear space $V_{3}(\rho_{3})$ is a $\rho_{3}((12))$-invariant subspace of (4.22)
(5.3) ${\rm Span}\{s_{1}(x)^{3},\quad s_{1}(x)s_{2}(x),\quad s_{3}(x),\quad s_{1}(x)y_{1}y_{2},\\ y_{1}^{3}+y_{2}^{3},\quad\omega^{2}f(x)y_{2}+\omega g(x)y_{1}\}.$
A general member of $V_{3}(\rho_{3})$ is smooth since it contains a smooth cubic
$s_{3}(x)+y_{1}^{2}y_{2}+y_{2}^{3}=0.$
We have $\dim\mathrm{C}_{\rho_{3}}(\mathrm{S}_{4})=1^{2}+1^{2}+1^{2}=3$ and
$\dim V_{3}(\rho_{3})-\dim\mathrm{C}_{\rho_{3}}(\mathrm{S}_{4})=3=20-r(\mathrm{S}_{4}).$
6. Groups of $r(G)=18$
We have eight groups:
$3^{1+4}_{+}:2,\quad\mathrm{A}_{4,3},\quad\mathrm{A}_{5},\quad 3^{2}.4,\quad\mathrm{S}_{3,3},\quad\mathrm{F}_{21},\quad\mathrm{Hol}(5),\quad\mathrm{QD}_{16}.$
6.1. The group $\mathbf{3^{1+4}_{+}:2,\ F_{21}}$ and $\mathbf{QD_{16}}$
These groups are studied in [LZ22].
(i)  $G\cong 3^{1+4}_{+}:2$. For an appropriate choice of coordinates, the defining equations of the corresponding cubic fourfolds are given by
(6.1) $F_{1}(x_{1},x_{2},x_{3})+F_{2}(x_{4},x_{5},x_{6})=0$
An element of order 3 in $G$ is $\mathrm{diag}(1,1,1,\omega,\omega,\omega)$. Moreover, any smooth cubic fourfold with a symplectic automorphism which can be diagonalized as $\mathrm{diag}(1,1,1,\omega,\omega,\omega)$ has this form, and a generic such cubic fourfold has symplectic automorphism group $23^{1+4}_{+}:$ 2.
Remark 6.1.
Extraspecial groups $G=p^{1+2n}_{\pm}$ are non-abelian $p$-groups with center $Z(G)\cong\mathrm{C}_{p}$ and $G/Z(G)\cong\mathrm{C}_{p}^{2n}$. The exponent of $p^{1+2n}_{+}$ is $1$, and that of $p^{1+2n}_{-}$ is $2$. We give an explicit matrix representation of $23^{1+4}:$ 2 in section 7.1.
(ii)  $G\cong\mathrm{F}_{21}$. For an appropriate choice of coordinates, the defining equations of the corresponding cubic fourfolds are given by
(6.2) $x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+x_{3}^{2}x_{4}+x_{4}^{2}x_{5}+x_{5}^{2}x_{6}+x_{6}^{2}x_{1}+ax_{1}x_{3}x_{5}+bx_{2}x_{4}x_{6}=0.$
The automorphisms $g_{1}=\mathrm{diag}(\zeta_{7},\zeta_{7}^{5},\zeta_{7}^{4},\zeta_{7}^{6},\zeta_{7}^{2},\zeta_{7}^{3})$ and $g_{2}:x_{i}\rightarrow x_{i+2}$ generate $\mathrm{F}_{21}\cong 7.3$. Moreover, any smooth cubic fourfold with symplectic automorphism of order 7 has this form, and a generic such cubic fourfold has symplectic automorphism group $\mathrm{F}_{21}$.
(iii)  $G\cong\mathrm{QD}_{16}$. For an appropriate choice of coordinates, the defining equations of the corresponding cubic fourfolds belong to
(6.3) $\mathrm{Span}\{x_{1}^{3},\ x_{1}x_{2}^{2},\ x_{2}x_{3}^{2},\ x_{2}x_{4}^{2},\ x_{1}x_{3}x_{4},\ x_{4}x_{5}^{2},\ x_{3}x_{6}^{2},\ x_{2}x_{5}x_{6}\}$
An element of order 8 in $G$ is $a=\mathrm{diag}(1,-1,i,-i,\zeta_{8},\zeta_{8}^{3})$. Moreover, any smooth cubic fourfold with a symplectic automorphism of order 8 has this form, and a generic such cubic fourfold has symplectic automorphism group $\mathrm{QD}_{16}$.
Remark 6.2.
The quasi-dihedral group $\mathrm{QD_{16}}$ is defined by
$\mathrm{QD_{16}}=\left<a,b\ |\ a^{8}=b^{2}=1,\ bab=a^{3}\right>.$
A general menber in (6.3) is normalized into
(6.4) $a_{1}x_{1}^{3}+a_{2}x_{1}^{2}x_{3}+a_{3}(x_{2}x_{3}^{2}+x_{2}x_{4}^{2})+a_{3}x_{1}x_{3}x_{4}\\ +a_{4}(x_{4}x_{5}^{2}+x_{3}x_{6}^{2})+a_{5}x_{2}x_{5}x_{6}$
by diagonal actions of (×$(\mathbb{C}^{\times})^{6}$. Then $\mathrm{QD}_{16}$ is generated by $a=\mathrm{diag}(1,-1,i,-i,\zeta_{8},\zeta_{8}^{3})$ and a permutation of coordinates by $(34)(56)$.
6.2. The group $\mathrm{A}_{4,3}=(\mathrm{S}_{4}\times\mathrm{S}_{3})\cap\mathrm{A}_{7}$
The Schur multiplier of $\mathrm{A}_{4,3}$ is $\mathrm{C}_{6}$, and we must consider a covering group $3.\mathrm{A}_{4,3}$. Firstly, we consider linear actions of $\mathrm{A}_{4,3}$ itself. Tthe character table of $\mathrm{A}_{4,3}$ is as follows:
$\begin{array}[]{c|rrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 3A&\rm 3B&\rm 3C&\rm 3D&\rm 4A&\rm 6A\cr\rm size&1&3&18&2&8&8&8&18&6\cr\hline\cr\chi_{1}&1&1&1&1&1&1&1&1&1\cr\chi_{2}&1&1&-1&1&1&1&1&-1&1\cr\chi_{3}&2&2&0&2&-1&-1&-1&0&2\cr\chi_{4}&2&2&0&-1&2&-1&-1&0&-1\cr\chi_{5}&2&2&0&-1&-1&-1&2&0&-1\cr\chi_{6}&2&2&0&-1&-1&2&-1&0&-1\cr\chi_{7}&3&-1&-1&3&0&0&0&1&-1\cr\chi_{8}&3&-1&1&3&0&0&0&-1&-1\cr\chi_{9}&6&-2&0&-3&0&0&0&0&1\cr\end{array}\quad\begin{array}[]{l}\rm 2A:(12)(34)\\ \rm 2B:(12(56)\\ \rm 3A:(567)\\ \rm 3B:(123)\\ \rm 3C:(123)(567)\\ \rm 3D:(123)(576)\\ \rm 4A:(1234)(67)\\ \rm 6A:(12)(34)(567)\end{array}$
Since we have $\widehat{\mathrm{A}_{4,3}}=\mathrm{C}_{2}$, we do not have to consider smi-invariants. A character $\phi$ satisfying conditions
$|\phi(2A)|=|\phi(2B)|=2,\qquad|\phi(4A)|=0$
is one of the followings:
$\phi_{i}=\chi_{1}+\chi_{i}+\chi_{8},\quad\phi_{i}\otimes\chi_{2}=\chi_{2}+\chi_{i}+\chi_{7}\quad(i=4,5,6).$
We have $\mathrm{Out}(\mathrm{A}_{4,3})\cong\mathrm{S}_{3}$ which permutes classes 3B, 3C and 3D, and hence $\phi_{4},\ \phi_{5}$ and $\phi_{6}$. Hence we consider only $\phi_{4}$ and $\phi_{4}\otimes\chi_{2}$.
Remark 6.3.
The conjugation by $(67)\in\mathrm{S}_{7}$ interchages classes 3C and 3D. However, the cycle type of 3B is different with that of 3B and 3C, and we need another kind of automorphisms to generate $\mathrm{Out}(\mathrm{A}_{4,3})$. One of them is given by
$[(3,4)(6,7),\ (5,6,7),\ (2,3,4),\ (1,4)(2,3),\ (1,3)(2,4)]\\ \longrightarrow[(2,4)(5,6),\ (5,6,7),\ (2,4,3)(5,6,7),\ (1,4)(2,3),\ (1,2)(3,4)],$
which is obtained by GAP.
Let $\rho$ be a representation affording $\phi_{4}$. Since Molien series of $\rho$ and $\rho\otimes\chi_{2}$ are
$\mu(\rho,t)=1+t+3t^{2}+5t^{3}+\cdots,\qquad\mu(\rho\otimes\chi_{2},t)=1+3t^{2}+t^{3}+\cdots,$
we have $\dim V_{3}(\rho\otimes\chi_{2})=1$ and this is not the case. The representaion $\rho$ is given as follows. Let $[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}:y_{3}]$ be projective coordinates of $\mathbb{P}^{6}$, and $H_{0}$ be a hyperplane
$s_{1}(x)+s_{1}(y)=(x_{1}+x_{2}+x_{3}+x_{4})+(y_{1}+y_{2}+y_{3})=0$
Now $\mathrm{A}_{4,3}$ acts on $H_{0}\cong\mathbb{P}^{5}$ as permutations of $x_{i}$ and $y_{i}$ respectively, and we have the following decomposition
$\chi_{1}\text{-space}\ [z:z:z:z:-z:-z:-z],$
$\chi_{8}\text{-space}\ [x_{1}:x_{2}:x_{3}:x_{4}:0:0:0]\ \text{with}\ s_{1}(x)=0,$
$\chi_{4}\text{-space}\ [0:0:0:0:y_{1}:y_{2}:y_{3}]\ \text{with}\ s_{1}(y)=0.$
The linear space $V_{3}(\rho)$ in the quotient ring $\mathbb{C}[x_{i},y_{i}]/(s_{1}(x)+s_{1}(y))$ is given by
(6.5) $\mathrm{Span}\{s_{1}(x)^{3},\ s_{1}(x)s_{2}(x),\ s_{1}(x)s_{2}(y),\ s_{3}(x),\ s_{3}(y)\}.$
Note that this family contains the Clebsch-Segre cubic (2.1), and a general member is smooth. We have $\dim\mathrm{C}_{\rho}(\mathrm{A}_{4,3})=1^{2}+1^{2}+1^{2}=3$ and
$\dim V_{3}(\rho)-\dim\mathrm{C}_{\rho}(\mathrm{A}_{4,3})=2=20-r(\mathrm{A}_{4,3}).$
Next we consider a triple cover $3.A_{4,3}=(\mathrm{A}_{4,3})^{*}/\mathrm{C}_{2}$ whose GAP-ID is (216,95). Since we have $\widehat{G^{*}}\cong\widehat{G}\cong\mathrm{C}_{2}$, we do not have to consider semi-invariants. Irreducible representaions of $G^{*}$ with a non-trivial action of the Schur multiplier are as follows.
$\begin{array}[]{l|rrrrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 3A&\rm 3B&\rm 3C&\rm 3D&\rm 3E&\rm 3F&\rm 4&\cdots\cr\rm size&1&3&18&1&1&6&24&24&24&18&\cdots\cr\hline\cr\chi_{10}&3&3&1&3\omega&3\overline{\omega}&0&0&0&0&1\cr\chi_{11}&3&-1&1&3\omega&3\overline{\omega}&0&0&0&0&-1\cr\chi_{12}=\chi_{10}\otimes\chi_{2}&3&3&-1&3\omega&3\overline{\omega}&0&0&0&0&-1\cr\chi_{13}=\chi_{11}\otimes\chi_{2}&3&-1&-1&3\omega&3\overline{\omega}&0&0&0&0&1\cr\chi_{14}=\overline{\chi_{10}}&3&3&1&3\overline{\omega}&3\omega&0&0&0&0&1\cr\chi_{15}=\overline{\chi_{11}}&3&-1&1&3\overline{\omega}&3\omega&0&0&0&0&-1\cr\chi_{16}=\overline{\chi_{10}}\otimes\chi_{2}&3&3&-1&3\overline{\omega}&3\omega&0&0&0&0&-1\cr\chi_{17}=\overline{\chi_{11}}\otimes\chi_{2}&3&-1&-1&3\overline{\omega}&3\omega&0&0&0&0&1\cr\chi_{18}&6&-2&0&6\omega&6\overline{\omega}&0&0&0&0&0\cr\chi_{19}&6&-2&0&6\overline{\omega}&6\omega&0&0&0&0&0\cr\end{array}$
(Conjugacy classes for order $6$ and $12$ are omitted.)
Classes 1, 3A and 3B form the Schur multiplier, and they act as scalar matrices for
$\chi_{i}+\chi_{j}\ (10\leq i,j\leq 13),\quad\chi_{k}+\chi_{l}\ (14\leq i,j\leq 17),\quad\chi_{18},\quad\chi_{19}.$
Among these characters, conditions
$|\phi(2A)|=|\phi(2B)|=2,\qquad|\phi(4A)|=0$
are satisfied only by
$\phi_{1}=\chi_{10}+\chi_{11},\quad\overline{\phi_{1}},\quad\phi_{2}=\chi_{12}+\chi_{13},\quad\overline{\phi_{2}}.$
Let $\rho_{i}$ be a representation affording $\phi_{i}$. We have Molien series
$\mu(\rho_{1},t)=1+4t^{3}+\cdots,\qquad\mu(\rho_{2},t)=1+17t^{6}+\cdots$
and
$\dim V_{3}(\rho_{1})=4,\qquad\dim V_{3}(\rho_{2})=0.$
For $a$ and $b$ in (2.2), the following elements
(6.6) $u_{1}=aba^{-1}bab^{-2}a^{-1}b,\quad u_{2}=b^{2}ab^{2}a^{-1}bab,\quad u_{3}=b^{2}aba^{-1}b^{-1}ab^{2}$
generate a group with GAP-ID (216, 95), and this gives one of $\rho_{1}$ and $\overline{\rho_{1}}$. Moreover we have
$\overline{u_{1}}=u_{1},\quad\overline{u_{2}}=u_{3}^{-1}(u_{2}u_{1})^{2},\quad\overline{u_{3}}=u_{3}u_{2}^{2}u_{3}u_{2},$
and hence $V_{3}(\rho_{1})=V_{3}(\overline{\rho_{1}})$. This space is given by
(6.7) $\mathrm{Span}\{p_{1}(x),\ p_{2}(x),\ p_{3}(x),\ p_{4}(x)\}$
where
$p_{1}(x)$ $=x_{2}^{3}+x_{4}^{3}-3x_{246}+x_{6}^{3}$
$p_{2}(x)$ $=8x_{135}+4x_{235}+4x_{145}+2x_{245}+4x_{136}+2x_{236}+2x_{146}+x_{246}$
$p_{3}(x)$ $=2x_{122}+2x_{344}+4x_{245}+4x_{236}+4x_{146}+9x_{246}+2x_{566}$
$p_{4}(x)$ $=2x_{1}^{3}+3x_{112}+2x_{3}^{3}+3x_{334}-3x_{245}+2x_{5}^{3}-3x_{236}-3x_{146}-6x_{246}+3x_{556}$
with $x_{ijk}=x_{i}x_{j}x_{k}$. This family contains a smooth cubic fourfold (2.3), and a general member is smooth. We have $\dim\mathrm{C}_{\rho_{1}}(3.\mathrm{A}_{4,3})=1^{2}+1^{2}=2$ and
$\dim V_{3}(\rho_{1})-\dim\mathrm{C}_{\rho_{1}}(3.\mathrm{A}_{4,3})=2=20-r(\mathrm{A}_{5}).$
6.3. The alternating group $\mathbf{\mathrm{A}_{5}}$
The Schur multiplier of $\mathrm{A}_{5}$ is $\mathrm{C}_{2}$, and we do not have to consider a Schur cover. The character tables of $\mathrm{A}_{5}$ is as follows:
$\begin{array}[]{c|rrrrr}\rm class&\rm 1&(12)(34)&(123)&(12345)&(13524)\cr\rm size&1&15&20&12&12\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&3&-1&0&\frac{1+\sqrt{5}}{2}&\frac{1-\sqrt{5}}{2}\cr\chi_{3}&3&-1&0&\frac{1-\sqrt{5}}{2}&\frac{1+\sqrt{5}}{2}\cr\chi_{4}&4&0&1&-1&-1\cr\chi_{5}&5&1&-1&0&0\cr\end{array}$
Since we have $\widehat{\mathrm{A}_{5}}=1$, we do not have to consider smi-invariants. If $\rho:\mathrm{A}_{5}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic condition, we have
$|\mathrm{Tr}_{\rho}(12)(34)|=2,\quad|\mathrm{Tr}_{\rho}(123)|=0,3,\quad|\mathrm{Tr}_{\rho}(12345)|=1.$
These conditions are satisfied only for
(6.8) $\theta_{1}=2\chi_{1}+\chi_{4},\quad\theta_{2}=\chi_{1}+\chi_{5},\quad\theta_{3}=\chi_{2}+\chi_{3}.$
For representations $\rho_{i}$ affording $\theta_{i}$, we have Molien series
$\mu(\rho_{1},t)$ $=1+2t+4t^{2}+7t^{3}+\cdots,$
$\mu(\rho_{2},t)$ $=1+t+2t^{2}+4t^{3}+\cdots,$
$\mu(\rho_{3},t)$ $=1+2t^{2}+6t^{4}+\cdots,$
and $V_{3}(\rho_{3})=0$. For projective coordinates $[x_{1}:\cdots:x_{5}:y]$, the representation $\rho_{1}$ is given as permutations of $x_{i}$, and we have
(6.9) $V_{3}(\rho_{1})=\mathrm{Span}\{y^{3},\ y^{2}s_{1}(x),\ ys_{1}(x)^{2},\ ys_{2}(x),\ s_{1}(s)^{3},\ s_{1}(x)s_{2}(x),\ s_{3}(x)\}$
Although $F\in V_{3}(\rho_{1})$ is stable under permutatoins by $\mathrm{S}_{5}$, the actoin of $(12)\in\mathrm{S}_{5}$ is non-symplectic. Replacing $x_{5}$ and $y$ by $y_{1}$ and $y_{2}$, we see that this is a subspace of (4.20). We have $\dim\mathrm{C}_{\rho_{1}}(\mathrm{A}_{5})=2^{2}+1^{2}=5$ and
$\dim V_{3}(\rho_{1})-\dim\mathrm{C}_{\rho_{1}}(\mathrm{A}_{5})=2=20-r(\mathrm{A}_{5}).$
The representation $\rho_{2}$ is given by permutations of projective coordinates $[x_{1}:\cdots:x_{6}]$ by “non-standard” $\mathrm{A}_{5}\subset\mathrm{A}_{6}$ generated by $(12345)$ and $(35)(46)$ (see Appendix). Putting $x_{ijk}=x_{i}x_{j}x_{k}$ and
$I_{1}$ $=x_{124}+x_{134}+x_{135}+x_{235}+x_{245}+x_{126}+x_{236}+x_{346}+x_{156}+x_{456},$
$I_{2}$ $=x_{123}+x_{234}+x_{125}+x_{145}+x_{345}+x_{136}+x_{146}+x_{246}+x_{256}+x_{356},$
we have
(6.10) $V_{3}(\rho_{2})=\mathrm{Span}\{s_{1}(x)^{3},\quad s_{1}(x)s_{2}(x),\quad s_{3}(x),\quad I_{1},\quad I_{2}\}$
with a linear relaton
$s_{1}^{3}-6(I_{1}+I_{2})-3s_{1}s_{2}+2s_{3}=0.$
A general $F\in V_{3}(\rho_{2})$ defines a smooth cubic fourfold since $s_{3}(x)\in V_{3}(\rho_{2})$ defines the Fermat cubic. We have $\mathrm{C}_{\rho_{2}}(\mathrm{A}_{5})=1^{2}+1^{2}=2$ and
$\dim V_{3}(\rho_{2})-\dim\mathrm{C}_{\rho_{2}}(\mathrm{A}_{5})=2=20-r(\mathrm{A}_{5})$
as desired.
6.4. The group $\mathbf{G\cong 3^{2}.4}$ with GAP-ID (36,9)
This group $G$ is represented as a (maximal) subgroup of $\mathrm{A}_{6}$ generated by $(123),\ (456),\ (1425)(36)$, and containing $\mathrm{A}_{3,3}=3^{2}:2$ as a subgroup of index 2. We have a Schur cover $G^{*}\cong 3.G$ with GAP-ID (108,15). Firstly, we consider linear representations of $G$ itself. The character table of $G$ is as follows.
$\begin{array}[]{c|rrrrrr}\rm class&1&\rm 2A&\rm 3A&\rm 3B&\rm 4A&\rm 4B\cr\rm size&1&9&4&4&9&9\cr\hline\cr\eta_{1}&1&1&1&1&1&1\cr\eta_{2}&1&1&1&1&-1&-1\cr\eta_{3}&1&-1&1&1&i&-i\cr\eta_{4}&1&-1&1&1&-i&i\cr\eta_{5}&4&0&1&-2&0&0\cr\eta_{6}&4&0&-2&1&0&0\cr\end{array}\qquad\begin{array}[]{l}\rm 2A:(12)(45)\\ \rm 3A:(123)\\ \rm 3B:(123)(456)\\ \rm 4A:(1425)(36)\\ \rm 4B:(1524)(36)\end{array}$
We have $\mathrm{Out}(G)\cong\mathrm{C}_{2}\times\mathrm{C}_{2}$, and it interchanges classes nA with nB for $n=3,4$. Since we have $\widehat{G}=\mathrm{C}_{4}$, we consider only invariants with the trivial character. Faithful characters satisfying the condition on absolute values are
$\phi_{1}=\eta_{1}+\eta_{2}+\eta_{5},\quad\phi_{2}=\eta_{1}+\eta_{2}+\eta_{6},$
$\phi_{1}\otimes\eta_{3}=\eta_{3}+\eta_{4}+\eta_{5},\quad\phi_{2}\otimes\eta_{3}=\eta_{3}+\eta_{4}+\eta_{6}.$
Since $\mathrm{Out}(G)$ interchanges $\eta_{5}$ with $\eta_{6}$, we consider only $\phi_{1}$ and $\phi_{1}\otimes\eta_{3}$. Note that
$\phi_{1}|\mathrm{A}_{3,3}$ $=(\eta_{1}+\eta_{2}+\eta_{5})|\mathrm{A}_{3,3}=(2\chi_{1}+\chi_{3}+\chi_{4})=\theta_{34},$
$\phi_{1}\otimes\eta_{3}|\mathrm{A}_{3,3}$ $=\theta_{34}\otimes\chi_{2}$
for $\chi_{i}$ and $\theta_{34}$ in section 4.2. A representaion $\rho$ affording $\phi_{1}$ is given by permutatioins of coordinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is $(1425)(36)$-invariants of (4.6):
(6.11) $\mathrm{Span}\{s_{1}(x)^{3}+s_{1}(y)^{3},\quad s_{2}(x)s_{1}(x)+s_{2}(y)s_{1}(y),\quad s_{3}(x)+s_{3}(y),\\ s_{1}(x)s_{2}(y)+s_{1}(y)s_{2}(x),\quad s_{1}(x)s_{1}(y)^{2}+s_{1}(x)^{2}s_{1}(y)\}.$
We have $\dim\mathrm{C}_{\rho}(G)=1^{2}+1^{2}+1^{2}=3$ and
$\dim V_{3}(\rho)-\dim\mathrm{C}_{\rho}(G)=2=20-r(G).$
For a representaion $\rho\otimes\eta_{3}$, there is no smooth cubic since $V(\rho\otimes\eta_{3})$ is a subspace of (4.7).
Next, we consider $G^{*}=3.G$. Since we have $\widehat{G^{*}}\cong\widehat{G}\cong\mathrm{C}_{4}$, we do not have to consider semi-invariants. Irreducible representaions of $G^{*}$ with a non-trivial action of the Schur multiplier are given as follows.
$\begin{array}[]{c|rrrrrrrrrrrrrr}\rm class&\rm 1&\rm 2&\rm 3A&\rm 3B&\rm 3C&\rm 3D&\rm 4A&\rm 4B&\rm 6A&\rm 6B&\rm 12A&\rm 12B&\rm 12C&\rm 12D\cr\rm size&1&9&1&1&12&12&9&9&9&9&9&9&9&9\cr\hline\cr\eta_{7}&3&1&\alpha&\overline{\alpha}&0&0&i&-i&\omega&\overline{\omega}&-\delta&\overline{\delta}&-\overline{\delta}&\delta\cr\eta_{8}&3&1&\alpha&\overline{\alpha}&0&0&-i&i&\omega&\overline{\omega}&\delta&-\overline{\delta}&\overline{\delta}&-\delta\cr\eta_{9}&3&-1&\alpha&\overline{\alpha}&0&0&1&1&-\omega&-\overline{\omega}&\omega&\overline{\omega}&\overline{\omega}&\omega\cr\eta_{10}&3&-1&\alpha&\overline{\alpha}&0&0&-1&-1&-\omega&-\overline{\omega}&-\omega&-\overline{\omega}&-\overline{\omega}&-\omega\cr\eta_{11}&3&1&\overline{\alpha}&\alpha&0&0&-i&i&\overline{\omega}&\omega&-\overline{\delta}&\delta&-\delta&\overline{\delta}\cr\eta_{12}&3&1&\overline{\alpha}&\alpha&0&0&i&-i&\overline{\omega}&\omega&\overline{\delta}&-\delta&\delta&-\overline{\delta}\cr\eta_{13}&3&-1&\overline{\alpha}&\alpha&0&0&1&1&-\overline{\omega}&-\omega&\overline{\omega}&\omega&\omega&\overline{\omega}\cr\eta_{14}&3&-1&\overline{\alpha}&\alpha&0&0&-1&-1&-\overline{\omega}&-\omega&-\overline{\omega}&-\omega&-\omega&-\overline{\omega}\cr\end{array}$
$\alpha=3\omega,\quad\delta=i\omega,\quad\overline{\delta}=-i\overline{\omega}$
Characters $\eta_{11},\cdots,\eta_{14}$ are complex conjugates of $\eta_{7},\cdots,\eta_{10}$. Classes 1, 3A and 3B form the Schur multiplier $\mathrm{C}_{3}$ which act as scalar matrices for
$\eta_{i}+\eta_{j}\quad(7\leq i,j\leq 10),\qquad\eta_{k}+\eta_{l}\quad(11\leq k,l\leq 14).$
The condition $(\eta_{i}+\eta_{j})(4A)=0$ is satisfied only by
$\eta_{7}+\eta_{8},\quad\eta_{9}+\eta_{10},\quad\eta_{11}+\eta_{12},\quad\eta_{13}+\eta_{14}.$
For representations $\rho_{i}$ affording $\eta_{i}$, we have Molien serieses
$\mu(\rho_{i}+\rho_{j},\,t)=\sum_{d=0}^{\infty}(\dim V_{d}(\rho_{i}+\rho_{j}))t^{d}=1+4t^{3}+\cdots\qquad((i,j)=(7,8),(11,12))$
$\mu(\rho_{i}+\rho_{j},\,t)=\sum_{d=0}^{\infty}(\dim V_{d}(\rho_{i}+\rho_{j}))t^{d}=1+20t^{6}+\cdots\qquad((i,j)=(9,10),(13,14)),$
and there is no invariant cubic for $\rho_{9}+\rho_{10}$ and $\rho_{13}+\rho_{14}$.
Let us consider a matrix $\tilde{h}=h\oplus(-h)$ where
$h=\frac{1}{\sqrt{3}}\begin{bmatrix}1&1&1\\ 1&\omega&\omega^{2}\\ 1&\omega^{2}&\omega\end{bmatrix}.$
[…]
$V_{3}(\rho_{7}+\rho_{8})=V_{3}(\overline{\rho_{7}}+\overline{\rho_{8}})=V_{3}(\rho_{11}+\rho_{13})$
The space $V_{3}(\rho_{7}+\rho_{8})$ is $\tilde{h}$-invariants of (4.11):
(6.12) $\mathrm{Span}\{s_{3}(x)-3u_{-}x_{1}x_{2}x_{3},\quad s_{3}(y)-3u_{+}y_{1}y_{2}y_{3},\quad I_{1}-u_{-}I_{2},\quad I_{3}-u_{+}I_{4}\}$
where $u_{\pm}=1\pm\sqrt{3}$. We have $\dim\mathrm{C}_{\rho_{7}+\rho_{8}}(G^{*})=1^{2}+1^{2}=2$ and
$\dim V_{3}(\rho_{7}+\rho_{8})-\dim\mathrm{C}_{\rho_{7}+\rho_{8}}(G^{*})=2=20-r(G).$
6.5. The groups $\mathbf{\mathrm{S}_{3,3}=\mathrm{S}_{3}\times\mathrm{S}_{3}}$
The Schur multiplier of $\mathrm{S}_{3,3}$ is $\mathrm{C}_{2}$, and we do not have to consider a Schur cover. The character table of $\mathrm{S}_{3}$ is
$\begin{array}[]{c|rrr}\rm class&\rm 1&(12)&(123)\cr\rm size&1&3&2\cr\hline\cr\eta_{1}&1&1&1\cr\eta_{2}&1&-1&1\cr\eta_{3}&2&0&-1\cr\end{array}$
and irreducible characters of $\mathrm{S}_{3,3}$ are given as tensor products $\eta_{ij}=\eta_{i}\boxtimes\eta_{j}$ where
$\eta_{ij}(a,b)=\eta_{i}(a)\eta_{j}(b)\qquad(a,b)\in\mathrm{S}_{3}\times\mathrm{S}_{3},$
and the character table of $\mathrm{S}_{3,3}$ is as follows.
$\begin{array}[]{c|rrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 2C&\rm 3A&\rm 3B&\rm 3C&\rm 6A&\rm 6B\cr\rm size&1&3&3&9&2&2&4&6&6\cr\hline\cr\eta_{11}&1&1&1&1&1&1&1&1&1\cr\eta_{21}&1&-1&1&-1&1&1&1&1&-1\cr\eta_{12}&1&1&-1&-1&1&1&1&-1&1\cr\eta_{22}&1&-1&-1&1&1&1&1&-1&-1\cr\eta_{13}&2&2&0&0&2&-1&-1&0&-1\cr\eta_{23}&2&-2&0&0&2&-1&-1&0&1\cr\eta_{31}&2&0&2&0&-1&2&-1&-1&0\cr\eta_{32}&2&0&-2&0&-1&2&-1&1&0\cr\eta_{33}&4&0&0&0&-2&-2&1&0&0\cr\end{array}\qquad\begin{array}[]{l}\rm 2A=(12)\\ \rm 2B=(45)\\ \rm 2C=(12)(45)\\ \rm 3A=(123)\\ \rm 3B=(456)\\ \rm 3C=(123)(456)\\ \rm 6A=(123)(45)\\ \rm 6B=(12)(456)\\ \end{array}$
Since we have $\widehat{\mathrm{S}_{3,3}}\cong C_{2}\times\mathrm{C}_{2}$,we consider only invariants with the trivial character. We see that $\mathrm{S}_{3,3}$-invariant cubics are given as subspaces of $\mathrm{A}_{3,3}$-invariant cubics $V_{3}(\rho)$ in (4.6). We have
$\eta_{11}|\mathrm{A}_{3,3}$ $=\eta_{22}|\mathrm{A}_{3,3}=\chi_{1},\quad\eta_{21}|\mathrm{A}_{3,3}=\eta_{12}|\mathrm{A}_{3,3}=\chi_{2},$
$\eta_{13}|\mathrm{A}_{3,3}$ $=\eta_{23}|\mathrm{A}_{3,3}=\chi_{3},\quad\eta_{31}|\mathrm{A}_{3,3}=\eta_{32}|\mathrm{A}_{3,3}=\chi_{4},\quad\eta_{33}|\mathrm{A}_{3,3}=\chi_{5}+\chi_{6}.$
for $\chi_{i}$ in section 4.2, and a character $\phi$ satisfying $|\phi(2A)|=2$ and
$\phi|\mathrm{A}_{3,3}=\theta_{34}=2\chi_{1}+\chi_{3}+\chi_{4}$
is one of the follwoings:
$\phi_{1}=\eta_{11}+\eta_{22}+\eta_{13}+\eta_{31},\quad\phi_{1}\otimes\eta_{22}=\eta_{11}+\eta_{22}+\eta_{23}+\eta_{32},$
$\phi_{2}=\eta_{11}+\eta_{22}+\eta_{13}+\eta_{32},\quad\phi_{2}\otimes\eta_{22}=\eta_{11}+\eta_{22}+\eta_{23}+\eta_{31}.$
Let us consider a representation $\rho_{1}$ affording $\phi_{1}$. Permutations of coordinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ by $\mathrm{A}_{3,3}$ is normalized by an action of a matrix
$b_{1}=\frac{1}{3}\begin{bmatrix}-2&1&-2\\ 1&-2&-2\\ -2&-2&1\end{bmatrix}\oplus\mathrm{diag}(1,1,1)$
with eigenvalues $(-1,-1,1,1,1,1)$. These actions generate $\mathrm{S}_{3,3}$ and this gives $\rho_{1}$. By the involution $b_{1}$, the space $V_{3}(\rho)$ in (4.6) is decomposed into the $(+1)$-eigenspace
(6.13) $V_{3}(\rho_{1})=\mathrm{Span}\{s_{1}(y)^{3},\ s_{2}(y)s_{1}(y),\ s_{3}(y),\ s_{1}(x)^{2}s_{1}(y),\ s_{2}(x)s_{1}(y),\ p(x)\}$
where $p(x)=2s_{1}(x)^{3}-9s_{1}(x)s_{2}(x)+9s_{3}(x)$, and the $(-1)$-eigenspace
(6.14) $V_{3}(\rho_{1},\eta_{22})=V_{3}(\rho_{1}\otimes\eta_{22})\\ =\mathrm{Span}\{s_{1}(x)^{3},\ s_{2}(x)s_{1}(x),\ s_{1}(x)s_{2}(y),\ s_{1}(x)s_{1}(y)^{2}\}.$
A general member of $V_{3}(\rho_{1})$ defines a smooth cubic fourfold since
$s_{3}(y)+s_{2}(x)s_{1}(y)+p(x)=0$
is smooth. On the other hand, $F\in V_{3}(\rho_{1}\otimes\eta_{22})$ is reducible. We have
$\dim\mathrm{C}_{\rho_{1}}(\mathrm{S}_{3,3})=1^{2}+1^{2}+1^{2}+1^{2}=4$
and
$\dim V_{3}(\rho_{1})-\dim\mathrm{C}_{\rho_{1}}(\mathrm{S}_{3,3})=2=20-r(\mathrm{S}_{3,3}).$
Similarly, $A_{3,3}$ and an involution
$b_{2}=\begin{bmatrix}0&-1&0\\ -1&0&0\\ 0&0&-1\end{bmatrix}\oplus\mathrm{diag}(1,1,1)$
generate $\mathrm{S}_{3,3}$, which gives a representation $\rho_{2}$ affording $\phi_{2}$. We can decompose $V_{3}(\rho)$ into $(+1)$-eigenspace of $b_{2}$
$V_{3}(\rho_{2})=\mathrm{Span}\{s_{1}(y)^{3},\ s_{2}(y)s_{1}(y),\ s_{3}(y),\ s_{1}(y)s_{2}(x),\ s_{1}(x)^{2}s_{1}(y)\},$
and $(-1)$-eigenspace of $b_{2}$
$V_{3}(\rho_{2},\eta_{22})=V_{3}(\rho_{2}\otimes\eta_{22})\\ =\mathrm{Span}\{s_{1}(x)^{3},\ s_{2}(x)s_{1}(x),\ s_{3}(x),\ s_{1}(x)s_{2}(y),\ s_{1}(x)s_{1}(y)^{2}\}.$
For $F\in V_{3}(\rho_{2})$, a cubic fourfold $F=0$ is singular along $s_{1}(y)=s_{3}(y)=0$. Similarly, $F\in V_{3}(\rho_{2}\otimes\eta_{22})$ defines a singular cubic.
6.6. The group $\mathbf{Hol}(5)$
[…]
$\begin{array}[]{c|rrrrr}\rm class&\rm 1&(12)(34)&(2345)&(2534)&(12354)\cr\rm size&1&5&5&5&4\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&-1&1\cr\chi_{3}&1&-1&-i&i&1\cr\chi_{4}&1&-1&i&-i&1\cr\chi_{5}&4&0&0&0&-1\cr\end{array}$
Since $\widehat{\mathrm{Hol}(5)}\cong\mathrm{C}_{4}$, we consider only invariants with the trivial character. Faithful characters satisfying the condition on absolute values are
$\theta=\chi_{1}+\chi_{2}+\chi_{5},\quad\theta\otimes\chi_{3}=\chi_{3}+\chi_{4}+\chi_{5}.$
Let $\rho$ be a representation $\rho$ affording $\theta$. We have Molien series
$\mu(\rho,t)=1+t+3t^{2}+5t^{3}+\cdots,\qquad\mu(\rho\otimes\chi,t)=1+2t^{2}+t^{3}+\cdots$
and
$\dim V_{3}(\rho)=5,\qquad\dim V_{3}(\rho\otimes\chi_{3})=1.$
Therefore $\rho\otimes\chi_{3}$ is not suitable. The representation $\rho$ is given by
$(12)(34)[x_{1}:\cdots:x_{5}:y]=[x_{2}:x_{1}:x_{4}:x_{3}:x_{5}:y],$
$(2345)[x_{1}:\cdots:x_{5}:y]=[x_{1}:x_{3}:x_{4}:x_{5}:x_{2}:-y],$
which satisfies symplectic conditions. Invariant cubics are
(6.15) $\mathrm{Span}\{s_{1}(x)^{3},\quad s_{1}(x)s_{2}(x),\quad s_{1}(x)y^{2},\quad s_{3}(x),\quad Q(x)y\}$
where
(6.16) $Q(x)=-x_{1}x_{2}+x_{1}x_{3}-x_{2}x_{3}-x_{1}x_{4}+x_{2}x_{4}\\ +x_{3}x_{4}+x_{1}x_{5}+x_{2}x_{5}-x_{3}x_{5}-x_{4}x_{5}.$
We have $\dim\mathrm{C}_{\rho}(\mathrm{Hol}(5))=1^{2}+1^{2}+1^{2}$ and
$\dim V_{3}(\rho)-\dim\mathrm{C}_{\rho}(\mathrm{Hol}(5))=2=20-r(\mathrm{Hol}(5)).$
7. Groups of $r(G)=19$
We have the following seven groups
$3^{1+4}:2.2,\quad\mathrm{A}_{6},\quad\mathrm{L}_{2}(7),\quad\mathrm{S}_{5},\quad\mathrm{M}_{9},\quad\mathrm{N}_{72},\quad\mathrm{T}_{48},$
and $\mathrm{A}_{6}$ was already considered in section 2.5.
7.1. The group $\mathbf{3^{1+4}_{+}:2.2}$ with GAP-ID (972, 776)
This group $G$ contains $H\cong 3^{1+4}_{+}:2$ as a normal subgroup of index 2, and therefore cubic fourfolds such that $G\subset\mathrm{Aut}^{s}(X)$ are given as subfamilies of (6.1) which are normalized into the Hesse normal forms
$F_{\alpha}(x_{1},x_{2},x_{3})=F_{\beta}(y_{1},y_{2},y_{3})\quad\text{where}\quad F_{\lambda}(x,y,z)=x^{3}+y^{3}+z^{3}-27\lambda xyz.$
For this normal form, $H\subset\mathrm{PGL}_{6}(\mathbb{C})$ is generated by natural actions of $\mathrm{A}_{3,3}$ and
$a=\mathrm{diag}(1,1,1,\omega,\omega,\omega),\quad b_{1}=\mathrm{diag}(1,1,1,1,\omega,\omega^{2}),\quad b_{2}=\mathrm{diag}(1,\omega,\omega^{2},1,1,1).$
The center $Z(H)$ of $H$ is $\{1,a,a^{2}\}$, but $Z(G)$ is trivial for the group $G$ of GAP-ID (972, 776). A normalizer $g\in\mathrm{PGL}_{6}(\mathbb{C})$ of $H$ acts on $Z(H)$, and hence $g$ takes the form of $\begin{bmatrix}A&\mathbf{0}\\ \mathbf{0}&D\end{bmatrix}$ or $\begin{bmatrix}\mathbf{0}&B\\ C&\mathbf{0}\end{bmatrix}$. In the fomer case, $g$ commutes with $Z(H)$ and we see that $g$ and $H$ generate a group with non-trivial center. In the later cases, two cubic curves $F_{\alpha}(x)=0$ and $F_{\beta}(y)=0$ are isomorphic if $g$ preserves the equation $F_{\alpha}(x)=F_{\beta}(y)$. Therefore, cubic fourfolds such that $G\subset\mathrm{Aut}^{s}(X)$ are given by equations
(7.1) $F_{\alpha}(x_{1},x_{2},x_{3})=F_{\alpha}(y_{1},y_{2},y_{3})$
and $G$ is generated by $H$ and
$g=\begin{bmatrix}\mathbf{0}&B\\ \mathbf{1}&\mathbf{0}\end{bmatrix}\quad\text{where}\quad B=\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{bmatrix}$
whose eigenvalues are $(1,1,-1,-1,i,-i)$.
7.2. The projective special linear group $\mathrm{L}_{2}(7)$
The group $\mathrm{L}_{2}(7)=\mathrm{PSL}_{2}(\mathbb{F}_{7})$ is isomorphic to a maximal subgroup of $\mathrm{A}_{7}$ generated by $(1,2)(3,6)$ and $(1,2,3,4,5,6,7)$. The Schur multiplier of $\mathrm{L}_{2}(7)$ is $\mathrm{C}_{2}$, and we do not have to consider a Schur cover. The character table of $\mathrm{L}_{2}(7)$ is as follows:
$\begin{array}[]{c|rrrrrr}\rm class&\rm 1&\rm 2A&\rm 3A&\rm 4A&\rm 7A&\rm 7B\cr\rm size&1&21&56&42&24&24\cr\hline\cr\chi_{1}&1&1&1&1&1&1\cr\chi_{2}&3&-1&0&1&\frac{-1+\sqrt{-7}}{2}&\frac{-1-\sqrt{-7}}{2}\cr\chi_{3}&3&-1&0&1&\frac{-1-\sqrt{-7}}{2}&\frac{-1+\sqrt{-7}}{2}\cr\chi_{4}&6&2&0&0&-1&-1\cr\chi_{5}&7&-1&1&-1&0&0\cr\chi_{6}&8&0&-1&0&1&1\cr\end{array}\qquad\begin{array}[]{l}\rm 2A:(12)(36)\\ \rm 3A:(235)(476)\\ \rm 4A:(2347)(56)\\ \rm 7A:(1234567)\\ \rm 7B:(1237645)\end{array}$
We see that $\chi_{4}$ is the unique character of degree 6 such that $\chi_{4}(4\mathrm{A})=0$. The corresponding cubic fourfolds are given as
(7.2) $\begin{cases}x_{1}+x_{2}+x_{3}+x_{4}+x_{5}+x_{6}+x_{7}=0\\ c_{1}(x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{4}^{3}+x_{5}^{3}+x_{6}^{3}+x_{7}^{3})\\ \qquad\qquad+c_{2}(x_{124}+x_{235}+x_{346}+x_{156}+x_{137}+x_{457}+x_{267})=0\end{cases}$
in $\mathbb{P}^{6}$ where $x_{ijk}=x_{i}x_{j}x_{k}$. This family contains the Clebsch-Segre cubic (2.1), and a general menber is smooth.
7.3. The symmetric group $\mathrm{S}_{5}$
The Schur multiplier of $\mathrm{S}_{5}$ is $\mathrm{C}_{2}$, and we do not have to consider a Schur cover. The character table of $\mathrm{S}_{5}$ is as follows.
$\begin{array}[]{c|rrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 3A&\rm 4A&\rm 5A&\rm 6A\cr\rm size&1&10&15&20&30&24&20\cr\hline\cr\eta_{1}&1&1&1&1&1&1&1\cr\eta_{2}&1&-1&1&1&-1&1&-1\cr\eta_{3}&4&-2&0&1&0&-1&1\cr\eta_{4}&4&2&0&1&0&-1&-1\cr\eta_{5}&5&1&1&-1&-1&0&1\cr\eta_{6}&5&-1&1&-1&1&0&-1\cr\eta_{7}&6&0&-2&0&0&1&0\cr\end{array}\qquad\begin{array}[]{l}\rm 2A:(12)\\ \rm 2B:(12)(34)\\ \rm 3A:(123)\\ \rm 4A:(1234)\\ \rm 5A:(12345)\\ \rm 6A:(12)(345)\end{array}$
Since we have $\widehat{\mathrm{S}_{5}}=\mathrm{C}_{2}$, we consider only invariants with the trivial character. Faithful characters satisfying the conditions on absolute values are
$\phi_{1}=\eta_{1}+\eta_{2}+\eta_{3},\quad\phi_{1}\otimes\eta_{2}=\eta_{1}+\eta_{2}+\eta_{4}.,$
$\phi_{2}=\eta_{1}+\eta_{5},\quad\phi_{2}\otimes\eta_{2}=\eta_{2}+\eta_{6}.$
Note that we have
$\phi_{1}|\mathrm{A}_{5}=(\phi_{1}\otimes\eta_{2})|\mathrm{A}_{5}=\theta_{1},\quad\phi_{2}|\mathrm{A}_{5}=(\phi_{2}\otimes\eta_{2})|\mathrm{A}_{5}=\theta_{2}$
for $\theta_{1}$ and $\theta_{2}$ in (6.8). By extending actions of $\mathrm{A}_{5}$ in section 6.3, we can obtain invariant cubics as subspaces of $V_{3}(\rho_{1})$ in (6.9) and $V_{3}(\rho_{2})$ in (6.10). The character $\phi_{1}$ is afforded by $\pi_{1}:\mathrm{S}_{5}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ such that
$\pi(\sigma)[x_{1}\cdots:x_{5}:y]=[x_{\sigma(1)}:\cdots:x_{\sigma(5)}:(\mathrm{sgn}\,\sigma)y],$
and this satisfies the symplectic conditions. We have $\pi_{1}|\mathrm{A}_{5}=\rho_{1}$ and $V_{3}(\rho_{1})$ is decomposed into (±$(\pm 1)$-eigenspaces for $(12)\in\mathrm{S}_{5}$
(7.3) $V_{3}(\rho_{1})_{+}$ $=V_{3}(\pi_{1})=\mathrm{Span}\{y^{2}s_{1}(x),\quad s_{1}(x)^{3},\quad s_{1}(x)s_{2}(x),\quad s_{3}(x)\}$
and
(7.4) $V_{3}(\rho_{1})_{-}$ $=V_{3}(\pi_{1}\otimes\eta_{2})=\mathrm{Span}\{y^{3},\quad ys_{1}(x)^{2},\quad ys_{2}(x)\}.$
A general menber of $V_{3}(\pi_{1})$ defines a smooth cubic 4-fold since $y^{2}s_{1}(x)+s_{3}(x)=0$ is smooth, but $F\in V_{3}(\pi_{1}\otimes\eta_{2})$ is reducible. We have $\dim\mathrm{C}_{\pi_{1}}(\mathrm{S}_{5})=1^{2}+1^{2}+1^{2}=3$ and
$\dim V_{3}(\pi_{1})-\dim\mathrm{C}_{\pi_{1}}(\mathrm{A}_{5})=1=20-r(\mathrm{S}_{5}).$
Next we consider a representation $\pi_{2}$ affording $\phi_{2}$. Let us consider the following involution with eigenvalues $(1,1,1,1,-1,-1)$:
(7.5) $a=\frac{1}{3}[a_{ij}]_{6\times 6}\qquad a_{ij}=\begin{cases}-2\ (i+j\equiv 3\mod 6)\\ 1\ (i+j\not\equiv 3\mod 6)\end{cases}.$
It normalizes $\rho_{2}(\mathrm{A}_{5})$ in section 6.3, and generates $\mathrm{S}_{5}$ with $\rho_{2}(\mathrm{A}_{5})$. This involution decomposes $V_{3}(\rho_{2})$ into
(7.6) $V_{3}(\rho_{2})_{+}$ $=V_{3}(\pi_{2})=\mathrm{Span}\{s_{1}(x)^{3},\quad s_{1}(x)s_{2}(x),\quad I_{1}-I_{2}\}$
and
(7.7) $V_{3}(\rho_{2})_{-}$ $=V_{3}(\pi_{2}\otimes\eta_{2})=\mathrm{Span}\{s_{1}(x)^{3}-9s_{1}(x)s_{2}(x)+18s_{3}(x)\}.$
A general menber of $V_{3}(\pi_{2})$ defines a smooth cubic 4-fold since $s_{1}(x)s_{2}(x)+(I_{1}-I_{2})=0$ is smooth. On the other hand,
$s_{1}(x)^{3}-9s_{1}(x)s_{2}(x)+18s_{3}(x)=0$
is singular at $[1:\cdots:1]$. We have $\dim\mathrm{C}_{\pi_{2}}(\mathrm{S}_{5})=1^{2}+1^{2}=2$ and
$\dim V_{3}(\pi_{2})-\dim\mathrm{C}_{\pi_{2}}(\mathrm{A}_{5})=1=20-r(\mathrm{S}_{5}).$
7.4. The Mathieu group $\mathrm{M}_{9}$
We have $\mathrm{M}_{9}\cong 3^{2}.\mathrm{Q}_{8}$ and this group contains $43^{2}.$4 in section 6.4 as a subgroup of index 2. The Schur multiplier of $\mathrm{M}_{9}$ is $\mathrm{C}_{3}$, and we must consider a Schur cover $3.\mathrm{M}_{9}$. The character table of $\mathrm{M}_{9}$ itself is as follows:
$\begin{array}[]{c|rrrrrr}\rm class&\rm 1&\rm 2A&\rm 3A&\rm 4A&\rm 4B&\rm 4C\cr\rm size&1&9&8&18&18&18\cr\hline\cr\chi_{1}&1&1&1&1&1&1\cr\chi_{2}&1&1&1&-1&1&-1\cr\chi_{3}&1&1&1&1&-1&-1\cr\chi_{4}&1&1&1&-1&-1&1\cr\chi_{5}&2&-2&2&0&0&0\cr\chi_{6}&8&0&-1&0&0&0\cr\end{array}$
For any 6-dimensional representation $\rho$, we have $\mathrm{Tr}_{\rho}(3\mathrm{A})=6$ and this is not the cases. Next, let us consider a Schur cover $3.\mathrm{M}_{9}$ with GAP-ID (216, 88). Since we have $\widehat{\mathrm{M}_{9}}=\mathrm{C}_{2}\times\mathrm{C}_{2}$, we consider only invariants with the trivial character. Irreducible representaions of $3.\mathrm{M}_{9}$ with a non-trivial action of the Schur multiplier are given as follows.
$\begin{array}[]{c|rrrrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 3A&\rm 3B&\rm 3C&\rm 4A&\rm 4B&\rm 4C&\rm 6A&\rm 6B&\cdots\cr\rm size&1&9&1&1&24&18&18&18&9&9&\cdots\cr\hline\cr\chi_{7}&3&-1&\alpha&\overline{\alpha}&0&-1&-1&1&\overline{\gamma}&\gamma\\ \chi_{8}&3&-1&\alpha&\overline{\alpha}&0&-1&1&-1&\overline{\gamma}&\gamma\\ \chi_{9}&3&-1&\alpha&\overline{\alpha}&0&1&-1&-1&\overline{\gamma}&\gamma\\ \chi_{10}&3&-1&\alpha&\overline{\alpha}&0&1&1&1&\overline{\gamma}&\gamma\\ \chi_{11}&3&-1&\overline{\alpha}&\alpha&0&-1&-1&1&\gamma&\overline{\gamma}\\ \chi_{12}&3&-1&\overline{\alpha}&\alpha&0&-1&1&-1&\gamma&\overline{\gamma}\\ \chi_{13}&3&-1&\overline{\alpha}&\alpha&0&1&-1&-1&\gamma&\overline{\gamma}\\ \chi_{14}&3&-1&\overline{\alpha}&\alpha&0&1&1&1&\gamma&\overline{\gamma}\\ \chi_{15}&6&2&\beta&\overline{\beta}&0&0&0&0&\overline{\delta}&\delta\\ \chi_{16}&6&2&\overline{\beta}&\beta&0&0&0&0&\delta&\overline{\delta}\end{array}$
$\alpha=3\omega,\quad\beta=6\omega,\quad\gamma=-\omega,\quad\delta=2\omega$
Classes 1, 3A and 3B form the Schur multiplier $\mathrm{C}_{3}$ which act as scalar matrices for
$\chi_{i}+\chi_{j}\quad(7\leq i,j\leq 10),\qquad\chi_{k}+\chi_{l}\quad(11\leq k,l\leq 14).\quad\chi_{15},\quad\chi_{16}.$
Among these characters, conditions
$\chi({\rm 4A})=\chi({\rm 4B})=\chi({\rm 4C})=0$
[…]
(7.8) $h_{2}=\frac{1}{\sqrt{3}}\begin{bmatrix}\omega&\omega^{2}&\omega^{2}\\ 1&1&\omega^{2}\\ 1&\omega^{2}&1\end{bmatrix}.$
The group $H_{2}$ together with $g$ generate a group $H_{3}$ with GAP-ID (216, 88), and this representation $\rho$ affords $\chi_{15}$ or $\overline{\chi_{15}}$. Moreover we have $\overline{H_{3}}=\delta H_{3}\delta^{-1}$ for $\delta=\mathrm{diag}(1,1,1,i\omega,i\omega,i\omega)$ since we have $\overline{H_{2}}=H_{2}$ (see section 6.4) and
$\overline{g}=\delta g^{-1}\tilde{h}\delta^{-1}\in\delta H_{3}\delta^{-1},\qquad\delta H_{2}\delta^{-1}=H_{2}$
where $\tilde{h}$ was given in section 6.4. Therefore we have $V_{3}(\overline{\rho})=\delta^{*}V_{3}(\rho)$, and we do not have to consider $\overline{\rho}$. The space $V_{3}(\rho)$ consists of $g$-invariants in (6.12):
(7.9) $\mathrm{Span}\{(s_{3}(x)-3u_{-}x_{1}x_{2}x_{3})+(\omega+i\omega^{2})(s_{3}(y)-3u_{+}y_{1}y_{2}y_{3}),\\ (i-\omega^{2})(I_{1}-u_{-}I_{2})+(I_{3}-u_{+}I_{4})\}.$
We have $\dim\mathrm{C}_{\rho}(3.\mathrm{M}_{9})=1$ and
$\dim V_{3}(\rho)-\dim\mathrm{C}_{\rho}(3.\mathrm{M}_{9})=1=20-r(\mathrm{M}_{9}).$
7.5. The group $\mathbf{N_{72}\cong(S_{3}\times S_{3}):2}$ with GAP-ID (72,40)
This is a subgroup of $\mathrm{S}_{6}$ generated by
$(12),\quad(123),\quad(45),\quad(456),\quad(14)(25)(36).$
The Schur multiplier of $\mathrm{N}_{72}$ is $\mathrm{C}_{2}$, and we do not have to consider a Schur cover. The character table of $\mathrm{N}_{72}$ is as follows.
$\begin{array}[]{c|rrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 2C&\rm 3A&\rm 3B&\rm 4A&\rm 6A&\rm 6B\cr\rm size&1&6&6&9&4&4&18&12&12\cr\hline\cr\xi_{1}&1&1&1&1&1&1&1&1&1\cr\xi_{2}&1&-1&-1&1&1&1&1&-1&-1\cr\xi_{3}&1&-1&1&1&1&1&-1&-1&1\cr\xi_{4}&1&1&-1&1&1&1&-1&1&-1\cr\xi_{5}&2&0&0&-2&2&2&0&0&0\cr\xi_{6}&4&-2&0&0&1&-2&0&1&0\cr\xi_{7}&4&0&-2&0&-2&1&0&0&1\cr\xi_{8}&4&0&2&0&-2&1&0&0&-1\cr\xi_{9}&4&2&0&0&1&-2&0&-1&0\cr\end{array}\qquad\begin{array}[]{l}\rm 2A=(12)\\ \rm 2B=(14)(25)(36)\\ \rm 2C=(12)(45)\\ \rm 3A=(123)\\ \rm 3B=(123)(456)\\ \rm 4A=(14)(2536)\\ \rm 6A=(12)(456)\\ \rm 6B=(142536)\\ \end{array}$
Since we have $\widehat{\mathrm{N}_{72}}\cong C_{2}\times\mathrm{C}_{2}$,we consider only invariants with the trivial character. We see that $\mathrm{N}_{72}$-invariant cubics are given as subspaces of $\mathrm{S}_{3,3}$-invariant cubics $V_{3}(\rho_{1})$ in (6.13). We have
$\xi_{1}|\mathrm{S}_{3,3}=\xi_{4}|\mathrm{S}_{3,3}=\eta_{11},\quad\xi_{2}|\mathrm{S}_{3,3}=\xi_{3}|\mathrm{S}_{3,3}=\eta_{22},\quad\xi_{5}|\mathrm{S}_{3,3}=\eta_{21}+\eta_{12},$
$\xi_{6}|\mathrm{S}_{3,3}=\eta_{23}+\eta_{32},\quad\xi_{9}|\mathrm{S}_{3,3}=\eta_{13}+\eta_{31},\quad\xi_{7}|\mathrm{S}_{3,3}=\xi_{8}|\mathrm{S}_{3,3}=\eta_{33}$
for $\eta_{ij}$ in section 6.5, and characters $\psi$ satisfying
$|\psi(2A)|=|\psi(2B)|=2,\quad\psi|\mathrm{S}_{3,3}=\phi_{1}=\eta_{11}+\eta_{22}+\eta_{13}+\eta_{31}$
are only
$\psi=\xi_{1}+\xi_{3}+\xi_{9},\quad\psi\otimes\xi_{2}=\xi_{2}+\xi_{4}+\xi_{9}.$
A representation $\pi$ affording the character $\psi$ is given by extending $\rho_{1}$ in secton 6.5. Let us consider a matrix $c=\begin{bmatrix}C_{1}&C_{2}\\ C_{2}&C_{1}\end{bmatrix}$ of order 2 with eigenvalues $(1,1,1,1,-1,-1)$, where
$C_{1}=\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{bmatrix},\quad C_{2}=\begin{bmatrix}-2&1&1\\ 1&-2&1\\ 1&1&-2\end{bmatrix}.$
The group $\rho_{1}(\mathrm{S}_{3,3})$ and $c$ generate a group with GAP-ID (72, 40), and this gives $\pi$. By the involution $c$, the space $V_{3}(\rho_{1})$ is decomposed into the $(+1)$-eigenspace
(7.10) $V_{3}(\pi)=\mathrm{Span}\{s_{1}(y)^{3},\ s_{1}(x)^{2}s_{1}(y),\ s_{2}(y)s_{1}(y)+s_{2}(x)s_{1}(y),\\ 9s_{2}(y)s_{1}(y)-9s_{3}(y)+p(x)\}$
and the $(-1)$-eigenspace
(7.11) $V_{3}(\pi,\xi_{2})=V_{3}(\pi\otimes\xi_{2})=\mathrm{Span}\{2s_{1}(y)^{3}-9s_{2}(y)s_{1}(y)-9s_{3}(y)+p(x),\\ s_{1}(y)^{3}-3s_{2}(y)s_{1}(y)-s_{1}(x)^{2}s_{1}(y)+3s_{2}(x)s_{1}(y)\}.$
A general member of $V_{3}(\pi)$ defines a smooth cubic fourfold since
$s_{1}(x)^{2}s_{1}(y)+(9s_{2}(y)s_{1}(y)-9s_{3}(y)+p(x))=0$
is smooth. For $F\in V_{3}(\pi\otimes\chi_{2})$, a cubic fourfold $F=0$ is singular along $x_{1}=x_{2}=x_{3},\ y_{1}=y_{2}=y_{3}$. We have $\dim\mathrm{C}_{\pi}(\mathrm{S}_{3,3}:2)=1^{2}+1^{2}+1^{2}=3$ and
$\dim V_{3}(\pi)-\dim\mathrm{C}_{\pi}(\mathrm{S}_{3,3}:2)=1=20-r(\mathrm{S}_{3,3}:2).$
7.6. The group $\mathbf{T_{48}\cong Q_{8}:S_{3}}$ with GAP-ID (48,29)
The Schur multiplier of $\mathrm{T}_{48}$ is 1, and hence we have $(\mathrm{T}_{48})^{*}=\mathrm{T}_{48}$. The character table of $\mathrm{T}_{48}$ is as follows.
$\begin{array}[]{c|rrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 3A&\rm 4A&\rm 6A&\rm 8A&\rm 8B\cr\rm size&1&1&12&8&6&8&6&6\cr\hline\cr\chi_{1}&1&1&1&1&1&1&1&1\cr\chi_{2}&1&1&-1&1&1&1&-1&-1\cr\chi_{3}&2&2&0&-1&2&-1&0&0\cr\chi_{4}&2&-2&0&-1&0&1&-\sqrt{-2}&\sqrt{-2}\cr\chi_{5}&2&-2&0&-1&0&1&\sqrt{-2}&-\sqrt{-2}\cr\chi_{6}&3&3&-1&0&-1&0&1&1\cr\chi_{7}&3&3&1&0&-1&0&-1&-1\cr\chi_{8}&4&-4&0&1&0&-1&0&0\cr\end{array}$
Since we have $\widehat{\mathrm{T}_{48}}=\mathrm{C}_{2}$, we consider only invariants with the trivial character. Characters $\theta$ of degree 6 satisfying
$|\theta(2A)|=|\theta(2B)|=2,\quad|\theta(4A)|=0,\quad|\theta(8A)|=|\theta(8B)|=\sqrt{2}$
are only
$\theta_{1}=\chi_{1}+\chi_{4}+\chi_{7},\quad\theta_{1}\otimes\chi_{2}=\chi_{2}+\chi_{5}+\chi_{6},$
$\overline{\theta_{1}}=\chi_{1}+\chi_{5}+\chi_{7},\quad\overline{\theta_{1}\otimes\chi_{2}}=\chi_{2}+\chi_{4}+\chi_{7}.$
Since we have Molien series
$\mu(\theta_{1},t)=1+t+2t^{2}+4t^{3}+\cdots,\qquad\mu(\theta_{1}\otimes\chi_{2},t)=1+2t^{2}+7t^{4}+\cdots,$
we consider only $\theta_{1}$ and $\overline{\theta_{1}}$. The group $T_{48}$ contains $\mathrm{Q}_{8}$ as a normal subgroup, and a representation $\pi$ affording $\theta_{1}$ is given by extending the representation $\rho$ in section 5.1. Let us consider two matrices
$n_{1}=\begin{bmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{bmatrix}\oplus\frac{1}{\sqrt{2}}\begin{bmatrix}1&-i\\ i&-1\end{bmatrix},\quad n_{2}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\oplus\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\oplus\begin{bmatrix}0&\zeta_{8}^{3}\\ \zeta_{8}^{5}&0\end{bmatrix}$
of order 2 with eigenvalues $(1,1,1,1,-1,-1)$. They normalize $\rho(\mathrm{Q}_{8})$, and $n_{1},n_{2}$ together with $\rho(\mathrm{Q}_{8})$ generate a group $H\cong\mathrm{Q}_{8}:\mathrm{S}_{3}$ with GAP-ID (48,29) (Conjugations by $n_{1}$ and $n_{2}$ permute conjugacy classes
$\{\rho(I),\ \rho(-I)\},\quad\{\rho(J),\ \rho(-J)\},\quad\{\rho(K),\ \rho(-K)\}$
of $\rho(\mathrm{Q}_{8})$, and they generate $\mathrm{Out}(\mathrm{Q}_{8})\cong\mathrm{S}_{3}$). Moreover we have $\overline{H}=\delta H\delta^{-1}$ for $\delta=\mathrm{diag}(1,1,-1,-1)\oplus\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$ since
$\overline{\rho(I)}=\delta\rho(I)^{3}\delta^{-1},\quad\overline{\rho(J)}=\delta\rho(J)\delta^{-1},\quad\overline{n_{1}}=\delta(n_{1}\rho(IJ))\delta^{-1},\quad\overline{n_{2}}=\delta(n_{2})\delta^{-1}.$
Therefore we have $V_{3}(\overline{\pi})=\delta^{*}V_{3}(\pi)$ and we do not have to consider $V_{3}(\overline{\pi})$. The space $V_{3}(\pi)$ is a subspace of (5.1):
(7.12) $\mathrm{Span}\{x_{1}^{3},\quad x_{1}(x_{2}^{2}+x_{3}^{2}+x_{4}^{2}),\quad x_{2}x_{3}x_{4},\\ ix_{3}(x_{5}^{2}+x_{6}^{2})+2x_{2}x_{5}x_{6}+x_{4}(x_{5}^{2}-x_{6}^{2})\}.$
A general menber of $V_{3}(\pi)$ is smooth since sum of the above four generator define a smooth cubic fourfold. We have $\dim\mathrm{C}_{\pi}(\mathrm{T}_{48})=1^{2}+1^{2}+1^{2}=3$ and
$\dim V_{3}(\pi)-\dim\mathrm{C}_{\pi}(\mathrm{T}_{48})=1=20-r(\mathrm{T}_{48}).$
8. Maximal groups
We have six maximal groups:
$3^{4}:\mathrm{A}_{6},\quad\mathrm{A}_{7},\quad 3^{1+4}_{+}:2.2^{2},\quad\mathrm{M}_{10},\quad\mathrm{L}_{2}(11),\quad\mathrm{A}_{3,5}.$
There exists the unique cubic fourfold $X$ such that $\mathrm{Aut}^{s}(X)=G$ except for $G=\mathrm{A}_{7}$ and $\mathrm{M}_{10}$. However, there exist two non-isomorphic cubic fourfolds $X$ such that $\mathrm{Aut}^{s}(X)$ is $\mathrm{A}_{7}$ or $\mathrm{M}_{10}$. Corresponding eight cubic fourfolds were identified in [HM19], [LZ22] and [YYZ23].
(i)  For $G=3^{4}:\mathrm{A}_{6}$, the corresponding cubic fourfold is the Fermat cubic
(8.1) $x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{4}^{3}+x_{5}^{3}+x_{6}^{3}=0.$
(ii)  For $G=\mathrm{A}_{7}$, corresponding two cubic fourfolds are (2.1) and (2.3).
(iii)  For $G=3^{1+4}_{+}:2.2^{2}$, the corresponding cubic fourfold is given by
$F(x_{1},x_{2},x_{3})=F(x_{4},x_{5},x_{6})$
where
(8.2) $F(x,y,z)=x^{3}+y^{3}+z^{3}-3(1+\sqrt{3})xyz.$
Note that $F(x,y,z)=0$ is an elliptic curve whose $j$-invariant is 1728 (see [Ko22]), and we may replace $F$ by
(8.3) $F(x,y,z)=y^{2}z-x^{3}-xz^{2}.$
In particular, this cubic fourfold is defined over $\mathbb{Q}$.
(iv)  For $G=\mathrm{L}_{2}(11)=\mathrm{PSL}_{2}(\mathbb{F}_{11})$, the corresponding cubic fourfold is given by
(8.4) $x_{6}^{3}=x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+x_{3}^{2}x_{4}+x_{4}^{2}x_{5}+x_{5}^{2}x_{1}.$
(v)  For $G=\mathrm{A}_{3,5}=(\mathrm{S}_{3}\times\mathrm{S}_{8})\cap\mathrm{A}_{8}$, the corresponding cubic fourfold is given by
(8.5) $\begin{cases}x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{4}^{3}+x_{5}^{3}+x_{6}^{3}+x_{7}^{3}+x_{8}^{3}=0\\ x_{1}+x_{2}+x_{3}=0\\ x_{4}+x_{5}+x_{6}+x_{7}+x_{8}=0\end{cases}$
(vi)  For $G=\mathrm{M}_{10}$, we give corresponding cubic fourfolds $X_{+}$ and $X_{-}$ defined over $\mathbb{Q}(\sqrt{6})$. The Mathieu group $\mathrm{M}_{10}$ contains $\mathrm{A}_{6}$ as a subgroup of index 2, and $X_{\pm}$ belongs to (2.6) or (2.9). As in the case of $\mathrm{M}_{9}\subset\mathrm{M}_{10}$, there is not a faithful 6-dimensional representation of $\mathrm{M}_{10}$ (see section 7.4). Therefore both of $X_{+}$ and $X_{-}$ belong to the family (2.9).
Let us consider a matrix
$r=\frac{1}{3\sqrt{6}}\begin{bmatrix}2\omega+3&-\omega&-2(\omega+1)&-3\omega-1&\omega-2&-7\\ -\omega-3&-\omega&7(\omega+1)&3\omega+2&1-2\omega&2\\ 2\omega&-4\omega&\omega-2&3\omega+2&\omega-2&3\omega+2\\ -7\omega&-4\omega&1-2\omega&-3\omega-1&1-2\omega&-3\omega-1\\ 2\omega+3&-\omega&\omega-2&-7&-2(\omega+1)&-3\omega-1\\ -\omega-3&-\omega&1-2\omega&2&7(\omega+1)&3\omega+2\end{bmatrix}$
of order 4 with eigenvalues $(1,1,-1,-1,i,-i)$. The group $H\cong 3.\mathrm{A}_{6}$ in section 2.5 is normalized by $r$, and $H$ together with $r$ generate $3.\mathrm{M}_{10}$. By the action of $r$, the space $V_{3}(\rho)=\mathbb{C}f(x)\oplus\mathbb{C}g(x)$ is decomposed into (±)plus-or-minus($\pm$)( ± )-eigenspaces $V_{3}(\rho)_{\pm}=\mathbb{C}F_{\pm}(x)$ where
(8.6) $F_{+}(x)=f(x)-(7+3\sqrt{6})g(x),\qquad F_{-}(x)=f(x)-(7-3\sqrt{6})g(x).$
Cubic fourfolds $X_{\pm}$ defined by $F_{\pm}(x)=0$ are smooth, defined over $\mathbb{Q}(\sqrt{6})$ and $\mathrm{Aut}^{s}(X_{\pm})=\mathrm{M}_{10}$. Since we have
$F_{+}(rx)=F_{+}(x),\qquad F(rx)=-F_{-}(x)$
these are not projectively equivalent, that is, not isomorphic. Moreover $X_{+}$ is $\mathrm{Gal}(\mathbb{Q}(\sqrt{6})/\mathbb{Q})$-conjugate to $X_{-}$.
Appendix A An outer automorphism of $\mathrm{S}_{6}$
It is known that automorphisms of $\mathrm{A}_{n}$ are given by conjugation by elements of $\mathrm{S}_{n}$ and $\mathrm{Aut}(\mathrm{A}_{n})=\mathrm{S}_{n}$ if $n\neq 6$ and $n\geq 4$. However we have $\mathrm{Aut}(\mathrm{A}_{6})=\mathrm{S}_{6}.2=\mathrm{A}_{6}.2^{2}$ and it contains three subgroups of index 2, that are isomorphic to $\mathrm{S}_{6},\ \mathrm{M}_{10}$ and $\mathrm{PGL}_{2}(\mathbb{F}_{9})$ (we have $\mathrm{A}_{6}\cong\mathrm{PSL}_{2}(\mathbb{F}_{9})$). Here we give an outer automorphism of $\mathrm{S}_{6}$ after [JR82] and related facts.
Let us define $\sigma:\mathrm{S}_{6}\rightarrow\mathrm{S}_{6}$ by
$\sigma((12))=(14)(25)(36),\quad\sigma((13))=(16)(24)(35),\quad\sigma((14))=(13)(26)(45),$
$\sigma((15))=(15)(23)(46),\quad\sigma((16))=(12)(34)(56).$
This is a well-defined automorphism of $\mathrm{S}_{6}$ and $\sigma^{2}$ is a conjugation by $(23)(46)$. The automorphism $\sigma$ is an outer automorphism since
  • the conjugacy class of $(12)$ is swiched with that of $(12)(34)(56)$,
  • the conjugacy class of $(123)$ is swiched with that of $(123)(456)$.
We have $\sigma(\mathrm{A}_{6})=\mathrm{A}_{6}$ and $\mathrm{Aut}(\mathrm{A}_{6})/\mathrm{S}_{6}\cong\mathrm{C}_{2}$ is generated by $\sigma$. Moreover we have
$\sigma((12)(45))=(13)(46),\quad\sigma((123))=(123)(465),\quad\sigma((456))=(132)(465)$
and hence $\sigma(\mathrm{A}_{3,3})=\mathrm{A}_{3,3}$ where $\mathrm{A}_{3,3}=(\mathrm{S}_{3}\times\mathrm{S}_{3})\cap\mathrm{A}_{6}$ is generated by
$(12)(45),\qquad(123),\qquad(456).$
Therefore $\sigma$ gives an outer automorphism of $\mathrm{A}_{3,3}$ (see section 4.2). Next let us consider $\mathrm{S}_{5}\subset\mathrm{S}_{6}$ generated by $x=(12345)$ and $y=(12)$. Then $\sigma(\mathrm{S}_{5})$ is generated by
$\sigma(x)=(14325),\quad\sigma(y)=(14)(25)(36)$
and acts on $\{1,2,3,4,5,6\}$ transitively. Therefore $\sigma(\mathrm{S}_{5})$ is not conjugate to a standard $\mathrm{S}_{5}\subset\mathrm{S}_{6}$. Taking intersections with $\mathrm{A}_{6}$, we obtain two types of $\mathrm{A}_{5}$ in $\mathrm{A}_{6}$. For this reason, we have two families of cubic fourfolds with symplectic actions of $\mathrm{A}_{5}$ (see section 6.3).

### v2

@@
-Cubic fourfolds with symplectic automorphisms
+arXiv:2409.08448v2[math.AG] 02 Mar 2026
 Abstract.
@@
 According to them, there are 15 maximal groups, and a finite group $G$ can acts faithfully on a smooth cubic fourfold if and only if $G$ is a subgroup of these 15 groups.
-However, explicit equatoins of cubic fourfolds with automorphisms are still unknown in many cases.
+However, explicit equations of cubic fourfolds with automorphisms are still unknown in many cases.
 In this paper, we study symplectic actions of 34 groups of Laza and Zheng, and detertmine invariant smooth cubic fourfolds.
 Since any automorphism of a smooth cubic fourfold is given by a projective transformation, our problem is reduced to classification of faithfull projective representations of 34 groups which satisfy symplectic conditions.
+We classify such projective representations by considering character tables of Schur covers for 34 groups.
+A Schur cover $G^{*}$ of a finite group $G$ is defined as a central extension $G^{*}=M.G$ such that $M\cong\mathrm{H}^{2}(G,\mathbb{C}^{\times})$ and $M\subset[G^{*},G^{*}]$.
 It is known that complex projective representations of a finite group $G$ lift to linear representations of a Schur cover $G^{*}$.
-We determine possible projective actions of 34 groups $G$ by considering character tables of $G^{*}$.
-As the result, we have
+Therefore, for a fixed $G$, we can obtain all families of cubic fourfolds such that $G\subset\mathrm{Aut}^{s}(X)$ in the following manner.
+(1) Determine linear representations $\rho:G^{*}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ which induces a faithful projective action of $G$ and satisfies symplectic conditions.
+(2) Check that the vector space of $\rho(G)$-invariant cubic forms $V_{3}(\rho)\subset\mathbb{C}[x_{1},\cdots,x_{6}]$ contains $F(x)\in V_{3}(\rho)$ which gives a smooth cubic fourfold $X_{F}$.
+(3) Check that $\mathrm{Aut}^{s}(X_{F})=G$ for a general member $F\in V_{3}(\rho)$.
+For (1), we can determine possible representations from character tables relying on absolute values of characters (see section 2.2).
+We can check (3) by counting dimensions of moduli spaces (see section 2.4).
+For each representation $\rho$ obtained in this manner, $V_{3}(\rho)$ defines an irreducible family of cubic fourfolds such that $G\subset\mathrm{Aut}^{s}(X)$.
+However, we may have $V_{3}(\rho_{1})=V_{3}(\rho_{2})$ even if $\rho_{1}$ is not projectively equivalent to $\rho_{2}$.
+This is caused by the action of $\mathrm{Out}(G)$.
+For an outer automorphism $\sigma$ of $G$, we may have $\rho\neq\rho\circ\sigma$ although we have $V_{3}(\rho)=V_{3}(\rho\circ\sigma)$.
+Excluding duplicated families, we have
 Theorem 1.1.
 The moduli space of cubic fourfolds with $G\subset\mathrm{Aut}^{s}(X)$ has two irreducible components if $G$ is one of the following groups:
-[…]
-For other cases, the moduli spaces are irreducible (see Table 1 in sectoin 2 ).
-For maximal cases, defining eqautions of corresponding cubic fourfolds are known, and they are defined over $\mathbb{Q}$ in most cases. In the last section, we show
+$3,\ \mathrm{A}_{3,3},\ \mathrm{D}_{12},\ \mathrm{A}_{4},\ \mathrm{A}_{4,3},\ \mathrm{A}_{5},\ 3^{2}.4,\ \mathrm{A}_{6},\ \mathrm{S}_{5},\ \mathrm{A}_{7},\ \mathrm{M}_{10}.$
+For other cases, the moduli spaces are irreducible (see Table 2 in Appendix B ).
+For maximal cases, defining equations of corresponding cubic fourfolds are known, and they are defined over $\mathbb{Q}$ in most cases. In the last section, we show
 Theorem 1.2.
@@
 Our results are heavily dependent on computer system GAP. Also we used databases [AtlasWeb] and [GroupNames] for information of finite groups such as generators and automorphisms.
+Acknowledgements.
+The author is grateful to referees for many suggestions to improve the manuscript.
+It was raised by referees that the number of irreducible components is (perhaps more easily) computed by computing the number of primitive embeddings of the lattice $S_{G}=(H^{4}(X,\mathbb{Z})_{prim})_{G}$ into $H^{4}(X,\mathbb{Z})_{prim}$.
+The number of embeddings is included in [MM24], Table 4.
+The groups of Laza and Zheng are included in this list, and each time they appear with a $\times$ in the cubic column corresponds to the number of primitive embeddings of $S_{G}$.
 Notations. We say that $G$ is an extension of $B$ by $A$ and denote $G\cong A.B$ if there is an exact sequence
-[…]
+$1\longrightarrow A\overset{\iota}{\longrightarrow}G\overset{\pi}{\longrightarrow}B\longrightarrow 1$
 (the structure of $A.B$ is not uniquely determined).
@@
   • $\mathrm{C}_{m}$ is often denoted by $m$ (e.g. $4$ denotes $\mathrm{C}_{4}$, and $2^{2}$ denotes $\mathrm{C}_{2}\times\mathrm{C}_{2}$).
-  • $p^{1+2n}_{\pm}$ denote extraspecial groups of order $p^{1+2n}$.
+  • $p^{1+2n}_{\pm}$ denote extraspecial groups of order $p^{1+2n}$ (see Remark 6.1).
   • $\mathrm{S}_{m,n}=\mathrm{S}_{m}\times\mathrm{S}_{n},\ \mathrm{A}_{m,n}=\mathrm{S}_{m,n}\cap\mathrm{A}_{m+n}$, where $\mathrm{A}_{n},\ \mathrm{S}_{n}$ denote alternating, symmetric groups of degree $n$.
@@
 (2.1) $\begin{cases}x_{1}+x_{2}+\cdots+x_{7}=0\\ x_{1}^{3}+x_{2}^{3}+\cdots+x_{7}^{3}=0\end{cases}\text{in}\ \mathbb{P}^{6},$
-and the symmetric group $\mathrm{S}_{7}$ acts on it as permutations of coordinates (However, the actoin of $(12)\in\mathrm{S}_{7}$ is not symplectic).
+and the symmetric group $\mathrm{S}_{7}$ acts on it as permutations of coordinates (However, the action of $(12)\in\mathrm{S}_{7}$ is not symplectic).
 Another $\mathrm{A}_{7}$-invariant cubic fourfold was identified by Yang, Yu and Zhu in [YYZ23] considering the “triple cover” of $\mathrm{A}_{7}$.
@@
 More precisely, a group $G^{*}$ is called a Schur covering group (or representation group) of $G$ if $G^{*}$ is a central extension
-[…]
+(2.4) $1\longrightarrow M\overset{\iota}{\longrightarrow}G^{*}\longrightarrow G\longrightarrow 1$
 such that $M\cong\mathrm{H}^{2}(G,\mathbb{C}^{\times})$ and $\iota(M)\subset[G^{*},G^{*}]$.
@@
 However, it is known that all projective representations of $G$ can be lifted to linear representations of one Schur covering group $G^{*}$.
-An abelian group $M=M(G)$ is called the Schur multiplier of $G$.
+The abelian group $M=M(G)$ in equation (2.4) is called the Schur multiplier of $G$.
 For example, we have
 $M(\mathrm{A}_{n})=\begin{cases}1\quad(n\leq 3)\\ \mathrm{C}_{6}\quad(n=6,7)\\ \mathrm{C}_{2}\quad(\text{others})\end{cases},\quad M(\mathrm{S}_{n})=\begin{cases}1\quad(n\leq 3)\\ \mathrm{C}_{2}\quad(n\geq 4)\end{cases}$
-(see [Ka85]). If $M(G)=1$, we have $G^{*}=G$ and any projective representation of $G$ is lifted to a linear representation of $G$ itself.
+(see [Ka85]).
+If $M(G)=1$, we have $G^{*}=G$ and any projective representation of $G$ is lifted to a linear representation of $G$ itself.
+For non-maximal cases of Laza and Zheng’s 34 groups, Schur covers are given in Appendix B.
 Remark 2.2.
-It is known that $G$ has only one Schur covering group if the orders of $G/[G,G]$ and $M(G)$ are coprime ([Ka85], Theoerm 4.4).
-For non-maximal cases of Laza and Zheng’s 34 groups, we have the following table:
-Table 1. Schur covers and multipliers
-$r(G)$ $G$ GAP-ID $M$ ID* Eq irr
-8 2 2,1 1 2,1 (3.1) 1
-12 $2^{2}$ 4,2 2 8,3 (3.2) 1
-12 3 3,1 1 3,1 (3.3), (3.4) 2
-14 4 4,1 1 4,1 (3.5) 1
-14 $\mathrm{S}_{3}$ 6,1 1 6,1 (3.6) 1
-15 $\mathrm{D}_{8}$ 8,3 2 16,7 (3.7) 1
-16 $\mathrm{A}_{3,3}$ 18,4 3 54,8 (4.6), (4.11) 2
-16 $\mathrm{D}_{12}$ 12,4 2 24,8 (4.2), (4.3) 2
-16 $\mathrm{A}_{4}$ 12,3 2 24,3 (4.20), (4.22) 2
-16 $\mathrm{D}_{10}$ 10, 1 1 10, 1 (4.1) 1
-17 $\mathrm{S}_{4}$ 24,12 2 48,28 (5.2), (5.3) 1
-17 $\mathrm{Q}_{8}$ 8,4 1 8,4 (5.1) 1
-18 $23^{1+4}_{+}:$ 2 486, 249 $3^{5}$ (6.1) 1
-18 $\mathrm{A}_{4,3}$ 72,43 6 432,258 (6.5), (6.7) 2
-18 $\mathrm{A}_{5}$ 60, 5 2 120,5 (6.9), (6.10) 2
-18 $43^{2}.$4 36,9 3 108,15 (6.11), (6.12) 2
-18 $\mathrm{S}_{3,3}$ 36,10 2 72,22 (6.13) 1
-18 $\mathrm{F}_{21}$ 21,1 1 21,1 (6.2) 1
-18 $\mathrm{Hol}(5)$ 20, 3 1 20,3 (6.15) 1
-18 $\mathrm{QD}_{16}$ 16,8 1 16,8 (6.3) 1
-19 $2.23^{1+4}_{+}:$ 2.2 972, 776 $3^{4}$ (7.1) 1
-19 $\mathrm{A}_{6}$ 360,118 6 - (2.6), (2.9) 2
-19 $\mathrm{L}_{2}(7)$ 168,42 2 336,114 (7.2) 1
-19 $\mathrm{S}_{5}$ 120,34 2 240, 89 (7.3), (7.6) 2
-19 $\mathrm{M}_{9}$ 72,41 3 216,88 (7.9) 1
-19 $\mathrm{N}_{72}$ 72,40 2 144,117 (7.10) 1
-19 $\mathrm{T}_{48}$ 48,29 1 48,29 (7.12) 1
-where $r(G)=\mathrm{rank}\ S_{G}$ (see [LZ22]), GAP-ID is “IdGroup($G$)” in GAP, $M$ is the Schur multiplier of $G$, and ID* is GAP-ID of a Schur covering group $G^{*}$ given by “SchurCover($G$)” in GAP.
-Equations of corresponding cubic 4-folds are given in Eq, and irr is the number of irreducible components of the moduli space $\mathcal{M}(G)$ of cubic fourfolds $X$ such that $G\subset\mathrm{Aut}^{s}(X)$.
+It is known that $G$ has only one Schur covering group if the orders of $G/[G,G]$ and $M(G)$ are coprime ([Ka85], Theorem 4.4).
 2.2. Symplectic automorphisms of cubic 4-folds
@@
 According to them, the order of a symplectic automorphism is one of
-$1,2,3,4,5,6,7,8,9,11,12,15$
-If a cubic fourfold $X$ has a symplectic automorphism of oder $n=9,11,12,15$, then we have
+$1,2,3,4,5,6,7,8,9,11,12,15.$
+If a cubic fourfold $X$ has a symplectic automorphism of order $n=9,11,12,15$, then we have
 $\mathrm{Aut}^{s}(X)=3^{4}:\mathrm{A}_{6},\quad\mathrm{L}_{2}(11),\quad 3^{1+4}:2.2^{2},\quad\mathrm{A}_{3,5},$
@@
 $\omega,\ \omega,\ \omega,\ \omega,\ -\omega,\ -\omega,$
-and $g$ is conjugate to $\mathrm{diag}(1,1,1,1,-1,-1)$ in $\mathrm{PGL}_{6}(\mathbb{C})$.
-Also note that, although we can not define the trace of $g\in\mathrm{PGL}_{6}(\mathbb{C})$, the absolute value $|\mathrm{Tr}\,\tilde{g}|$ for a lift $\tilde{g}\in\mathrm{GL}_{6}(\mathbb{C})$ does not depend on a lifting if $\tilde{g}$ is of finite order.
-These values are usefull information to classify symplectic actions from character tables.
+[…]
 2.3. Invariants
@@
 If $\chi$ is trivial, we say that $F(x)$ is a $\rho(G)$-invariant.
-We denote the vectore space of $\rho(G)$-semi-invariants of degree $d$ with character $\chi$, by $V_{d}(\rho,\chi)$.
+We denote the vector space of $\rho(G)$-semi-invariants of degree $d$ with character $\chi$, by $V_{d}(\rho,\chi)$.
 If $\chi$ is trivial, it is denoted by $V_{d}(\rho)$.
@@
 Therefore we have $\widehat{G^{*}}\cong\widehat{G}$.
+(3) We have Molien’s formula (see [DK15])
+$\frac{1}{|G|}\sum_{g\in G}\frac{1}{\det(\mathbf{1}_{n}-t\rho(g))}=\sum_{d=0}^{\infty}(\dim V_{d}(\rho))t^{d},$
+and we can compute $\dim V_{d}(\rho)$ using a function “MolienSeries” in GAP.
 Now let $\rho:G\rightarrow\mathrm{PGL}_{n}(\mathbb{C})$ be a projective representation, and $\tilde{\rho}:G^{*}\rightarrow\mathrm{GL}_{n}(\mathbb{C})$ be a lifting of $\rho$, where $G^{*}$ is a Schur cover with the multiplier $M\subset G^{*}$.
@@
 In particular, $\rho$ and $\rho\otimes\theta$ have same semi-invariants if $\theta^{d}=1$.
-(3)  Let us assume that $d$ and $|\widehat{G}|$ are co-prime to each other. Then, for any $\chi\in\widehat{G}$, there exist $\theta\in\widehat{G}$ such that
+[…]
 $V_{d}(\rho,\,\chi)=V_{d}(\rho\otimes\theta).$
@@
 (3) By the assumption, there is $e\in\mathbb{N}$ such that
-$ed\equiv 1\mod|\widehat{G}|.$
+[…]
 For $\theta=\chi^{-e}$, we have
@@
 For a representation $\rho:G\rightarrow\mathrm{GL}_{n}(\mathbb{C})$, we have the centralizer and the normalizer of $\rho(G)$ in $\mathrm{GL}_{n}(\mathbb{C})$:
-$\mathrm{C}_{\rho}(G)$ =$=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ hgh^{-1}=g\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\},$
-$\mathrm{N}_{\rho}(G)$ =$=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ h\rho(G)h^{-1}=\rho(G)\}.$
+$\mathrm{C}_{\rho}(G)$ $=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ hgh^{-1}=g\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\},$
+$\mathrm{N}_{\rho}(G)$ $=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ h\rho(G)h^{-1}=\rho(G)\}.$
 Moreover we define the stabilizer of $V_{d}(\rho,\,\chi)$:
-$\mathrm{N}_{\rho,\chi}(G)$ =$=\{h\in\mathrm{N}_{\rho}(G)\ |\ \chi_{\rho}(hgh^{-1})=\chi_{\rho}(g)\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\}$
+$\mathrm{N}_{\rho,\chi}(G)$ $=\{h\in\mathrm{N}_{\rho}(G)\ |\ \chi_{\rho}(hgh^{-1})=\chi_{\rho}(g)\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\}$
 where $\chi_{\rho}\circ\rho=\chi$ (see Remark 2.4).
@@
 Then the quotient $V_{d}(\rho,\,\chi)^{sm}/\mathrm{N}_{\rho,\chi}(G)$ is the moduli space $\mathcal{M}_{d}(\rho,\chi)$ of smooth hypersurfaces with the action of $G$ with respect to $(\rho,\,\chi)$.
-Now let $G$ be one of Laza and Zheng’s 34 groups, $G^{*}$ be a Schur cover and $\rho:G^{*}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ be a lifting of a faithful projective actoin of $G$ on $\mathbb{P}^{5}$.
+Now let $G$ be one of Laza and Zheng’s 34 groups, $G^{*}$ be a Schur cover and $\rho:G^{*}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ be a lifting of a faithful projective action of $G$ on $\mathbb{P}^{5}$.
 If $\rho(G^{*})$ acts simplectically on a smooth cubic fourfold $F(x)=0$ with $F\in V_{3}(\rho,\,\chi)$, we have
 (2.5) $20-r(G)=\dim\mathcal{M}_{3}(\rho,\chi)=\dim V_{3}(\rho,\,\chi)-\dim\mathrm{N}_{\rho,\chi}(G^{*}).$
-Note that we have $\dim\mathrm{N}_{\rho,\chi}(G^{*})\geq 1$ since ×$\mathbb{C}^{\times}\subset\mathrm{N}_{\rho,\chi}(G^{*})$. Therefore we have the following criterion to exclude unnecessary cases.
+Note that we have $\dim\mathrm{N}_{\rho,\chi}(G^{*})\geq 1$ since $\mathbb{C}^{\times}\subset\mathrm{N}_{\rho,\chi}(G^{*})$. Therefore we have the following criterion to exclude unnecessary cases.
 Lemma 2.7.
@@
 Let us look at the character table of $3.\mathrm{A}_{6}$.
-$\begin{array}[]{c|rrrrrrrrrrrrrrrrr}\hline\cr\rm class&\rm 1&\rm 3A&\rm 3B&\rm 2A&\rm 6A&\rm 6B&\rm 3C&\rm 3D&\rm 12A&\rm 12B&\rm 4A&\cdots\\ \rm size&1&1&1&45&45&45&120&120&90&90&90&\cdots\\ \hline\cr\chi_{1}&1&1&1&1&1&1&1&1&1&1&1&\cdots\\ \chi_{2}&3&3\overline{\omega}&3\omega&-1&-\overline{\omega}&-\omega&0&0&\overline{\omega}&\omega&1&\\ \chi_{3}&3&3\overline{\omega}&3\omega&-1&-\overline{\omega}&-\omega&0&0&\overline{\omega}&\omega&1&\\ \chi_{4}&3&3\omega&3\overline{\omega}&-1&-\omega&-\overline{\omega}&0&0&\omega&\overline{\omega}&1&\\ \chi_{5}&3&3\omega&3\overline{\omega}&-1&-\omega&-\overline{\omega}&0&0&\omega&\overline{\omega}&1&\\ \chi_{6}&5&5&5&1&1&1&-1&2&-1&-1&-1&\\ \chi_{7}&5&5&5&1&1&1&2&-1&-1&-1&-1&\\ \chi_{8}&6&6\overline{\omega}&6\omega&2&2\overline{\omega}&2\omega&0&0&0&0&0&\\ \chi_{9}&6&6\omega&6\overline{\omega}&2&2\omega&2\overline{\omega}&0&0&0&0&0&\\ \vdots&\vdots&\\ \hline\cr\end{array}$
+[…]
 The following infomations are omitted:
@@
 These characters are afforded by the following two representations.
-One is permutatons of projective coordinates $[x_{1}:x_{2}:x_{3}:x_{4}:x_{5}:x_{6}]$ denoted by $\rho$, and another is the composition of $\rho$ and an outer automorphism $\sigma$ of $\mathrm{A}_{6}$ (see Appendix).
+One is permutatons of projective coordinates $x_{1}:x_{2}:x_{3}:x_{4}:x_{5}:x_{6}]$ denoted by $\rho$, and another is the composition of $\rho$ and an outer automorphism $\sigma$ of $\mathrm{A}_{6}$ (see Appendix A).
 We have $V_{3}(\rho)=V_{3}(\rho\circ\sigma)$ (see Proposition 2.6) and this vector space is given by
 (2.6) $\mathrm{Span}\{x_{1}^{3}+\cdots+x_{6}^{3},\quad(x_{1}^{2}+\cdots+x_{6}^{2})(x_{1}+\cdots+x_{6}),\quad(x_{1}+\cdots+x_{6})^{3}\}.$
-This linear system contains the Fermat cubic, and a general menber is smooth.
+This linear system contains the Fermat cubic, and a general member is smooth.
 Since $\mathrm{A}_{6}$ is a non-commutative simple group, any action is symplectic (Although $\mathrm{S}_{6}$ acts on these cubics, the action of $(12)\in\mathrm{S}_{6}$ is not symplectic since it is diagonalized into $\mathrm{diag}(1,1,1,1,1,-1)$, see Table 1 ).
@@
 (2.8) $g(x)=3(x_{1}x_{2}x_{3}+x_{1}x_{2}x_{4}+x_{1}x_{2}x_{5}+x_{1}x_{2}x_{6}+x_{1}x_{3}x_{4}+x_{1}x_{3}x_{5}+x_{1}x_{5}x_{6}\\ +x_{2}x_{3}x_{4}+x_{2}x_{5}x_{6}+x_{3}x_{5}x_{6}+x_{3}x_{4}x_{6}+x_{3}x_{4}x_{5}+x_{4}x_{5}x_{6})\\ +4(x_{1}x_{4}x_{5}+x_{1}x_{3}x_{6}+x_{2}x_{3}x_{5}+x_{2}x_{4}x_{6})\\ +5(x_{1}x_{4}x_{6}+x_{2}x_{3}x_{6}+x_{2}x_{4}x_{5})+8x_{1}x_{3}x_{5}.$
-This linear system contains a smooth cubic fourfolds $f(x)=0$, and a general menber is smooth. Moreover we have $H=\overline{H}$ since
+This linear system contains a smooth cubic fourfolds $f(x)=0$, and a general member is smooth. Moreover we have $H=\overline{H}$ since
 $\overline{s}=t^{2}st^{-1}(st)^{2}t,\qquad\overline{t}=st(st^{-1})^{2}stst^{-1}s.$
@@
 From the above, we see that $\mathcal{M}(A_{6})$ has two irreducible components.
+2.6. Proof of Theorem 1.1
+As is the case of $\mathrm{A}_{6}$, we determine all admisible projective representations (excluding $\rho$ for which $V_{3}(\rho)$ is same with invariants already obtained) for 34 groups $G$ in the following sections.
+The result is listed in Table 2 in Appendix B.
+To complete the proof of Theorem 1.1, we need to show that two families for groups
+$3,\ \mathrm{A}_{3,3},\ \mathrm{D}_{12},\ \mathrm{A}_{4},\ \mathrm{A}_{4,3},\ \mathrm{A}_{5},\ 3^{2}.4,\ \mathrm{A}_{6},\ \mathrm{S}_{5},\ \mathrm{A}_{7},\ \mathrm{M}_{10}$
+[…]
+[…]
+where horizontal arrows are group actions.
+Therefore if two family $V_{3}(\rho_{1},\chi_{1})$ and $V_{3}(\rho_{2},\chi_{2})$ for $G$ listed in the above are same, then there exist $\phi\in\mathrm{Aut}(G)$ such that $\rho_{1}$ is projectively equivalent to $\rho_{2}\circ\phi$.
+However, we see that this is impossible by considering actions of Schur multiplier or comparing dimensions of irreducible components of two representations.
+For example, we have two families (2.6) and (2.9) for $G=\mathrm{A}_{6}$.
+We can distinguish them from each other since the Schur multiplier $\mathrm{C}_{3}$ acts trvially for (2.6) and non-trivially for (2.9).
+We see also that irreducible decomposition of the action of $\mathrm{A}_{6}$ consists of a 1-dimensional representation and a 5-dimensional representatoin for (2.6) but an irreducible 6-dimensional representation for (2.9).
 3. Groups of small $r(G)$
@@
 (v) $G=\mathrm{S}_{3}$. We have the following normal forms. The symmetric group $\mathrm{S}_{3}$ acts on projective coordinates
-$[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$
+$x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$
 by permuting $x_{i}$ and $y_{i}$ simultaneously. Invariant cubics are
@@
 and the charactet table is as follows.
-$\begin{array}[]{c|rrrrr}\hline\cr\rm class&1&\rm 2A&\rm 2B&\rm 2C&\rm 4A\cr&\rm 1&a^{2}&\{b,ba^{2}\}&\{ba,ba^{3}\}&\{a,a^{3}\}\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&-1&1\cr\chi_{3}&1&1&1&-1&-1\cr\chi_{4}&1&1&-1&1&-1\cr\chi_{5}&2&-2&0&0&0\cr\hline\cr\end{array}.$
+$\begin{array}[]{c|rrrrr}\hline\cr\rm class&1&\rm 2A&\rm 2B\quad&\rm 2C\quad&\rm 4A\quad\cr&\rm 1&a^{2}&\{b,ba^{2}\}&\{ba,ba^{3}\}&\{a,a^{3}\}\cr\hline\cr\chi_{1}&1&1&1\quad&1\quad&1\quad\cr\chi_{2}&1&1&-1\quad&-1\quad&1\quad\cr\chi_{3}&1&1&1\quad&-1\quad&-1\quad\cr\chi_{4}&1&1&-1\quad&1\quad&-1\quad\cr\chi_{5}&2&-2&0\quad&0\quad&0\quad\cr\hline\cr\end{array}.$
 We have $\widehat{\mathrm{D}_{8}}=\{\chi_{1},\chi_{2},\chi_{3},\chi_{4}\}\cong\mathrm{C}_{2}\times\mathrm{C}_{2}$, and we do not have to consider semi-invariants by Proposition 2.6 (3).
@@
 Therefore invariants for $\theta_{3}$ and $\theta_{4}$ are same.
-Let us construct representatoins $\rho_{i}$ affording $\theta_{i}\ (i=1,2,3)$.
+Let us construct representations $\rho_{i}$ affording $\theta_{i}\ (i=1,2,3)$.
 We may assume that $a=\mathrm{diag}(1,1,-1,-1,i,-i)$, and then $\rho_{i}$ is given by $a$ and the following $b_{i}$:
@@
 Moreover, a generic cubic fourfold admitting such an order 6 automorphism has symplectic automorphism group $\mathrm{D}_{12}$.
-The action of $\mathrm{D}_{12}$ is given by the follwoing way. A general member of (4.2) is normalized into
+The action of $\mathrm{D}_{12}$ is given by the following way. A general member of (4.2) is normalized into
 (4.4) $a_{1}(x_{1}^{2}x_{3}+x_{2}^{2}x_{3})+a_{2}(x_{1}^{2}x_{4}+x_{2}^{2}x_{4})+a_{3}(x_{5}^{6}+x_{6}^{6})+\text{other terms}$
-by diagonal actions of (×$(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(-1,-1,1,1,\omega,\omega^{2})$ and a permutation of coordinates by $(12)(56)$.
+by diagonal actions of $(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(-1,-1,1,1,\omega,\omega^{2})$ and a permutation of coordinates by $(12)(56)$.
 A general member of (4.3) is normalized into
 (4.5) $a_{1}(x_{1}^{3}+x_{3}^{3})+a_{2}(x_{1}x_{2}^{2}+x_{3}x_{4}^{2})+\text{other terms}$
-by diagonal actions of (×$(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(1,-1,\omega,-\omega,\omega^{2},\omega^{2})$ and a permutation of coordinates by $(13)(24)$.
+by diagonal actions of $(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(1,-1,\omega,-\omega,\omega^{2},\omega^{2})$ and a permutation of coordinates by $(13)(24)$.
 4.2. The groups $\mathrm{A}_{3,3}=(\mathrm{S}_{3}\times\mathrm{S}_{3})\cap\mathrm{A}_{6}$
@@
 The characters teble of $A_{3,3}$ is as follows.
-$\begin{array}[]{c|rrrrrr}\rm class&1&\rm 2A&\rm 3A&\rm 3B&\rm 3C&\rm 3D\cr\rm size&1&9&2&2&2&2\cr\hline\cr\chi_{1}&1&1&1&1&1&1\cr\chi_{2}&1&-1&1&1&1&1\cr\chi_{3}&2&0&2&-1&-1&-1\cr\chi_{4}&2&0&-1&2&-1&-1\cr\chi_{5}&2&0&-1&-1&2&-1\cr\chi_{6}&2&0&-1&-1&-1&2\end{array}\quad\begin{array}[]{l}\rm 2A:(12)(45)\\ \rm 3A:(123)\\ \rm 3B:(456)\\ \rm 3C:(123)(456)\\ \rm 3D:(123)(465)\\ \end{array}$
+[…]
 Since we have $\widehat{A_{3,3}}\cong C_{2}$, we do not have to consider semi-invariants.
 If $\rho:\mathrm{A}_{3,3}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ is faithful and satisfies the symplectic conditions, we have
-$|\mathrm{Tr}_{\rho}(g)|=\begin{cases}2\quad(\text{the order of $g$ is 2})\\ 0,\,3\quad(\text{the order of $g$ is 3})\end{cases},$
+[…]
 and these conditions are satisfied only by
 $\theta_{ij}=2\chi_{1}+\chi_{i}+\chi_{j},\quad\theta_{ij}\otimes\chi_{2}\qquad(3\leq i<j\leq 6).$
-The outer automorphism group $\mathrm{Out}(A_{3,3})$ is isomorphic to $\mathrm{S}_{4}$, and it permutes classes 3A, 3B, $\rm 3A,\ 3B,\ 3C$ and $\rm 3D$.
+The outer automorphism group $\mathrm{Out}(A_{3,3})$ is isomorphic to $\mathrm{S}_{4}$, and it permutes classes $\rm 3A,\ 3B,\ 3C$ and $\rm 3D$.
 We consider only $\theta_{34}$ and $\theta_{34}\otimes\chi_{2}$ since $\mathrm{Out}(\mathrm{A}_{3,3})$ permutes $\theta_{ij}$ (see Proposition 2.6 (1)).
 Remark 4.1.
-We see that $\mathrm{Out}(A_{3,3})$ is generated by conjugations by $(12),(13)(24)(36)$, and an automorphism $\sigma$ in Appendix.
-A representation $\rho$ affording $\theta_{34}$ is given by permutations of corrdinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is
+We see that $\mathrm{Out}(A_{3,3})$ is generated by conjugations by $(12),(13)(24)(36)$, and an automorphism $\sigma$ in Appendix A.
+A representation $\rho$ affording $\theta_{34}$ is given by permutations of coordinates $x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is
 (4.6) $\mathrm{Span}\{s_{1}(x)^{3},\ s_{2}(x)s_{1}(x),\ s_{3}(x),\ s_{1}(y)^{3},\ s_{2}(y)s_{1}(y),\ s_{3}(y),\\ s_{1}(x)s_{2}(y),\ s_{1}(y)s_{2}(x),\ s_{1}(x)s_{1}(y)^{2},\ s_{1}(x)^{2}s_{1}(y)\},$
-which is a subspace of (3.6). Although these are stable under permutatoins by $\mathrm{S}_{3,3}$, the actoin of $(12)\in\mathrm{S}_{3,3}$ is non-symplectic. We have
+which is a subspace of (3.6). Although these are stable under permutatoins by $\mathrm{S}_{3,3}$, the action of $(12)\in\mathrm{S}_{3,3}$ is non-symplectic. We have
 $\dim\mathrm{C}_{\rho}(\mathrm{A}_{3,3})=2^{2}+1^{2}+1^{2}=6$
@@
 $(\chi_{7}+\chi_{10})(2\mathrm{A})=(\chi_{8}+\chi_{9})(2\mathrm{A})=0,$
-$\chi_{7}+\chi_{10}$ and $\chi_{8}+\chi_{9}$ are not suitable. Let $\rho_{i}$ be a representation affording $\chi_{i}$. We have Molien serieses
+$\chi_{7}+\chi_{10}$ and $\chi_{8}+\chi_{9}$ are not suitable. Let $\rho_{i}$ be a representation affording $\chi_{i}$. We have Molien series
 $\mu(2\rho_{i},\,t)=\sum_{d=0}^{\infty}(\dim V_{d}(2\rho_{i}))t^{d}=1+8t^{3}+\cdots\qquad(i=7,8),$
@@
 and hence $\overline{H}=H$, that is, $V_{3}(2\rho_{7})=V_{3}(2\rho_{8})$. From eigenvalues, we see that matrices
-(4.10) $\tilde{a}=\begin{bmatrix}a&{\bf 0}\\ {\bf 0}&a\end{bmatrix},\quad\tilde{b}=\begin{bmatrix}b&{\bf 0}\\ {\bf 0}&b\end{bmatrix},\quad\tilde{c}=\begin{bmatrix}c&{\bf 0}\\ {\bf 0}&c\end{bmatrix},\quad$
-satisfy symplectic conditions, and $V_{3}(2\rho_{7})$ with projective coordinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ is given by
+(4.10) $\tilde{a}=\begin{bmatrix}a&{\bf 0}\\ {\bf 0}&a\end{bmatrix},\quad\tilde{b}=\begin{bmatrix}b&{\bf 0}\\ {\bf 0}&b\end{bmatrix},\quad\tilde{c}=\begin{bmatrix}c&{\bf 0}\\ {\bf 0}&c\end{bmatrix}$$\quad$
+satisfy symplectic conditions, and $V_{3}(2\rho_{7})$ with projective coordinates $x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ is given by
 (4.11) ${\rm Span}\{s_{3}(x),\ x_{1}x_{2}x_{3},\ s_{3}(y),\ y_{1}y_{2}y_{3},\ I_{1},\ I_{2},\ I_{3},\ I_{4}\}$
@@
 since $\chi_{i}^{3}=1$ for $i=1,2,3$.
-Taking projective coordinates $[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by permutations of $x_{i}$, and it satisfies the symplectic condition.
+Taking projective coordinates $x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by permutations of $x_{i}$, and it satisfies the symplectic condition.
 The linear space $V_{3}(\rho_{1})$ is
@@
 $g(x)$ $=(x_{1}-x_{2})(x_{3}-x_{4})-\omega^{2}(x_{1}-x_{3})(x_{2}-x_{4})+\omega(x_{1}-x_{4})(x_{2}-x_{3}).$
-A general menber of $V_{3}(\rho_{5})$ is smooth since it contains the Fermat cubic. We have
+A general member of $V_{3}(\rho_{5})$ is smooth since it contains the Fermat cubic. We have
 $\dim\mathrm{C}_{\rho_{5}}(\mathrm{A}_{4})=1^{2}+1^{2}+1^{2}+1^{2}=4$
@@
 $\dim V_{3}(\rho_{5},\ \chi_{i})-\dim\mathrm{C}_{\rho_{5}}(\mathrm{A}_{4})=6-4=2\quad(i=2,3),$
-a general menber of $V_{3}(\rho_{5},\chi_{i})$ is not smooth for $i=2,3$.
+a general member of $V_{3}(\rho_{5},\chi_{i})$ is not smooth for $i=2,3$.
 5. Groups of $r(G)=17$
@@
 $\mathrm{Q}_{8}=\{\pm 1,\ \pm I,\ \pm J,\ \pm K\}$
-with usual rules $I^{2}=J^{2}=K^{2}=-1$,  $IJ=-JI=K$. The character table of $\mathrm{Q}_{8}$ is the following:
+with usual rules $I^{2}=J^{2}=K^{2}=-1$, IJ=JI=KIJ=-JI=K. The character table of $\mathrm{Q}_{8}$ is the following:
 $\begin{array}[]{c|rrrrr}\rm class&\rm 1&-1&\pm I&\pm J&\pm K\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&1&-1\cr\chi_{3}&1&1&1&-1&-1\cr\chi_{4}&1&1&-1&-1&1\cr\chi_{5}&2&-2&0&0&0\cr\end{array}.$
-Since we have $\widehat{\mathrm{Q}_{8}}=\mathrm{C}_{2}\times\mathrm{C}_{2}$, we do not have to consider smi-invariants.
+Since we have $\widehat{\mathrm{Q}_{8}}=\mathrm{C}_{2}\times\mathrm{C}_{2}$, we do not have to consider semi-invariants.
 If $\rho:\mathrm{Q}_{8}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic conditions, we have
@@
 The Schur multiplier of $\mathrm{S}_{4}$ is $\mathrm{C}_{2}$, and we do not have to consider a Schur cover. The character tables of $\mathrm{S}_{4}$ is as follows:
-$\begin{array}[]{c|rrrrr}\rm class&1&(12)(34)&(12)&(123)&(1234)\cr\rm size&1&3&6&8&6\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&1&-1\cr\chi_{3}&2&2&0&-1&0\cr\chi_{4}&3&-1&-1&0&1\cr\chi_{5}&3&-1&1&0&-1\cr\end{array}$
-Since we have $\widehat{\mathrm{S}_{4}}=\mathrm{C}_{2}$, we do not have to consider smi-invariants.
+$\begin{array}[]{c|rrrrr}\rm class&1&(12)(34)&(12)&(123)&(1234)\cr\rm size&1&3\quad&6&8\quad&6\quad\cr\hline\cr\chi_{1}&1&1\quad&1&1\quad&1\quad\cr\chi_{2}&1&1\quad&-1&1\quad&-1\quad\cr\chi_{3}&2&2\quad&0&-1\quad&0\quad\cr\chi_{4}&3&-1\quad&-1&0\quad&1\quad\cr\chi_{5}&3&-1\quad&1&0\quad&-1\quad\cr\end{array}$
+Since we have $\widehat{\mathrm{S}_{4}}=\mathrm{C}_{2}$, we do not have to consider semi-invariants.
 If $\rho:\mathrm{S}_{4}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic condition, we have
@@
 $\dim V_{3}(\rho_{3})=6,\qquad\dim V_{3}(\rho_{4})=\dim V_{3}(\rho_{3},\chi_{2})=2,$
-and $\rho_{4}$ is not suitable. Taking projective coordinates $[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by
+and $\rho_{4}$ is not suitable. Taking projective coordinates $x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by
 $(1234)[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]=[x_{2}:x_{3}:x_{4}:x_{1}:y_{1}:-y_{2}],$
@@
 An element of order 3 in $G$ is $\mathrm{diag}(1,1,1,\omega,\omega,\omega)$.
-Moreover, any smooth cubic fourfold with a symplectic automorphism which can be diagonalized as $\mathrm{diag}(1,1,1,\omega,\omega,\omega)$ has this form, and a generic such cubic fourfold has symplectic automorphism group $23^{1+4}_{+}:$ 2.
+Moreover, any smooth cubic fourfold with a symplectic automorphism which can be diagonalized as $\mathrm{diag}(1,1,1,\omega,\omega,\omega)$ has this form, and a generic such cubic fourfold has symplectic automorphism group $3^{1+4}_{+}:2$.
 Remark 6.1.
@@
 The exponent of $p^{1+2n}_{+}$ is $1$, and that of $p^{1+2n}_{-}$ is $2$.
-We give an explicit matrix representation of $23^{1+4}:$ 2 in section 7.1.
+We give an explicit matrix representation of $3^{1+4}_{+}2$ in section 7.1.
 (ii) $G\cong\mathrm{F}_{21}$. For an appropriate choice of coordinates, the defining equations of the corresponding cubic fourfolds are given by
@@
 $\mathrm{QD_{16}}=\left<a,b\ |\ a^{8}=b^{2}=1,\ bab=a^{3}\right$>.
-A general menber in (6.3) is normalized into
+A general member in (6.3) is normalized into
 (6.4) $a_{1}x_{1}^{3}+a_{2}x_{1}^{2}x_{3}+a_{3}(x_{2}x_{3}^{2}+x_{2}x_{4}^{2})+a_{3}x_{1}x_{3}x_{4}\\ +a_{4}(x_{4}x_{5}^{2}+x_{3}x_{6}^{2})+a_{5}x_{2}x_{5}x_{6}$
-by diagonal actions of (×$(\mathbb{C}^{\times})^{6}$. Then $\mathrm{QD}_{16}$ is generated by $a=\mathrm{diag}(1,-1,i,-i,\zeta_{8},\zeta_{8}^{3})$ and a permutation of coordinates by $(34)(56)$.
+by diagonal actions of $(\mathbb{C}^{\times})^{6}$. Then $\mathrm{QD}_{16}$ is generated by $a=\mathrm{diag}(1,-1,i,-i,\zeta_{8},\zeta_{8}^{3})$ and a permutation of coordinates by $(34)(56)$.
 6.2. The group $\mathrm{A}_{4,3}=(\mathrm{S}_{4}\times\mathrm{S}_{3})\cap\mathrm{A}_{7}$
@@
 $\begin{array}[]{c|rrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 3A&\rm 3B&\rm 3C&\rm 3D&\rm 4A&\rm 6A\cr\rm size&1&3&18&2&8&8&8&18&6\cr\hline\cr\chi_{1}&1&1&1&1&1&1&1&1&1\cr\chi_{2}&1&1&-1&1&1&1&1&-1&1\cr\chi_{3}&2&2&0&2&-1&-1&-1&0&2\cr\chi_{4}&2&2&0&-1&2&-1&-1&0&-1\cr\chi_{5}&2&2&0&-1&-1&-1&2&0&-1\cr\chi_{6}&2&2&0&-1&-1&2&-1&0&-1\cr\chi_{7}&3&-1&-1&3&0&0&0&1&-1\cr\chi_{8}&3&-1&1&3&0&0&0&-1&-1\cr\chi_{9}&6&-2&0&-3&0&0&0&0&1\cr\end{array}\quad\begin{array}[]{l}\rm 2A:(12)(34)\\ \rm 2B:(12(56)\\ \rm 3A:(567)\\ \rm 3B:(123)\\ \rm 3C:(123)(567)\\ \rm 3D:(123)(576)\\ \rm 4A:(1234)(67)\\ \rm 6A:(12)(34)(567)\end{array}$
-Since we have $\widehat{\mathrm{A}_{4,3}}=\mathrm{C}_{2}$, we do not have to consider smi-invariants. A character $\phi$ satisfying conditions
+Since we have $\widehat{\mathrm{A}_{4,3}}=\mathrm{C}_{2}$, we do not have to consider semi-invariants. A character $\phi$ satisfying conditions
 $|\phi(2A)|=|\phi(2B)|=2,\qquad|\phi(4A)|=0$
@@
 One of them is given by
-$[(3,4)(6,7),\ (5,6,7),\ (2,3,4),\ (1,4)(2,3),\ (1,3)(2,4)]\\ \longrightarrow[(2,4)(5,6),\ (5,6,7),\ (2,4,3)(5,6,7),\ (1,4)(2,3),\ (1,2)(3,4)],$
+[(3,4)(6,7),(5,6,7),(2,3,4),(1,4)(2,3),(1,3)(2,4)][(2,4)(5,6),(5,6,7),(2,4,3)(5,6,7),(1,4)(2,3),(1,2)(3,4)],[(3,4)(6,7),\ (5,6,7),\ (2,3,4),\ (1,4)(2,3),\ (1,3)(2,4)]\\ $\longrightarrow$[(2,4)(5,6),\ (5,6,7),\ (2,4,3)(5,6,7),\ (1,4)(2,3),\ (1,2)(3,4)],
 which is obtained by GAP.
@@
 The representaion $\rho$ is given as follows.
-Let $[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}:y_{3}]$ be projective coordinates of $\mathbb{P}^{6}$, and $H_{0}$ be a hyperplane
+Let $x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}:y_{3}]$ be projective coordinates of $\mathbb{P}^{6}$, and $H_{0}$ be a hyperplane
 $s_{1}(x)+s_{1}(y)=(x_{1}+x_{2}+x_{3}+x_{4})+(y_{1}+y_{2}+y_{3})=0$
@@
 $\begin{array}[]{c|rrrrr}\rm class&\rm 1&(12)(34)&(123)&(12345)&(13524)\cr\rm size&1&15&20&12&12\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&3&-1&0&\frac{1+\sqrt{5}}{2}&\frac{1-\sqrt{5}}{2}\cr\chi_{3}&3&-1&0&\frac{1-\sqrt{5}}{2}&\frac{1+\sqrt{5}}{2}\cr\chi_{4}&4&0&1&-1&-1\cr\chi_{5}&5&1&-1&0&0\cr\end{array}$
-Since we have $\widehat{\mathrm{A}_{5}}=1$, we do not have to consider smi-invariants. If $\rho:\mathrm{A}_{5}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic condition, we have
+Since we have $\widehat{\mathrm{A}_{5}}=1$, we do not have to consider semi-invariants. If $\rho:\mathrm{A}_{5}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic condition, we have
 $|\mathrm{Tr}_{\rho}(12)(34)|=2,\quad|\mathrm{Tr}_{\rho}(123)|=0,3,\quad|\mathrm{Tr}_{\rho}(12345)|=1.$
@@
 $\mu(\rho_{3},t)$ $=1+2t^{2}+6t^{4}+\cdots,$
-and $V_{3}(\rho_{3})=0$. For projective coordinates $[x_{1}:\cdots:x_{5}:y]$, the representation $\rho_{1}$ is given as permutations of $x_{i}$, and we have
+and $V_{3}(\rho_{3})=0$. For projective coordinates $x_{1}:\cdots:x_{5}:y]$, the representation $\rho_{1}$ is given as permutations of $x_{i}$, and we have
 (6.9) $V_{3}(\rho_{1})=\mathrm{Span}\{y^{3},\ y^{2}s_{1}(x),\ ys_{1}(x)^{2},\ ys_{2}(x),\ s_{1}(s)^{3},\ s_{1}(x)s_{2}(x),\ s_{3}(x)\}$
-Although $F\in V_{3}(\rho_{1})$ is stable under permutatoins by $\mathrm{S}_{5}$, the actoin of $(12)\in\mathrm{S}_{5}$ is non-symplectic.
+Although $F\in V_{3}(\rho_{1})$ is stable under permutatoins by $\mathrm{S}_{5}$, the action of $(12)\in\mathrm{S}_{5}$ is non-symplectic.
 Replacing $x_{5}$ and $y$ by $y_{1}$ and $y_{2}$, we see that this is a subspace of (4.20).
@@
 $\dim V_{3}(\rho_{1})-\dim\mathrm{C}_{\rho_{1}}(\mathrm{A}_{5})=2=20-r(\mathrm{A}_{5}).$
-The representation $\rho_{2}$ is given by permutations of projective coordinates $[x_{1}:\cdots:x_{6}]$ by “non-standard” $\mathrm{A}_{5}\subset\mathrm{A}_{6}$ generated by $(12345)$ and $(35)(46)$ (see Appendix).
+The representation $\rho_{2}$ is given by permutations of projective coordinates $x_{1}:\cdots:x_{6}]$ by “non-standard” $\mathrm{A}_{5}\subset\mathrm{A}_{6}$ generated by $(12345)$ and $(35)(46)$ (see Appendix A).
 Putting $x_{ijk}=x_{i}x_{j}x_{k}$ and
@@
 6.4. The group $\mathbf{G\cong 3^{2}.4}$ with GAP-ID (36,9)
-This group $G$ is represented as a (maximal) subgroup of $\mathrm{A}_{6}$ generated by $(123),\ (456),\ (1425)(36)$, and containing $\mathrm{A}_{3,3}=3^{2}:2$ as a subgroup of index 2.
+This group $G$ is represented as a (maximal) subgroup of $\mathrm{A}_{6}$ generated by (123),(456),(1425)(36)(123),\ (456),\ (1425)(36), and containing $\mathrm{A}_{3,3}=3^{2}:2$ as a subgroup of index 2.
 We have a Schur cover $G^{*}\cong 3.G$ with GAP-ID (108,15).
@@
 for $\chi_{i}$ and $\theta_{34}$ in section 4.2.
-A representaion $\rho$ affording $\phi_{1}$ is given by permutatioins of coordinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is $(1425)(36)$-invariants of (4.6):
+A representaion $\rho$ affording $\phi_{1}$ is given by permutatioins of coordinates $x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is $(1425)(36)$-invariants of (4.6):
 (6.11) $\mathrm{Span}\{s_{1}(x)^{3}+s_{1}(y)^{3},\quad s_{2}(x)s_{1}(x)+s_{2}(y)s_{1}(y),\quad s_{3}(x)+s_{3}(y),\\ s_{1}(x)s_{2}(y)+s_{1}(y)s_{2}(x),\quad s_{1}(x)s_{1}(y)^{2}+s_{1}(x)^{2}s_{1}(y)\}.$
@@
 $\eta_{7}+\eta_{8},\quad\eta_{9}+\eta_{10},\quad\eta_{11}+\eta_{12},\quad\eta_{13}+\eta_{14}.$
-For representations $\rho_{i}$ affording $\eta_{i}$, we have Molien serieses
+For representations $\rho_{i}$ affording $\eta_{i}$, we have Molien series
 $\mu(\rho_{i}+\rho_{j},\,t)=\sum_{d=0}^{\infty}(\dim V_{d}(\rho_{i}+\rho_{j}))t^{d}=1+4t^{3}+\cdots\qquad((i,j)=(7,8),(11,12))$
@@
 $h=\frac{1}{\sqrt{3}}\begin{bmatrix}1&1&1\\ 1&\omega&\omega^{2}\\ 1&\omega^{2}&\omega\end{bmatrix}$.
-[…]
+Eigenvalues of $h$ are 1,1,i1,-1,i, and hence $\tilde{h}$ satisfies the symplectic condition.
+The group $H\cong 3.A_{3,3}$ generated by $\tilde{a},\tilde{b},\tilde{c}$ in (4.10) is normalized by $\tilde{h}$, and $H$ together with $\tilde{h}$ generate a group $H_{2}\cong 3.(3^{2}.4)$ with GAP-ID (108,15) (we have $\tilde{h}^{2}=\tilde{a}$ and $H_{2}:H]=2$).
+This gives one of $\rho_{7}+\rho_{8}$ and $\rho_{11}+\rho_{13}$.
+Moreover we have $\overline{H}=H$ (see (4.9)) and $\overline{h}=h^{3}$, that is, $\overline{H_{2}}=H_{2}$.
+Therefore we have
 $V_{3}(\rho_{7}+\rho_{8})=V_{3}(\overline{\rho_{7}}+\overline{\rho_{8}})=V_{3}(\rho_{11}+\rho_{13})$
@@
 and the character table of $\mathrm{S}_{3,3}$ is as follows.
-$\begin{array}[]{c|rrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 2C&\rm 3A&\rm 3B&\rm 3C&\rm 6A&\rm 6B\cr\rm size&1&3&3&9&2&2&4&6&6\cr\hline\cr\eta_{11}&1&1&1&1&1&1&1&1&1\cr\eta_{21}&1&-1&1&-1&1&1&1&1&-1\cr\eta_{12}&1&1&-1&-1&1&1&1&-1&1\cr\eta_{22}&1&-1&-1&1&1&1&1&-1&-1\cr\eta_{13}&2&2&0&0&2&-1&-1&0&-1\cr\eta_{23}&2&-2&0&0&2&-1&-1&0&1\cr\eta_{31}&2&0&2&0&-1&2&-1&-1&0\cr\eta_{32}&2&0&-2&0&-1&2&-1&1&0\cr\eta_{33}&4&0&0&0&-2&-2&1&0&0\cr\end{array}\qquad\begin{array}[]{l}\rm 2A=(12)\\ \rm 2B=(45)\\ \rm 2C=(12)(45)\\ \rm 3A=(123)\\ \rm 3B=(456)\\ \rm 3C=(123)(456)\\ \rm 6A=(123)(45)\\ \rm 6B=(12)(456)\\ \end{array}$
+[…]
 Since we have $\widehat{\mathrm{S}_{3,3}}\cong C_{2}\times\mathrm{C}_{2}$,we consider only invariants with the trivial character.
@@
 $\phi|\mathrm{A}_{3,3}=\theta_{34}=2\chi_{1}+\chi_{3}+\chi_{4}$
-is one of the follwoings:
+is one of the followings:
 $\phi_{1}=\eta_{11}+\eta_{22}+\eta_{13}+\eta_{31},\quad\phi_{1}\otimes\eta_{22}=\eta_{11}+\eta_{22}+\eta_{23}+\eta_{32},$
 $\phi_{2}=\eta_{11}+\eta_{22}+\eta_{13}+\eta_{32},\quad\phi_{2}\otimes\eta_{22}=\eta_{11}+\eta_{22}+\eta_{23}+\eta_{31}.$
-Let us consider a representation $\rho_{1}$ affording $\phi_{1}$. Permutations of coordinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ by $\mathrm{A}_{3,3}$ is normalized by an action of a matrix
+Let us consider a representation $\rho_{1}$ affording $\phi_{1}$. Permutations of coordinates $x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ by $\mathrm{A}_{3,3}$ is normalized by an action of a matrix
 $b_{1}=\frac{1}{3}\begin{bmatrix}-2&1&-2\\ 1&-2&-2\\ -2&-2&1\end{bmatrix}\oplus\mathrm{diag}(1,1,1)$
-with eigenvalues $(-1,-1,1,1,1,1)$.
+with eigenvalues (1,1,1,1,1,1)(-1,-1,1,1,1,1).
 These actions generate $\mathrm{S}_{3,3}$ and this gives $\rho_{1}$.
@@
 (6.13) $V_{3}(\rho_{1})=\mathrm{Span}\{s_{1}(y)^{3},\ s_{2}(y)s_{1}(y),\ s_{3}(y),\ s_{1}(x)^{2}s_{1}(y),\ s_{2}(x)s_{1}(y),\ p(x)\}$
-where $p(x)=2s_{1}(x)^{3}-9s_{1}(x)s_{2}(x)+9s_{3}(x)$, and the $(-1)$-eigenspace
+where $p(x)=2s_{1}(x)^{3}-9s_{1}(x)s_{2}(x)+9s_{3}(x)$, and the (1)(-1)-eigenspace
 (6.14) $V_{3}(\rho_{1},\eta_{22})=V_{3}(\rho_{1}\otimes\eta_{22})\\ =\mathrm{Span}\{s_{1}(x)^{3},\ s_{2}(x)s_{1}(x),\ s_{1}(x)s_{2}(y),\ s_{1}(x)s_{1}(y)^{2}\}.$
@@
 $V_{3}(\rho_{2})=\mathrm{Span}\{s_{1}(y)^{3},\ s_{2}(y)s_{1}(y),\ s_{3}(y),\ s_{1}(y)s_{2}(x),\ s_{1}(x)^{2}s_{1}(y)\},$
-and $(-1)$-eigenspace of $b_{2}$
+and (1)(-1)-eigenspace of $b_{2}$
 $V_{3}(\rho_{2},\eta_{22})=V_{3}(\rho_{2}\otimes\eta_{22})\\ =\mathrm{Span}\{s_{1}(x)^{3},\ s_{2}(x)s_{1}(x),\ s_{3}(x),\ s_{1}(x)s_{2}(y),\ s_{1}(x)s_{1}(y)^{2}\}.$
@@
 6.6. The group $\mathbf{Hol}(5)$
-[…]
-$\begin{array}[]{c|rrrrr}\rm class&\rm 1&(12)(34)&(2345)&(2534)&(12354)\cr\rm size&1&5&5&5&4\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&-1&1\cr\chi_{3}&1&-1&-i&i&1\cr\chi_{4}&1&-1&i&-i&1\cr\chi_{5}&4&0&0&0&-1\cr\end{array}$
+The group $\mathrm{Hol}(5)\cong\mathrm{D}_{10}.2$ is given as a subgroup of $\mathrm{S}_{5}$ generated by $(12)(34)$ and $(2,3,4,5)$.
+The Schur multiplier of $\mathrm{Hol}(5)$ is 1, and hence we have $\mathrm{Hol}(5)^{*}=\mathrm{Hol}(5)$.
+The character table of $\mathrm{Hol}(5)$ is as follows:
+$\begin{array}[]{c|rrrrr}\rm class&\rm 1&(12)(34)&(2345)&(2534)&(12354)\cr\rm size&1&5\quad&5\quad&5\quad&4\quad\cr\hline\cr\chi_{1}&1&1\quad&1\quad&1\quad&1\quad\cr\chi_{2}&1&1\quad&-1\quad&-1\quad&1\quad\cr\chi_{3}&1&-1\quad&-i\quad&i\quad&1\quad\cr\chi_{4}&1&-1\quad&i\quad&-i\quad&1\quad\cr\chi_{5}&4&0\quad&0\quad&0\quad&-1\quad\cr\end{array}$
 Since $\widehat{\mathrm{Hol}(5)}\cong\mathrm{C}_{4}$, we consider only invariants with the trivial character. Faithful characters satisfying the condition on absolute values are
@@
 $g=\begin{bmatrix}\mathbf{0}&B\\ \mathbf{1}&\mathbf{0}\end{bmatrix}\quad\text{where}\quad B=\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{bmatrix}$
-whose eigenvalues are $(1,1,-1,-1,i,-i)$.
+whose eigenvalues are (1,1,1,1,i,i)(1,1,-1,-1,i,-i).
 7.2. The projective special linear group $\mathrm{L}_{2}(7)$
@@
 (7.2) $\begin{cases}x_{1}+x_{2}+x_{3}+x_{4}+x_{5}+x_{6}+x_{7}=0\\ c_{1}(x_{1}^{3}+x_{2}^{3}+x_{3}^{3}+x_{4}^{3}+x_{5}^{3}+x_{6}^{3}+x_{7}^{3})\\ \qquad\qquad+c_{2}(x_{124}+x_{235}+x_{346}+x_{156}+x_{137}+x_{457}+x_{267})=0\end{cases}$
-in $\mathbb{P}^{6}$ where $x_{ijk}=x_{i}x_{j}x_{k}$. This family contains the Clebsch-Segre cubic (2.1), and a general menber is smooth.
+in $\mathbb{P}^{6}$ where $x_{ijk}=x_{i}x_{j}x_{k}$. This family contains the Clebsch-Segre cubic (2.1), and a general member is smooth.
 7.3. The symmetric group $\mathrm{S}_{5}$
@@
 Since we have $\widehat{\mathrm{S}_{5}}=\mathrm{C}_{2}$, we consider only invariants with the trivial character. Faithful characters satisfying the conditions on absolute values are
-$\phi_{1}=\eta_{1}+\eta_{2}+\eta_{3},\quad\phi_{1}\otimes\eta_{2}=\eta_{1}+\eta_{2}+\eta_{4}.,$
+$\phi_{1}=\eta_{1}+\eta_{2}+\eta_{4},\quad\phi_{1}\otimes\eta_{2}=\eta_{1}+\eta_{2}+\eta_{3},$
 $\phi_{2}=\eta_{1}+\eta_{5},\quad\phi_{2}\otimes\eta_{2}=\eta_{2}+\eta_{6}.$
@@
 $\pi(\sigma)[x_{1}\cdots:x_{5}:y]=[x_{\sigma(1)}:\cdots:x_{\sigma(5)}:(\mathrm{sgn}\,\sigma)y],$
-and this satisfies the symplectic conditions. We have $\pi_{1}|\mathrm{A}_{5}=\rho_{1}$ and $V_{3}(\rho_{1})$ is decomposed into (±$(\pm 1)$-eigenspaces for $(12)\in\mathrm{S}_{5}$
+and this satisfies the symplectic conditions. We have $\pi_{1}|\mathrm{A}_{5}=\rho_{1}$ and $V_{3}(\rho_{1})$ is decomposed into $(\pm 1)$-eigenspaces for $(12)\in\mathrm{S}_{5}$
 (7.3) $V_{3}(\rho_{1})_{+}$ $=V_{3}(\pi_{1})=\mathrm{Span}\{y^{2}s_{1}(x),\quad s_{1}(x)^{3},\quad s_{1}(x)s_{2}(x),\quad s_{3}(x)\}$
@@
 (7.4) $V_{3}(\rho_{1})_{-}$ $=V_{3}(\pi_{1}\otimes\eta_{2})=\mathrm{Span}\{y^{3},\quad ys_{1}(x)^{2},\quad ys_{2}(x)\}.$
-A general menber of $V_{3}(\pi_{1})$ defines a smooth cubic 4-fold since $y^{2}s_{1}(x)+s_{3}(x)=0$ is smooth, but $F\in V_{3}(\pi_{1}\otimes\eta_{2})$ is reducible.
+A general member of $V_{3}(\pi_{1})$ defines a smooth cubic 4-fold since $y^{2}s_{1}(x)+s_{3}(x)=0$ is smooth, but $F\in V_{3}(\pi_{1}\otimes\eta_{2})$ is reducible.
 We have $\dim\mathrm{C}_{\pi_{1}}(\mathrm{S}_{5})=1^{2}+1^{2}+1^{2}=3$ and
 $\dim V_{3}(\pi_{1})-\dim\mathrm{C}_{\pi_{1}}(\mathrm{A}_{5})=1=20-r(\mathrm{S}_{5}).$
-Next we consider a representation $\pi_{2}$ affording $\phi_{2}$. Let us consider the following involution with eigenvalues $(1,1,1,1,-1,-1)$:
+Next we consider a representation $\pi_{2}$ affording $\phi_{2}$. Let us consider the following involution with eigenvalues (1,1,1,1,1,1)(1,1,1,1,-1,-1):
 (7.5) $a=\frac{1}{3}[a_{ij}]_{6\times 6}\qquad a_{ij}=\begin{cases}-2\ (i+j\equiv 3\mod 6)\\ 1\ (i+j\not\equiv 3\mod 6)\end{cases}.$
@@
 (7.7) $V_{3}(\rho_{2})_{-}$ $=V_{3}(\pi_{2}\otimes\eta_{2})=\mathrm{Span}\{s_{1}(x)^{3}-9s_{1}(x)s_{2}(x)+18s_{3}(x)\}.$
-A general menber of $V_{3}(\pi_{2})$ defines a smooth cubic 4-fold since $s_{1}(x)s_{2}(x)+(I_{1}-I_{2})=0$ is smooth. On the other hand,
+A general member of $V_{3}(\pi_{2})$ defines a smooth cubic 4-fold since $s_{1}(x)s_{2}(x)+(I_{1}-I_{2})=0$ is smooth. On the other hand,
 $s_{1}(x)^{3}-9s_{1}(x)s_{2}(x)+18s_{3}(x)=0$
-is singular at $[1:\cdots:1]$. We have $\dim\mathrm{C}_{\pi_{2}}(\mathrm{S}_{5})=1^{2}+1^{2}=2$ and
+is singular at $1:\cdots:1]$. We have $\dim\mathrm{C}_{\pi_{2}}(\mathrm{S}_{5})=1^{2}+1^{2}=2$ and
 $\dim V_{3}(\pi_{2})-\dim\mathrm{C}_{\pi_{2}}(\mathrm{A}_{5})=1=20-r(\mathrm{S}_{5}).$
 7.4. The Mathieu group $\mathrm{M}_{9}$
-We have $\mathrm{M}_{9}\cong 3^{2}.\mathrm{Q}_{8}$ and this group contains $43^{2}.$4 in section 6.4 as a subgroup of index 2.
+We have $\mathrm{M}_{9}\cong 3^{2}.\mathrm{Q}_{8}$ and this group contains $3^{2}.4$ in section 6.4 as a subgroup of index 2.
 The Schur multiplier of $\mathrm{M}_{9}$ is $\mathrm{C}_{3}$, and we must consider a Schur cover $3.\mathrm{M}_{9}$.
@@
 $\chi({\rm 4A})=\chi({\rm 4B})=\chi({\rm 4C})=0$
-[…]
+are satisfied only by $\chi_{15}$ and $\chi_{16}=\overline{\chi_{15}}$.
+Representaions affording $\chi_{15}$ and $\overline{\chi_{15}}$ are given by extending the action of $H_{2}\cong 3.(3^{2}.4)$ in section 6.4.
+Let us consider a matrix $g=\begin{bmatrix}{\bf 0}&h_{2}\\ \omega h_{2}&{\bf 0}\end{bmatrix}$ of order 4 with eigenvalues (1,1,1,1,i,i)(1,1,-1,-1,i,-i) where
 (7.8) $h_{2}=\frac{1}{\sqrt{3}}\begin{bmatrix}\omega&\omega^{2}&\omega^{2}\\ 1&1&\omega^{2}\\ 1&\omega^{2}&1\end{bmatrix}$.
@@
 The Schur multiplier of $\mathrm{N}_{72}$ is $\mathrm{C}_{2}$, and we do not have to consider a Schur cover. The character table of $\mathrm{N}_{72}$ is as follows.
-$\begin{array}[]{c|rrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 2C&\rm 3A&\rm 3B&\rm 4A&\rm 6A&\rm 6B\cr\rm size&1&6&6&9&4&4&18&12&12\cr\hline\cr\xi_{1}&1&1&1&1&1&1&1&1&1\cr\xi_{2}&1&-1&-1&1&1&1&1&-1&-1\cr\xi_{3}&1&-1&1&1&1&1&-1&-1&1\cr\xi_{4}&1&1&-1&1&1&1&-1&1&-1\cr\xi_{5}&2&0&0&-2&2&2&0&0&0\cr\xi_{6}&4&-2&0&0&1&-2&0&1&0\cr\xi_{7}&4&0&-2&0&-2&1&0&0&1\cr\xi_{8}&4&0&2&0&-2&1&0&0&-1\cr\xi_{9}&4&2&0&0&1&-2&0&-1&0\cr\end{array}\qquad\begin{array}[]{l}\rm 2A=(12)\\ \rm 2B=(14)(25)(36)\\ \rm 2C=(12)(45)\\ \rm 3A=(123)\\ \rm 3B=(123)(456)\\ \rm 4A=(14)(2536)\\ \rm 6A=(12)(456)\\ \rm 6B=(142536)\\ \end{array}$
+[…]
 Since we have $\widehat{\mathrm{N}_{72}}\cong C_{2}\times\mathrm{C}_{2}$,we consider only invariants with the trivial character.
@@
 A representation $\pi$ affording the character $\psi$ is given by extending $\rho_{1}$ in secton 6.5.
-Let us consider a matrix $c=\begin{bmatrix}C_{1}&C_{2}\\ C_{2}&C_{1}\end{bmatrix}$ of order 2 with eigenvalues $(1,1,1,1,-1,-1)$, where
+Let us consider a matrix $c=\begin{bmatrix}C_{1}&C_{2}\\ C_{2}&C_{1}\end{bmatrix}$ of order 2 with eigenvalues (1,1,1,1,1,1)(1,1,1,1,-1,-1), where
 $C_{1}=\begin{bmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{bmatrix},\quad C_{2}=\begin{bmatrix}-2&1&1\\ 1&-2&1\\ 1&1&-2\end{bmatrix}$.
@@
 (7.10) $V_{3}(\pi)=\mathrm{Span}\{s_{1}(y)^{3},\ s_{1}(x)^{2}s_{1}(y),\ s_{2}(y)s_{1}(y)+s_{2}(x)s_{1}(y),\\ 9s_{2}(y)s_{1}(y)-9s_{3}(y)+p(x)\}$
-and the $(-1)$-eigenspace
+and the (1)(-1)-eigenspace
 (7.11) $V_{3}(\pi,\xi_{2})=V_{3}(\pi\otimes\xi_{2})=\mathrm{Span}\{2s_{1}(y)^{3}-9s_{2}(y)s_{1}(y)-9s_{3}(y)+p(x),\\ s_{1}(y)^{3}-3s_{2}(y)s_{1}(y)-s_{1}(x)^{2}s_{1}(y)+3s_{2}(x)s_{1}(y)\}.$
@@
 $n_{1}=\begin{bmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{bmatrix}\oplus\frac{1}{\sqrt{2}}\begin{bmatrix}1&-i\\ i&-1\end{bmatrix},\quad n_{2}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\oplus\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\oplus\begin{bmatrix}0&\zeta_{8}^{3}\\ \zeta_{8}^{5}&0\end{bmatrix}$
-of order 2 with eigenvalues $(1,1,1,1,-1,-1)$.
+of order 2 with eigenvalues (1,1,1,1,1,1)(1,1,1,1,-1,-1).
 They normalize $\rho(\mathrm{Q}_{8})$, and $n_{1},n_{2}$ together with $\rho(\mathrm{Q}_{8})$ generate a group $H\cong\mathrm{Q}_{8}:\mathrm{S}_{3}$ with GAP-ID (48,29) (Conjugations by $n_{1}$ and $n_{2}$ permute conjugacy classes
@@
 (7.12) $\mathrm{Span}\{x_{1}^{3},\quad x_{1}(x_{2}^{2}+x_{3}^{2}+x_{4}^{2}),\quad x_{2}x_{3}x_{4},\\ ix_{3}(x_{5}^{2}+x_{6}^{2})+2x_{2}x_{5}x_{6}+x_{4}(x_{5}^{2}-x_{6}^{2})\}.$
-A general menber of $V_{3}(\pi)$ is smooth since sum of the above four generator define a smooth cubic fourfold. We have $\dim\mathrm{C}_{\pi}(\mathrm{T}_{48})=1^{2}+1^{2}+1^{2}=3$ and
+A general member of $V_{3}(\pi)$ is smooth since sum of the above four generator define a smooth cubic fourfold. We have $\dim\mathrm{C}_{\pi}(\mathrm{T}_{48})=1^{2}+1^{2}+1^{2}=3$ and
 $\dim V_{3}(\pi)-\dim\mathrm{C}_{\pi}(\mathrm{T}_{48})=1=20-r(\mathrm{T}_{48}).$
@@
 $r=\frac{1}{3\sqrt{6}}\begin{bmatrix}2\omega+3&-\omega&-2(\omega+1)&-3\omega-1&\omega-2&-7\\ -\omega-3&-\omega&7(\omega+1)&3\omega+2&1-2\omega&2\\ 2\omega&-4\omega&\omega-2&3\omega+2&\omega-2&3\omega+2\\ -7\omega&-4\omega&1-2\omega&-3\omega-1&1-2\omega&-3\omega-1\\ 2\omega+3&-\omega&\omega-2&-7&-2(\omega+1)&-3\omega-1\\ -\omega-3&-\omega&1-2\omega&2&7(\omega+1)&3\omega+2\end{bmatrix}$
-of order 4 with eigenvalues $(1,1,-1,-1,i,-i)$.
+of order 4 with eigenvalues (1,1,1,1,i,i)(1,1,-1,-1,i,-i).
 The group $H\cong 3.\mathrm{A}_{6}$ in section 2.5 is normalized by $r$, and $H$ together with $r$ generate $3.\mathrm{M}_{10}$.
-By the action of $r$, the space $V_{3}(\rho)=\mathbb{C}f(x)\oplus\mathbb{C}g(x)$ is decomposed into (±)plus-or-minus($\pm$)( ± )-eigenspaces $V_{3}(\rho)_{\pm}=\mathbb{C}F_{\pm}(x)$ where
+By the action of $r$, the space $V_{3}(\rho)=\mathbb{C}f(x)\oplus\mathbb{C}g(x)$ is decomposed into $(\pm)$-eigenspaces $V_{3}(\rho)_{\pm}=\mathbb{C}F_{\pm}(x)$ where
 (8.6) $F_{+}(x)=f(x)-(7+3\sqrt{6})g(x),\qquad F_{-}(x)=f(x)-(7-3\sqrt{6})g(x).$
@@
 For this reason, we have two families of cubic fourfolds with symplectic actions of $\mathrm{A}_{5}$ (see section 6.3).
+Appendix B Table
+For non-maximal cases of Laza and Zheng’s 34 groups, we have the following table, where $r(G)=\mathrm{rank}\ S_{G}$ (see [LZ22]), GAP-ID is “IdGroup(GG)” in GAP, $M$ is the Schur multiplier of $G$, and ID* is GAP-ID of a Schur covering group $G^{*}$ given by “SchurCover(GG)” in GAP.
+Equations of corresponding cubic 4-folds are given in Eq, and irr is the number of irreducible components of the moduli space $\mathcal{M}(G)$ of cubic fourfolds $X$ such that $G\subset\mathrm{Aut}^{s}(X)$.
+Table 2. Schur covers and multipliers
+$r(G)$ $G$ GAP-ID $M$ ID* Eq irr
+8 2 2,1 1 2,1 (3.1) 1
+12 $2^{2}$ 4,2 2 8,3 (3.2) 1
+12 3 3,1 1 3,1 (3.3), (3.4) 2
+14 4 4,1 1 4,1 (3.5) 1
+14 $\mathrm{S}_{3}$ 6,1 1 6,1 (3.6) 1
+15 $\mathrm{D}_{8}$ 8,3 2 16,7 (3.7) 1
+16 $\mathrm{A}_{3,3}$ 18,4 3 54,8 (4.6), (4.11) 2
+16 $\mathrm{D}_{12}$ 12,4 2 24,8 (4.2), (4.3) 2
+16 $\mathrm{A}_{4}$ 12,3 2 24,3 (4.20), (4.22) 2
+16 $\mathrm{D}_{10}$ 10, 1 1 10, 1 (4.1) 1
+17 $\mathrm{S}_{4}$ 24,12 2 48,28 (5.2), (5.3) 1
+17 $\mathrm{Q}_{8}$ 8,4 1 8,4 (5.1) 1
+18 $3^{1+4}_{+}:2$ 486, 249 $3^{5}$ - (6.1) 1
+18 $\mathrm{A}_{4,3}$ 72,43 6 432,258 (6.5), (6.7) 2
+18 $\mathrm{A}_{5}$ 60, 5 2 120,5 (6.9), (6.10) 2
+18 $3^{2}.4$ 36,9 3 108,15 (6.11), (6.12) 2
+18 $\mathrm{S}_{3,3}$ 36,10 2 72,22 (6.13) 1
+18 $\mathrm{F}_{21}$ 21,1 1 21,1 (6.2) 1
+18 $\mathrm{Hol}(5)$ 20, 3 1 20,3 (6.15) 1
+18 $\mathrm{QD}_{16}$ 16,8 1 16,8 (6.3) 1
+19 $3^{1+4}_{+}:2.2$ 972, 776 $3^{4}$ - (7.1) 1
+19 $\mathrm{A}_{6}$ 360,118 6 - (2.6), (2.9) 2
+19 $\mathrm{L}_{2}(7)$ 168,42 2 336,114 (7.2) 1
+19 $\mathrm{S}_{5}$ 120,34 2 240, 89 (7.3), (7.6) 2
+19 $\mathrm{M}_{9}$ 72,41 3 216,88 (7.9) 1
+19 $\mathrm{N}_{72}$ 72,40 2 144,117 (7.10) 1
+19 $\mathrm{T}_{48}$ 48,29 1 48,29 (7.12) 1